### v1

# Affine Chord Sobolev Inequalities from Integral Geometry

Abstract.
Functional analogues of isoperimetric inequalities for chord power integrals are established for $\alpha>0$, together with a complete characterization of the equality cases. For $\alpha\in(0,n)$, these inequalities—referred to as chord Sobolev inequalities— serves as a counterpart to the fractional Sobolev inequality of Almgren and Lieb, which corresponds to the functional isoperimetric inequality for chord power integrals with $\alpha\in(-1,0)$.
Moreover, affine isoperimetric inequalities for chord power integrals are derived, that are significantly stronger than their Euclidean counterparts. Extending the framework to functional and affine settings, functional extensions of chord power integrals and radial mean bodies are introduced. The inclusion relations for classical radial mean bodies due to Gardner and Zhang are extended to both log-concave and $s$-concave function, and the extensions are shown to be sharp.
1. Introduction
For $\alpha>-1$, the chord power integral of a convex body $K\subset\mathbb{R}^{n}$ (a convex compact set with non-empty interior), is defined by
$I_{\alpha+1}(K)=\int_{{\rm Aff}(n,1)}|K\cap l|^{\alpha+1}dl,$
where ${\rm Aff}(n,1)$ is the affine Grassmannian consistent of $1$-dimensional lines in $\mathbb{R}^{n}$, and $dl$ is the Haar measure on it. The quantity $|K\cap l|$ represents the length of the intersection. The chord power integral has attracted increasing attention in recent years, mainly in connection with Minkowski-type problems (see for example [10, 17, 40, 48, 49]). In this paper, we study inequalities from the viewpoint of chord power integrals, while previous studies on related inequalities approached their problems from different perspectives.
The isoperimetric inequalities for chord power integrals (cf. [46]) state that, given a convex body $K\subset\mathbb{R}^{n}$,
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.1)
for $\alpha\in(-1,0)\cup(n,\infty)$, and
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.2)
for $\alpha\in(0,n)$. Equality holds in both cases if and only if $K$ is a ball. Throughout, $|K|$ denotes the volume of $K$, that is, the $n$-dimensional Lebesuge measure of $K$, and $B^{n}$ stands for the Euclidean unit ball in $\mathbb{R}^{n}$ with volume $\omega_{n}$.
For $\alpha\in(-1,0)$, the chord power integral $I_{\alpha+1}(K)$ is also referred to as the fractional perimeter, which is closely related to the fractional Sobolev norm (cf. [13, 32, 33]):
$\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a measurable function, and $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{n}$. Replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$, which takes value $1$ on $K$ and 0 elsewhere, yields
$I_{\alpha+1}(K)=\frac{-\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K^{c}}\frac{1}{|x-y|^{n-\alpha}}dxdy,$ (1.3)
which follows from the Blaschke-Petkantschin formula (cf. [46, Theorem 7.2.7]).
Almgren and Lieb [1] established the Pólya–Szegö inequalities for fractional Sobolev norms, and the equality case was subsequently characterized by Frank and Seiringer [13]. A direct consequence is the fractional Sobolev inequality: for $\alpha\in(-1,0)$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$,
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$ (1.4)
where $\|f\|_{p}$ denotes the $L^{p}$ norm of $f$ and $W^{-\alpha,1}(\mathbb{R}^{n})$ denotes the fractional Sobolev space of $L^{1}$ functions with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$. Equality holds in (1.4) if and only if $f$ is a constant multiple of the characteristic function of a ball. The best constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{|\alpha|(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}=\frac{n2^{\alpha}\pi^{\frac{n-\alpha-1}{2}}\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{n}{2}+1)^{\frac{\alpha}{n}}}{|\alpha|\Gamma(\frac{n+\alpha}{2}+1)},$ (1.5)
where $\Gamma$ denotes the gamma function. The constant can be obtained from the results in [12] as described in [16].
Given the sharp fractional Sobolev inequality introduced above—together with its best constant and equality cases—and the geometric isoperimetric inequality for chord power integrals, it becomes natural to regard (1.4) as a functional analogue of (1.1) for $\alpha\in(-1,0)$. More precisely, replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$ reduces the inequality to the isoperimetric inequality for $I_{\alpha+1}(K)$ with $\alpha\in(-1,0)$.
Building on this correspondence, it is natural to ask whether the isoperimetric inequalities established for $\alpha\in(0,n)$ and $\alpha>n$ also admit functional counterparts.
Question. Does there exist a functional inequality for a suitable class of functions $f$ whose geometric instance, corresponding to $f=\operatorname{1}_{K}$, recovers the isoperimetric inequalities for chord power integrals with $\alpha>0$?
Noting that for $\alpha>0$, the quantity $I_{\alpha+1}(K)$ also admits a double integral representation, as shown in [40, Lemma 3.3]:
$I_{\alpha+1}(K)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K}\frac{1}{|x-y|^{n-\alpha}}dxdy.$
This representation suggests a connection with the Hardy-Littlewood-Sobolev inequality (HLS inequality) established by Lieb [30]: for $0<\alpha<n$ and non-negative $f\in L^{p}(\mathbb{R}^{n})$ with $p=2n/(n+\alpha)$,
$\gamma_{n,\alpha}\|f\|_{\frac{2n}{n+\alpha}}^{2}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{f(x)f(y)}{|x-y|^{n-\alpha}}dxdy.$ (1.6)
Equality in (1.6) occurs if and only if $f(x)=a(1+\lambda|x-x_{0}|^{2})^{-(n+\alpha)/2}$ for $x\in\mathbb{R}^{n}$ with $a\geq 0,\lambda>0$ and $x_{0}\in\mathbb{R}^{n}$. This observation leads to the unexpected conclusion that the HLS inequality is not the functional counterpart of (1.2). On further inspection, the essential obstruction appears to be the $2$-homogeneity of the HLS inequality in $f$. Motivated by the 1-homogeneous structure arising in the case $\alpha\in(-1,0)$, we turn to a formulation that reflects the intrinsic homogeneity underlying the geometric setting. This leads naturally to the following 1-homogeneous analogue of (1.6).
Theorem 1 (Chord Sobolev inequalities).
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,$
where the sharp constant $\sigma_{n,\alpha}$ is given by (1.5). Equality holds if and only if $f$ is a constant multiple of the characteristic function of a ball.
With the double integral appearing in Theorem 1 and the fractional Sobolev norm (1.3), we can naturally extend the notion of the chord power integral to functions $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ by setting
$I_{\alpha+1}(f)=\frac{-\alpha(\alpha+1)}{2n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha\in(-1,0),$
and
$I_{\alpha+1}(f)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha>n.$
A more unified formulation can be written as
$I_{\alpha+1}(f)=\int_{0}^{\infty}I_{\alpha+1}(\{f\geq r\})dr,$
which follows from the coarea formula when $\alpha\in(-1,0)$ and from the layer-cake representation of $\min\{f(x),f(y)\}$ when $\alpha>n$. Here $f$ is typically assumed to be an integrable log-concave function, although the log-concavity assumption is not required in general. For instance, when $\alpha\in(-1,0)$, it suffices that $f$ is a non-negative function in the corresponding fractional Sobolev space.
The inequality in Theorem 1 is a functional counterpart of the classical isoperimetric inequality (1.2). It is invariant under translations and rotations, but not under volume-preserving linear transformations. Affine inequalities, which remain invariant under translations and volume-preserving linear transformations, have proved to be very powerful, as they often strengthen the corresponding Euclidean inequalities. A famous example is the Petty projection inequality for convex bodies in $\mathbb{R}^{n}$. It strengthens the Euclidean isoperimetric inequality and also directly implies it. See, for example, [14, 45], and see [7, 19, 23, 29, 36, 37, 39, 47, 51] for extensions in the $L^{p}$ Brunn–Minkowski theory and for more general sets. Recently, Milman and Yehudayoff [42] established isoperimetric inequalities for affine quermassintegrals. These new inequalities are significantly stronger than the isoperimetric inequalities for the classic quermassintegrals, thereby settling a long-standing conjecture of Lutwak [35].
More recently, Haddad and Ludwig [25] established the following sharp affine fractional Sobolev inequality, which is significantly stronger than the classical inequality (1.4).
Affine fractional Sobolev inequalities: For $1<\alpha<0$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$, that is, $f$ with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$,
$\begin{split}2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha}\Big\|\frac{f(\cdot+tu)-f(\cdot)}{t}\Big\|_{1}dt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}\\ \leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.\end{split}$ (1.7)
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality for radially symmetric functions.
In the limit as $\alpha\rightarrow-1^{+}$, the inequality (1.7) recovers the well-known affine Sobolev inequality established by Zhang [51], which is an affine version of the classic $L^{1}$ Sobolev inequality. For a Sobolev function $f$, Zhang [51] defined the polar projection body $\Pi^{*}f$ by
$\|\xi\|_{\Pi^{*}f}=\frac{1}{2}\int_{\mathbb{R}^{n}}|\nabla f(x)\cdot\xi|dx,~~\xi\in\mathbb{S}^{n-1},$
where $\|\cdot\|_{K}$ denotes the gauge function of $K$. Wang [47] extended the definition to functions of bounded variation and obtained the generalized Sobolev-Zhang inequality. The extended version corresponds to the functional form of the Petty projection inequality. Using (1.7), Haddad and Ludwig derived the fractional Petty projection inequalities by taking $f=\operatorname{1}_{K}$.
Motivated by the same central ideas as in their earlier work [25], Haddad and Ludwig [26, 27] established the affine fractional $L^{p}$ Sobolev inequalities as well as affine HLS inequalities. For further results on affine functional inequalities, we refer the reader to [9, 11, 22].
The following Theorem 2 presents an affine version of our functional isoperimetric inequality for chord power integrals, and thus strengthens Theorem 1.
Theorem 2.
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}$ (1.8)
$\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality if $f$ is radially symmetric.
In Section 4, we introduce an anisotropic version of the right side of (1.8), which gives rise to a new star-shaped set ${\rm L}_{\alpha}f$, defined via its radial function by
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,~~u\in\mathbb{S}^{n-1}.$
With the definition, the first inequality in Theorem 2 can be written as
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$ (1.9)
Since both sides of (1.9) are invariant under translations of $f$ and
${\rm L}_{\alpha}(f\circ\phi^{-1})=\phi{\rm L}_{\alpha}f$
for every volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, it follows that inequality (1.9) is indeed affine invariant.
According to (1.1) and (1.2) the isoperimetic inequality for $I_{\alpha+1}(K)$ reverses when $\alpha>n$. Theorem 1 establishes the functional analogue for $\alpha\in(0,n)$. However, the case $\alpha>n$ does not admit a straightforward reversal, which will be discussed in detail in Section 6.
The next result presents the affine inequality we obtain for $\alpha>n$, where ${\rm L}_{\alpha}f$ has been introduced above and $L^{\infty}(\mathbb{R}^{n})$ is the space of essentially bounded measurable functions, equipped with the essential supremum norm $\|\cdot\|_{\infty}$. By eliminating the intermediate term, it serves as the functional analogue of (1.1) for $\alpha>n$.
Theorem 3.
Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid. There is equality in the second inequality if $f$ is radially symmetric.
In Section 7 and Section 8, we relate the set ${\rm L}_{\alpha}f$ to radial mean bodies originally introduced by Gardner and Zhang [18]. For $\alpha>-1$ and a convex body $K\subset\mathbb{R}^{n}$, the radial $\alpha$-th mean body ${\rm R}_{\alpha}K$ is defined via its radial function by
$\rho_{{\rm R}_{\alpha}K}(u)=\Big(\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx\Big)^{1/\alpha},\quad\quad\alpha\neq 0,$
and
$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx,\quad\quad\alpha=0.$
For $\alpha\in(-1,0)$, Haddad and Ludwig [25] showed that the radial mean body ${\rm R}_{\alpha}K$ is proportional to $\Pi^{*,-\alpha}\operatorname{1}_{K}$, where $\Pi^{*,-\alpha}\operatorname{1}_{K}$ denotes the fractional polar projection body arsing from the affine fractional Sobolev inequality (1.7) (see Section 2.6 for details). Of particular interest, for $f$ with bounded variation, Haddad and Ludwig [25] proved that
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|u\|_{\Pi^{*,-\alpha}f}^{-\alpha}=\|u\|_{\Pi^{*}f}.$
In Section 7, we show that ${\rm R}_{\alpha}K$ is proportional to ${\rm L}_{\alpha}\operatorname{1}_{K}$ when $\alpha>n$. Motivated by this relation, we introduce an explicit definition of the functional counterpart ${\rm R}_{\alpha}f$, based on the fractional polar projection body $\Pi^{*,-\alpha}f$ and the star-shaped set ${\rm L}_{\alpha}f$. We note that Haddad and Ludwig [27] also used this notation for radial mean bodies, which in their setting are defined via the fractional $L^{2}$ polar projection body and the star-shaped sets appearing in affine HLS inequalities. Here, we use the same notation ${\rm R}_{\alpha}f$ for an $L^{1}$- extension to keep the notation simple, even though the definition in [27] is slightly different.
In our setting, in particular, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$, we relate ${\rm R}_{\alpha}f$ to the radial mean bodies of its superlevel sets $\{f\geq r\}$. More precisely, for $\alpha>-1$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),\quad\quad\quad\alpha\neq 0,$
and
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r),\quad\quad\alpha=0,$
where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
This unexpected relation naturally leads to the definition of ${\rm R}_{\infty}f$,
${\rm R}_{\infty}f={\rm R}_{\infty}~{\rm supp}(f)={\rm D}~{\rm supp}(f),$
where $f\in L^{1}(\mathbb{R}^{n})$ is a log-concave function with compact support, denoted by ${\rm supp}(f)$. The notation ${\rm D}K$ is the difference body of $K$, defined by ${\rm D}K=\{x-y:x,y\in K\}$. This observation yields the following monotonicity property:
${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f$
and
$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta<\infty$. When $f=\operatorname{1}_{K}$, these relations reduce to the geometric case of Gardner and Zhang [18].
Gardner and Zhang also established the reversed inclusion in [18, Theorem 5.5], which requires a more involved argument. For a convex body $K\subset\mathbb{R}^{n}$, they proved
${\rm D}K\subset c_{n,\beta}{\rm R}_{\beta}K\subset c_{n,\alpha}{\rm R}_{\alpha}K\subset n|K|\Pi^{*}K,$ (1.10)
for $1<\alpha<\beta<\infty$. In each inclusion, equality holds if and only if $K$ is a simplex, denoted by $\triangle$. Here $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$, where $B(\alpha+1,n)$ denotes the Beta function.
In (1.10), when $\beta=n$, the left side inclusion implies the Rogers-Shephard inequality [44], which is
$|{\rm D}K|\leq\binom{2n}{n}|K|.$
When $\alpha=n$, the right side inclusion yields Zhang’s inequality [50] (see also Ball [5]):
$n^{-n}\binom{2n}{n}=|\triangle|^{n-1}|\Pi^{*}\triangle|\leq|K|^{n-1}|\Pi^{*}K|,$
which is a reversed form of Petty projection inequality.
To extend these geometric inequalities to their functional analogues, there are two natural ways to associate a convex body with a log-concave function:
$K\mapsto\operatorname{1}_{K}\quad\quad\text{and}\quad\quad K\mapsto e^{-\|\cdot\|_{K}}.$
Using the second correspondence, Alonso-Gutiérrez, Bernués and González Merino [2] extended Zhang’s inequality to log-concave functions, showing that for a log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq 2^{n}n!\|f\|_{1}^{n+1}|\Pi^{*}f|.$ (1.11)
If $f(0)=\|f\|_{\infty}$, equality holds if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$. For further results on Zhang’s inequality and on radial mean bodies, we refer the reader to [3, 4, 20, 21, 24, 28].
In Section 9, we extend (1.10) to both log-concave and $s$-concave functions, which are stated below as the following two theorems.
Theorem 4.
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$
for $1<\alpha<\beta<\infty$. In each inclusion, there is equality if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$.
When $\alpha=n$, the inclusion
$\frac{1}{\Gamma(n+1)^{1/n}}{\rm R}_{n}f\subset 2\|f\|_{1}|\Pi^{*}f|,$
recovers (1.11). The inclusion relations in Theorem 4 can be deduced from the results of Alonso-Gutiérrez, Bernués and González Merino [2, Lemma 3.2, Remark 5], but the equality cases were not addressed there. In Section 9, we employ a different approach that allows us to establish Theorem 4 together with a complete characterization of equality.
In proving Theorem 4, we first establish an inclusion relation between ${\rm R}_{\alpha}f$ and ${\rm R}_{\beta}f$, and then consider the limits $\alpha\to-1$ and $\beta\to\infty$ to obtain inclusions between ${\rm R}_{\infty}f$ and $\Pi^{*}f$. However, in this limiting process $\beta\to\infty$ degenerates to the singular set $\{o\}$, so that ${\rm R}_{\infty}f$ does not appear in the general case. This motivates us to consider the embedding $K\mapsto\operatorname{1}_{K}$, leading to the following natural question:
Open problem. For a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, does
${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f$
hold? Here $c_{n,\alpha}$ is the constant in (1.10).
Unfortunately, this inclusion fails for general log-concave integrable functions with compact support, as we show by a counterexample in Section 9. This failure leads us naturally to consider $s$-concave functions.
Our main result concerns $s$-concave functions $f\in L^{1}(\mathbb{R}^{n})$, where we denote the positive part of $f$ by $f_{+}(x):=\max\{f(x),0\}$. The equality cases in this setting are fully characterized as well.
Theorem 5.
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$,
${\rm R}_{\infty}f\subset c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f\subset(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$ (1.12)
where $c_{n,\alpha}(s)=(nB(\alpha+1,n+\frac{1}{s}))^{-1/\alpha}$. In each inclusion, there is equality if and only if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
Theorems 4 and 5 share the same general structure and rely on similar methods of proof. As noted above, in Theorem 4, $\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f$ degenerates to the singular set $\{o\}$ as $\beta\to\infty$, so that the set ${\rm R}_{\infty}f$ does not appear. In contrast, Theorem 5 avoids this degeneration, and in the limit $s\to\infty$, the inclusion (1.12) recovers (1.10).
2. Preliminary
We collect some basic notation and results on $L^{p}$ spaces, Schwarz symmetrals, star-shaped sets, log-concave functions, radial mean bodies and fractional polar projection bodies. The books of Gardner [14], Schneider [45], and Schneider and Weil [46] are good general references.
2.1. $L^{p}$ spaces
Let p>1p>-1 and $(X,\mathcal{B},\mu)$ be a measure space. In this paper, we mainly consider the case $X=\mathbb{R}^{n}$ or $X\subset\mathbb{R}^{n}$ with $\mu$ being the $n$-dimensional Lebesgue measure.
For $p\neq 0$ and a measurable function $f:X\rightarrow\mathbb{R}$, let
$\|f\|_{p}=\Big(\int_{X}|f(x)|^{p}d\mu(x)\Big)^{1/p}.$
In particular, if $p\in(-1,0)$ and $\mu(\{x\in X:f(x)=0\})>0$, we set $\|f\|_{p}=+\infty$. The space $L^{p}(X)$ is then defined as
$L^{p}(X)=\{f:X\rightarrow\mathbb{R}:f~\text{is measurable},~\|f\|_{p}<\infty\},$
where the $f=g$ in $L^{p}(X)$ means that they are equal almost everywhere.
When $\mu(x)=1$, the $L^{p}$-norm is monotone in $p$, that is,
$\|f\|_{p}\leq\|f\|_{q},$
for 1<p<q<0-1<p<q<0 or $0<p<q<\infty$. This follows from Jensen’s inequality. In fact, the monotonicity also holds for $p<0<q$. To see this, under the assumption $\mu(X)=1$, we first note the limit
$\lim_{p\rightarrow 0}\log\|f\|_{p}=\int_{X}\log|f(x)|d\mu(x),$
which follows from the continuity of $\|f\|_{p}$ with respect to $p$. Applying Jensen’s inequality then gives
$\|f\|_{p}\leq\int_{X}\log|f(x)|d\mu(x)\leq\|f\|_{q},$
for $p<0<q$.
2.2. Symmetrization
For $E\subset\mathbb{R}^{n}$, the characteristic function $\operatorname{1}_{E}$ is defined by $\operatorname{1}_{E}(x)=1$ for $x\in E$, and $\operatorname{1}_{E}(x)=0$ for $x\notin E$. If $E$ is a Borel set of finite measure, the Schwarz symmetral of $E$, denoted by $E^{\star}$, is the closed, centered ball with the same volume as $E$.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a non-negative measurable function. The superlevel set of $f$ is defined by $\{f\geq t\}=\{x\in X:f(x)\geq t\}$ for $t>0$. We say $f$ is non-zero if $\{f\neq 0\}$ has positive measure. For $f$ with superlevel sets of finite measure, by the layer cake formula,
$f(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}}(x)dt$
for almost every $x\in\mathbb{R}^{n}$.
The Schwarz symmetral of $f$, denoted by $f^{\star}$, is defined as
$f^{\star}(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}^{\star}}(x)dt$
for $x\in\mathbb{R}^{n}$. Hence $f^{\star}$ is radially symmetric and the superlevel set $\{f^{\star}\geq t\}$ has the same volume as $\{f\geq t\}$ for every $t>0$.
One main technique in our proofs is the Riesz rearrangement inequality (see, for example [31, Theorem 3.7]) and the characterization of equality case due to Burchard [8].
Theorem 2.1 (Riesz’s rearrangement inequality.).
For measurable $f,g,k:\mathbb{R}^{n}\rightarrow[0,\infty)$ with superlevel sets of finite measure.
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}f(x)k(x-y)g(y)dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}f^{\star}(x)k^{\star}(x-y)g^{\star}(y)dxdy.$
Theorem 2.2 (Burchard).
Let $A,B$ and $C$ be sets of finite positive measure in $\mathbb{R}^{n}$ and denote by $a,b$ and $c$ the radii of their Schwarz symmetrals $A^{\star},B^{\star}$ and $C^{\star}$. For $|\alpha-\beta|<\gamma<\alpha+\beta$, there is equality in
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{A}(x)\operatorname{1}_{B}(x-y)\operatorname{1}_{C}(y)dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{A^{\star}}(x)\operatorname{1}_{B^{\star}}(x-y)\operatorname{1}_{C^{\star}}(y)dxdy$
if and only if, up to sets of measure zero,
$A=x_{0}+aD,~B=x_{1}+bD,~C=x_{2}+cD,$
where $D$ is a centered ellipsoid, and $x_{0},x_{1}$ and $x_{2}=x_{0}+x_{1}$ are vectors in $\mathbb{R}^{n}$.
2.3. Star-shaped sets and dual mixed volumes
We say $K\subset\mathbb{R}^{n}$ is a star-shaped set (with respect to the origin) if for any $x\in K$ the line segment $[0,x]$ is contained in $K$. The gauge function $\|\cdot\|_{K}:\mathbb{R}^{n}\rightarrow[0,+\infty]$ of a star-shaped set $K$ is defined by
$\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\},$
[…]
$\rho_{K}(x)=\|x\|_{K}^{-1}=\sup\{\lambda\geq 0:\lambda x\in K\}.$
In particular, when $K$ is the unit ball, $\|\cdot\|_{K}$ is the Euclidean norm $|\cdot|$.
For a star-shaped set $K$ with measurable radial function, the $n$-dimensional volume, or the $n$-dimensional Lebesgue measure of $K$ is given by
$|K|=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n}du,$
where $du$ is the spherical Lebesgue measure. The dual mixed volume for star-shaped sets $K$ and $L$ by Lutwak [34] is given by
$\tilde{V}_{\alpha}(K,L)=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{L}(u)^{\alpha}du,$
[…]
For $\alpha\in(0,n)$, and star-shaped sets $K$ and $L$ of finite volume, the dual mixed volume inequality states that
$\tilde{V}_{\alpha}(K,L)\leq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.1)
where equality holds if and only if $K$ and $L$ are dilates, that is, $\rho_{K}(u)=c\rho_{L}(u)$ for some $c\geq 0$ and almost all $u\in\mathbb{S}^{n-1}$. For $\alpha<0$ and $\alpha>n$, the dual mixed volume inequality states that
$\tilde{V}_{\alpha}(K,L)\geq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.2)
where equality holds if and only if $K$ and $L$ are dilates.
2.4. Log-concave functions
A function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is log-concave if $\log f$ is a concave function on $\mathbb{R}^{n}$, with values in $[-\infty,\infty)$. For $s>0$, a function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is $s$-concave if $f^{s}$ is concave on its support.
Lemma 2.3.
[2, Lemma 2.1] Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a log-concave and integrable function. Then the function $g:\mathbb{R}^{n}\rightarrow[0,\infty)$ defined by
$g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy$
is even, log-concave, with $\|g\|_{\infty}=g(o)=\|f\|_{1}$ and ${\rm supp}(g)$ containing the origin point in its interior.
For $\lambda\in[0,1]$, $a,b>0$ and $s>0$, we denote by
[…]
For an $s$-concave function $f$, we have $f((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(f(x),f(y))$ for $\lambda\in[0,1]$. We say $f$ is $(\lambda_{0},s)$-concave, if $f((1-\lambda_{0})x+\lambda_{0}y)\geq\mathcal{M}_{\lambda_{0}}^{s}(f(x),f(y))$ for fixed $\lambda_{0}\in(0,1)$.
The following theorem from Balogh and Kristály [6] gives the equality characterization for Borell-Brascamp-Lieb inequality.
Theorem 2.4.
[6, Theorem 3.1] Let $\lambda\in(0,1)$ and $s>-\frac{1}{n}$. Let $f,g,h:\mathbb{R}^{n}\rightarrow[0,\infty)$ be three non-zero, compactly supported integrable functions satisfying
$h((1-\lambda)x+\lambda y)\geq{\mathcal{M}}_{\lambda}^{s}(f(x),g(y)).$
Then there is equality in
$\int_{\mathbb{R}^{n}}h\geq{\mathcal{M}}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}f,\int_{\mathbb{R}^{n}}g\Big),$
if and only if there is $x_{0}\in\mathbb{R}^{n}$, an $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(f)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ and $c_{0}=(\frac{|{\rm supp}(g)|}{|{\rm supp}(f)|})^{1/n}$ such that up to null measure sets,
${\rm supp}(g)=c_{0}\,{\rm supp}(f)+x_{0},~~\text{and}~~{\rm supp}(h)=(1-\lambda+\lambda c_{0})\,{\rm supp}(f)+sx_{0},$
and for almost all $x\in{\rm supp}(f)$,
$\left\{\begin{aligned} &f(x)=\Phi(x)&&,\\ &g(c_{0}x+x_{0})=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &h((1-\lambda+\lambda c_{0})x+\lambda x_{0})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$. (2.3)
2.5. Convex bodies and radial mean bodies
A convex body in $\mathbb{R}^{n}$ is a convex, compact set with non-empty interior. Let $K\subset\mathbb{R}^{n}$ be a convex body. For $\alpha>-1$ and $\alpha\neq 0$, Gardner and Zhang [18] defined the radial $\alpha$-th mean body of $K$ via its radial function by
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx.$
For $\alpha=0$,
$\log\rho_{{\rm R}_{0}f}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx.$
For $x=x^{\prime}+tu$ with $x^{\prime}\in K|u^{\perp}$, we have
$\int_{K}\rho_{K-x}(u)^{\alpha}dx=\int_{K|u^{\perp}}\int_{t_{1}}^{t_{2}}(t_{2}-t)^{\alpha}dtdx^{\prime},$
where $t_{1}=\min\{t:x^{\prime}+tu\in K\}$ and $t_{2}=\max\{t:x^{\prime}+tu\in K\}$. Since $t_{2}-t_{1}=|K\cap(x^{\prime}+\mathbb{R}u)|$, we then obtain
$\int_{K}\rho_{K-x}(u)^{\alpha}du=\frac{1}{\alpha+1}\int_{K|u^{\perp}}|K\cap(x^{\prime}+\mathbb{R}u)|^{\alpha+1}dx^{\prime}.$
2.6. Fractional polar projection bodies
For $\alpha\in(-1,0)$ and measurable $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$,
$\rho_{\Pi^{*,-\alpha}f}(u)^{-\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$ (2.4)
Then the affine fractional Sobolev inequality (1.7) can be written as
$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.$
Following the calculation in [27], we obtain that
$\rho_{\Pi^{*,-\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}|K\triangle(K+tu)|dt$ (2.5)
$=\int_{0}^{\infty}t^{\alpha-1}\int_{K|u^{\perp}}2\min\{|K\cap(y+\mathbb{R}u),t|\}dydt$
$=2\int_{K|u^{\perp}}\int_{0}^{\infty}t^{\alpha-1}\min\{|K\cap(y+\mathbb{R}u)|,t\}dtdy$
$=2\int_{K|u^{\perp}}\Big(\int_{0}^{|K\cap(y+\mathbb{R}u)|}t^{\alpha}dt+\int_{|K\cap(y+\mathbb{R}u)|}^{\infty}|K\cap(y+\mathbb{R}u)|t^{\alpha-1}dt\Big)dy$
$=-\frac{2}{\alpha(\alpha+1)}\int_{K|u^{\perp}}|K\cap(y+\mathbb{R}u)|^{\alpha+1}dy,$
where $K\triangle L$ is the symmetric difference of $K,L\subset\mathbb{R}^{n}$. Hence,
$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K,$ (2.6)
for $\alpha\in(-1,0)$.
3. Proof of Theorem 1
In this section, we focus on the proof of Theorem 1. Before presenting the proof, we first establish the following lemma, which is based on the Minkowski inequality (see, for example, [31, Theorem 2.4]). Let $F$ be a non-negative, measurable function on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and $p\geq 1$. The Minkowski inequality states that
$\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)^{p}dx\Big)^{1/p}dy\geq\Big(\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)dy\Big)^{p}dx\Big)^{1/p}.$ (3.1)
Equality and finiteness in (3.1) for $p\in(1,\infty)$ imply the existence of measurable functions $\varphi,\psi:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that $F(x,y)=\varphi(x)\psi(y)$ almost everywhere.
Lemma 3.1.
For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
Equality holds if and only if $f=c\operatorname{1}_{E}$ for some measurable $E$ of finite measure and some constant $c>0$.
Proof.
By the Minkowski inequality (3.1),
$\int_{\mathbb{R}^{n}}f(x)^{\frac{n}{n+\alpha}}dx$ $=\int_{\mathbb{R}^{n}}\Big(\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}dx$ (3.2)
$\geq\Big(\int_{0}^{\infty}\Big(\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx\Big)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}$
$=\Big(\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}},$
that is
$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f>t\}|^{\frac{n+\alpha}{n}}dt.$
It is clear that equality is attained when $f=c\operatorname{1}_{E}$. Conversely, if equality holds in (3.2), there are non-negative measurable functions $\varphi$ and $\psi$ such that
$\operatorname{1}_{\{f\geq r\}}(x)=\varphi(x)\psi(t).$
This implies that
$\{(x,r)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)\geq r\}=\{x\in\mathbb{R}^{n}:\varphi(x)>0\}\times\{r\in\mathbb{R}:\psi(r)>0\}.$
Therefore, for every fixed $r\in\mathbb{R}$ such that $r>0$ and $r\leq f(x)$ for some $x$, the superlevel set $\{f\geq r\}$ is a set independent of $r$, which implies that $f=c\operatorname{1}_{E}$ for some measurable set $E$ of finite measure and some constant $c>0$. ∎
We are now ready to prove Theorem 1.
Proof of Theorem 1.
Note that
$\min\{f(x),f(y)\}=\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)dr.$
By Fubini’s theorem and Riesz rearrangement inequality, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdydr$
$\leq\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{\{f\geq r\}^{\star}}(y)}{|x-y|^{n-\alpha}}dxdydr$
$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr,$
where the last equation is from the definition of chord power integrals and $|E^{\star}|=|E|$ for any $E\subset\mathbb{R}^{n}$ with finite measure. The best constant $\sigma_{n,\alpha}$ is given by
$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{\alpha(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}.$
Then Lemma 3.1 implies that
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
When the equality holds, it follows from Lemma 3.1 that $f$ is a characteristic function. Then by the isoperimetric inequality of chord integrals, we have $f=c\operatorname{1}_{B}$, up to a set of measure zero. ∎
4. The star-shaped set ${\rm L}_{\alpha}f$
Building on the approach of Haddad and Ludwig [25, 26, 27], we introduce the following anisotropic version of $I_{\alpha+1}(f)$. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. We consider
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy,$
where $K\subset\mathbb{R}^{n}$ is a star-shaped set with measurable radial function and $\alpha>0$.
Here, we first recall the definition of ${\rm L}_{\alpha}f$:
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt.$
Using Fubini’s theorem and polar coordinate, we obtain that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy$ $=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(y+z),f(y)\}}{\|z\|_{K}^{n-\alpha}}dzdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\rho_{K}(tu)^{n-\alpha}t^{n-1}\min\{f(y),f(y+tu)\}dtdudy$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(y),f(y+tu)\}dydtdu$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}du.$
That is,
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f),$ (4.1)
for measurable $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ and star-shaped set $K\subset\mathbb{R}^{n}$ with measurable radial function. In particular,
$|{\rm L}_{n}f|$ $=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{{\rm L}_{n}f}(u)^{n}du$
$=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}t^{n-1}\min\{f(x),f(x+tu)\}dxdtdu$
$=\frac{1}{n}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy.$
Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. Then the following hold:
For $\alpha\in(0,n)$, the dual mixed volume inequality (2.1) together with (4.1) implies
$\sup\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
For $\alpha>n$, the dual mixed volume inequality (2.2) together with (4.1) yields
$\inf\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
Moreover, in either regime a suitable dilate of ${\rm L}_{\alpha}f$ attains the supremum (respectively, infimum), provided $|{\rm L}_{\alpha}f|$ is finite.
We conclude this section with a result on the convexity of ${\rm L}_{\alpha}f$ when $f$ is log-concave.
Proposition 4.1.
If $f:\mathbb{R}^{n}\to[0,\infty)$ is log-concave and in $L^{1}(\mathbb{R}^{n})$, then ${\rm L}_{\alpha}f$ is a convex set for every $\alpha>0$.
Proof.
For $u\in\mathbb{S}^{n-1}$, we define
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx.$
Since $f$ is log-concave, Lemma 2.3 implies that $g$ is also log-concave.
Note that
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}g(tu)dt.$
Applying a result of Ball [5, Theorem 5], we obtain that
$u\mapsto\left(\int_{0}^{\infty}t^{\alpha-1}g(tu)dt\right)^{-\frac{1}{\alpha}}$
defines a norm on $\mathbb{R}^{n}$, which concludes the proof. ∎
5. Affine isoperimetric inequalities for $\alpha\in(0,n)$
In this section, we aim to establish an affine version of Theorem 1. To this end, we first prove the following crucial lemma, which uses the Riesz rearrangement inequality. We say that $A\subset\mathbb{R}^{n}$ is homothetic to $B\subset\mathbb{R}^{n}$ if there exist $x\in\mathbb{R}^{n}$ and $a>0$ such that $A=x+aB$.
Lemma 5.1.
Let $q>0$ and $K\subset\mathbb{R}^{n}$ be a star-shaped set with measurable radial function and $|K|>0$. For a measurable function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy<\infty,$
there is
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
The equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost all $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to a set of measure zero.
Proof.
[…]
$\|z\|_{K}^{-q}=\int_{0}^{\infty}\operatorname{1}_{t^{-1/q}K}(z)dt.$
Using the Fubini’s theorem, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{\|x-y\|_{K}^{q}}dxdydr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydtdr.$
By the Riesz rearrangement inequality, Theorem 2.2, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq r\}}(y)dxdy,$
which implies that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
If the equality holds, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
for almost all $(r,t)\in(0,\infty)^{2}$. For sufficiently large $t>0$, the assumptions of Theorem 2.2 are fulfilled. Therefore, up to sets of measure zero, we have
$\{f\geq r\}=x+aD,~~t^{-1/q}K=bD,$
where $D$ is an ellipsoid and $x\in\mathbb{R}^{n}$. Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is not dependent on $t$ and $r$, which concludes the proof. ∎
Lemma 5.2.
Let $\alpha\in(0,n)$. For non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
$|{\rm L}_{\alpha}f|\leq|{\rm L}_{\alpha}f^{\star}|.$
If $|{\rm L}_{\alpha}f^{\star}|<\infty$, there is equality if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi$.
Proof.
We first assume that $|{\rm L}_{\alpha}f|<\infty$. By letting $q=n-\alpha$ in Lemma 5.1 and combining it with (4.1), we obtain
$\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\leq\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
With $K={\rm L}_{\alpha}f$ and the dual mixed volume inequality (2.1), we obtain
$|{\rm L}_{\alpha}f|$ $=\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
$\leq\tilde{V}_{\alpha}(({\rm L}_{\alpha}f)^{\star},{\rm L}_{\alpha}f^{\star})$
$\leq|({\rm L}_{\alpha}f)^{\star}|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$=|{\rm L}_{\alpha}f|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}},$
which implies that
$|{\rm L}_{\alpha}f|\leq|{\rm L}_{\alpha}f^{\star}|.$
The equality case follows from Lemma 5.1.
If $|{\rm L}_{\alpha}f|=\infty$, define
$f_{k}(x)=f(x)\operatorname{1}_{kB^{n}}(x),~~k\in\mathbb{N}.$
By the monotone convergence theorem,
$\lim_{k\rightarrow\infty}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f_{k}(x),f_{k}(x+tu)\}dxdt=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
with the convergence being monotone increasing. Applying the monotone convergence theorem once more then gives
$\lim_{k\rightarrow\infty}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{$ $f_{k}(x),f_{k}(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du$
$=\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du.$
That is,
$\lim_{k\rightarrow\infty}|{\rm L}_{\alpha}f_{k}|=|{\rm L}_{\alpha}f|=\infty.$ (5.1)
Since $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$ and $f_{k}$ has compact support, $|{\rm L}_{\alpha}f_{k}|<\infty$. By the definition of the Schwarz symmetral, we have $(f_{k})^{\star}\leq f^{\star}$. Then the first part of the proof implies that
$|{\rm L}_{\alpha}f_{k}|\leq|{\rm L}_{\alpha}(f_{k})^{\star}|\leq|{\rm L}_{\alpha}f^{\star}|.$
Taking the limit as $k\rightarrow\infty$ in the previous inequality, we obtain from (5.1) that
$|{\rm L}_{\alpha}f^{\star}|=\infty,$
which gives the desired result. ∎
Proof of Theorem 2.
By (4.1) and the dual mixed volume inequality (2.1), we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f)\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
There is equality precisely when ${\rm L}_{\alpha}f$ is a ball, which is the case for radially symmetric functions.
For the second inequality, noting that $\|f\|_{\frac{n}{n+\alpha}}=\|f^{\star}\|_{\frac{n}{n+\alpha}}$ and ${\rm L}_{\alpha}f^{\star}$ is a ball, we obtain from Theorem Theorem 1, (4.1), and Lemma 5.2 that
$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$
$=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f^{\star})$
$=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
If equality holds throughout, it follows from Theorem Theorem 1 that $f=c\operatorname{1}_{B}$ with $B$ a ball. By Lemma 5.2, $f=f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. This characterizes the equality cases. ∎
6. Affine isoperimetric inequalities for $\alpha>n$
As mentioned in the introduction, Theorem 1 extends the classical isoperimetric inequality for chord power integrals to function. It also provides an $L^{1}$ analogue of the HLS inequalities, for which both inequalities reverse when $\alpha>n$. In our setting, however, this reversal does not appear in such a simple form here.
Before addressing this nontrivial case, we first present the following lemma, which serves as a counterpart to Lemma 5.1
Lemma 6.1.
Let $q>0$ and $K\subset\mathbb{R}^{n}$ be a star-shaped set with measurable radial function and $|K|>0$. If $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is an integrable function such that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy<\infty,$
then
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$
Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to sets of measure zero.
Proof.
[…]
$\|z\|_{K}^{q}=\int_{0}^{\infty}\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(z)dt.$
Using Fubini’s theorem, we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),$ $f(y)\}\|x-y\|_{K}^{q}dxdy$ (6.1)
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\|x-y\|_{K}^{q}\operatorname{1}_{\{f\geq r\}}(y)dxdydr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)(1-\operatorname{1}_{t^{1/q}K}(x-y))\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt.$
The Riesz rearrangement inequality, Theorem 2.2, implies that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{1/q}K}(x-y)\operatorname{1}$ ${}_{\{f\geq r\}}(y)dxdy$ (6.2)
$\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy.$
Note that $\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx<\infty$ for almost all $r>0$, as $f\in L^{1}(\mathbb{R}^{n})$. Combing (6.2) with the calculation in (6.1), we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$ (6.3)
Moreover, if equality holds in (6.3), then there exists a null set $N\subset(0,\infty)^{2}$ such that equality holds in (6.2) for all $(r,t)\in(0,\infty)^{2}\setminus N$.
For almost every $t\in(0,\infty)$, we have $(r,t)\in(0,\infty)^{2}\setminus N$ for almost every $r>0$. For $t$ sufficiently small, Theorem 2.2 guarantees the existence of a centered ellipsoid $D$ and $x\in\mathbb{R}^{n}$ such that
$\{f\geq r\}=x+aD\quad\text{and }\quad t^{1/q}K=bD.$
Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is independent of $r$ and $t$, which concludes the proof. ∎
With the previous lemma established, for $\alpha>n$, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$ (6.4)
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{\{f^{\star}\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdy$
$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
However, Lemma 3.2 is not applicable in this case because the inequality holds in the opposite direction. We therefore apply a different approach to obtain the result as follows.
Theorem 6.2.
Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
Equality holds if and only if $f=c\operatorname{1}_{B}$ with $B$ a ball and $c>0$.
Proof.
As we calculated in (6.4),
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy\geq\sigma_{n,\alpha}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
By Jensen’s inequality, we have
$\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\geq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|dr\Big)^{\frac{n+\alpha}{n}}=\|f\|_{\infty}^{-\frac{n+\alpha}{n}}\|f\|_{1}^{\frac{n+\alpha}{n}},$ (6.5)
and thus
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$ (6.6)
If equality holds in (6.6), then equality must occur in the application of Jensen’s inequality in (6.5). This in turn implies that $|\{f\geq r\}|$ is constant for almost every $r\in(0,\|f\|_{\infty})$. Equivalently, this happens if and only if $f=c\operatorname{1}_{E}$, where $E$ is a measurable set of finite measure and $c>0$.
When $f=c\operatorname{1}_{E}$, the inequality (6.6) reduces to the geometric case:
$I_{\alpha+1}(E)\geq\Big(\frac{I_{\alpha+1}(B^{n})}{\omega_{n}^{(n+\alpha)/n}}\Big)|E|^{\frac{n+\alpha}{n}},$
with equality if and only if $E$ is a ball. This then completes the characterization of the equality cases. ∎
Lemma 6.3.
Let $\alpha>n$ and $f:\mathbb{R}^{n}\to[0,\infty)$ be integrable. If $|{\rm L}_{\alpha}f|<\infty$, then
$|{\rm L}_{\alpha}f|\geq|{\rm L}_{\alpha}f^{\star}|.$
Equality holds if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
Proof.
By letting $q=n-\alpha$ in Lemma 6.2 and combining it with (4.1), we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\geq n\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
With $K={\rm L}_{\alpha}f$ and the dual mixed volume inequality (2.2), we have
$|{\rm L}_{\alpha}f|$ $\geq\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
$\geq\tilde{V}_{\alpha}(({\rm L}_{\alpha}f)^{\star},{\rm L}_{\alpha}f^{\star})$
$\geq|({\rm L}_{\alpha}f)^{\star}|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}$
$=|{\rm L}_{\alpha}f|^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}},$
which implies that
$|{\rm L}_{\alpha}f|\geq|{\rm L}_{\alpha}f^{\star}|.$
The equality cases follows from Lemma 6.1. ∎
Building on the results established above, we proceed to the proof of Theorem 3.
Proof of Theorem 3.
For the first inequality, if the left side is $\infty$, the inequality holds trivially. We then assume that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy<\infty.$
It follows from (4.1) and the dual mixed volume inequality (2.2) that
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f)\geq n\omega^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
By the equality case in the dual mixed volume inequality, ${\rm L}_{\alpha}f$ must be a ball, which occurs when $f$ is radially symmetric.
For the second inequality, we may assume that $|{\rm L}_{\alpha}f|$ is finite. Note that ${\rm L}_{\alpha}f^{\star}$ is a ball. Then Lemma 6.3 and (4.1) implies that
$n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f^{\star})=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy.$
Combining Theorem 6.2 for the right-hand side with the facts $\|f\|_{1}=\|f^{\star}\|_{1}$ and $\|f\|_{\infty}=\|f^{\star}\|_{\infty}$, we deduce
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
The equality characterization is given by Theorem 6.2 and Lemma 6.3. ∎
7. Radial mean bodies
In this section, we begin by recalling some basic notation and properties of the classical radial mean bodies. Let $K\subset\mathbb{R}^{n}$ be a convex body. For $\alpha>-1$ and $u\in\mathbb{S}^{n-1}$, Gardner and Zhang [18] defined the radial $\alpha$-th mean body of $K$ by its radial function:
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx$
for $\alpha\neq 0$ and by
$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx$
for $\alpha=0$.
They further defined ${\rm R}_{\infty}K$ consistently by
$\rho_{{\rm R}_{\infty}K}(u)=\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}K}(u)=\max_{x\in K}\rho_{K-x}(u),$
so that ${\rm R}_{\infty}K$ coincides with the difference body of $K$, namely
${\rm R}_{\infty}K=DK=K-K.$
As introduced in Section 2, $\rho_{{\rm R}_{\alpha}K}(u)$ can be expressed as the integral of the chord length $|K\cap(x^{\prime}+\mathbb{R}u)|$. It is also closely related to the covariogram of $K$, that is
$g_{K}(x)=|K\cap(K+x)|.$
In fact
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\rho_{DK}(u)}t^{\alpha-1}g_{K}(tu)dt$
for $\alpha>0$.
Note that, for $\alpha>0$,
$\rho_{{\rm L}_{\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{K}(x),\operatorname{1}_{K}(x+tu)\}dxdt$
$=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\operatorname{1}_{K}(x)\operatorname{1}_{K}(x+tu)dxdt$
$=\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt$
$=\int_{0}^{\infty}t^{\alpha-1}g_{K}(tu)dt.$
Hence,
${\rm L}_{\alpha}\operatorname{1}_{K}=\Big(\frac{|K|}{\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.1)
for $\alpha>0$. Recall that in Section 2.5,
$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.2)
for $\alpha\in(-1,0)$.
Combining (7.1) and (7.2) with Theorem 2 and Theorem 3, we obtain the following Pólya–Szegö inequality:
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
for $\alpha\in(0,n)$ and
$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
for $\alpha>n$. These two formulas can also be regarded as the affine isoperimetric inequalities for chord power integrals.
8. Radial mean bodies for log-concave functions
By (7.1) and (7.2), we define the radial mean bodies for log-concave functions, by
${\rm R}_{\alpha}f=\Big(\frac{-\alpha}{2\|f\|_{1}}\Big)^{1/\alpha}\Pi^{*,-\alpha}f$ (8.1)
for $\alpha\in(-1,0)$ and
${\rm R}_{\alpha}f=\Big(\frac{\alpha}{\|f\|_{1}}\Big)^{1/\alpha}{\rm L}_{\alpha}f$
for $\alpha>0$.
The following proposition describes the relation between ${\rm R}_{\alpha}f$ and the radial $\alpha$-th mean body of the superlevel set $\{f\geq r\}$, for $\alpha\in(-1,0)\cup(0,\infty)$.
Proposition 8.1.
Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. For $\alpha\in(-1,0)\cup(0,\infty)$,
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),$
where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
Proof.
First we consider $\alpha>0$. Recall that
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
where
$\min\{f(x),f(x+tu)\}=\int_{0}^{\|f\|_{\infty}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(x+tu)dr.$
By Fubini’s theorem, we then obtain
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dtdr.$
Note that for a convex body $K$,
$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt.$
We have
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}\cdot\frac{|\{f\geq r\}|}{\|f\|_{1}}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
Second, when $\alpha\in(-1,0)$, the definition (8.1) and (2.4) imply that
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$
Since
$|f(x+tu)-f(x)|=\int_{0}^{\infty}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|dr,$
we obtain
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}$ $=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\int_{0}^{\|f\|_{\infty}}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|drdxdt$
$=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dtdr.$
By the calculation in (2.5) and (2.6), we have
$\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}=\frac{-\alpha}{2\{f\geq r\}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dt,$
which implies
$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\frac{2|\{f\geq r\}|}{-\alpha}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
This completes the proof. ∎
By the continuity of the $L^{p}$-norm in $p$, we define
$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r).$ (8.2)
A more explicit formula for ${\rm R}_{0}f$ will be presented in Section 9.
We recall the following convergence result from [25], which shows that as $\alpha\rightarrow-1$, the radial mean body ${\rm R}_{\alpha}f$ converges to the polar projection body of $f$. The opposite endpoint $\alpha=\infty$ will be considered below.
Lemma 8.2.
[25, Theorem 3] For $\alpha\in(-1,0)$ and log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1}$
for $u\in\mathbb{S}^{n-1}$.
Lemma 8.3.
Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave and non-zero. If $f$ has compact support, then
$\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)=\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u),$
for $u\in\mathbb{S}^{n-1}$.
Before proving the lemma below, we recall a result from [18, Theorem 5.5], which states that for every convex body $K$
${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K,$
where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. The complete statement of the theorem will be given in Section 9.
Proof.
Since Jensen’s inequality implies that ${\rm R}_{\alpha}K\subset{\rm D}K$ for $\alpha>-1$, we have
$\rho_{{\rm R}_{\alpha}f}(u)\leq\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha}.$
Letting $\alpha\rightarrow\infty$, we obtain
$\limsup_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\leq\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.3)
On the other hand, by the reverse inclusion ${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K$,
$\rho_{{\rm R}_{\alpha}f}(u)\geq c_{n,\alpha}^{-1}\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f>r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha},$
where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. Note that $\lim_{\alpha\rightarrow\infty}c_{n,\alpha}=1$. Therefore we have
$\liminf_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\geq\sup_{r}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.4)
Combining (8.3) with (8.4), we then conclude the proof. ∎
Definition 8.4.
Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave with compact support. The body ${\rm R}_{\infty}f$ is given by
${\rm R}_{\infty}f={\rm D}{\rm supp}(f).$
If $f$ is not compactly supported, ${\rm R}_{\infty}f=\mathbb{R}^{n}$.
Recall that in the geometric case, Jensen’s inequality implies that
${\rm R}_{\alpha}K\subset{\rm R}_{\beta}K\subset{\rm D}K,$ (8.5)
where Gardner and Zhang [18] further showed that
$\Pi^{*}K\subset((1+\alpha)|K|)^{1/\alpha}{\rm R}_{\alpha}K\subset((1+\beta)|K|)^{1/\beta}{\rm R}_{\beta}K\subset{\rm D}K$ (8.6)
for $1<\alpha<\beta$.
Applying Jensen’s inequality together with (8.5), we immediately obtain
${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta$. The following theorem provides a functional analogue of (8.6).
Theorem 8.5.
For non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
for $1<\alpha<\beta$.
Proof.
If $f$ is not compactly supported, the right inclusion is trivial since ${\rm R}_{\infty}f=\mathbb{R}^{n}$. We now assume that $f$ is with compact support.
By Proposition 8.1,
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}(1+\alpha)|\{f\geq r\}|\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
$=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\alpha,r}{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha},$
where $l_{\alpha,r}=((1+\alpha)|\{f\geq r\}|)^{1/\alpha}$ It follows from (8.6) that
$l_{\alpha,r}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)\leq l_{\beta,r}\rho_{{\rm R}_{\beta}\{f\geq r\}}(u),$
for $1<\alpha<\beta$.
Jensen’s inequality then implies that
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
$\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\beta}dr\Big)^{1/\beta}$
$=\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}\rho_{{\rm R}_{\beta}f}(u),$
which gives
$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f.$
Letting $\beta\rightarrow\infty$, together with $\lim_{\beta\rightharpoonup\infty}(1+\beta)^{1/\beta}=1$, we obtain the right inclusion. Letting $\alpha\rightarrow-1$, by Lemma 8.2, we obtain the left inclusion. ∎
9. Reversed inequalities
In the previous section, we recalled the monotonicity relations (8.5) and (8.6) for radial mean bodies and extended them to their functional analogues. In this section, we turn to the reverse inclusion, which is considerably more delicate.
We first recall the following theorem established by Gardner and Zhang [18]. In particular, for $\alpha=n$,
$c_{n,n}{\rm R}_{n}K\subset n|K|\Pi^{*}K$
which yields Zhang’s inequality [50].
Theorem 9.1.
[18, Theorem 5.5] Let $K$ be a convex body in $\mathbb{R}^{n}$. If $1<\alpha<\beta$, then
${\rm D}K\subset c_{n,\beta}{\rm R}_{\beta}K\subset c_{n,\alpha}{\rm R}_{\alpha}K\subset n|K|\Pi^{*}K.$
In each inclusion equality holds if and only if $K$ is a simplex.
To extend Theorem 9.1 to the functional setting, we rely on the following result from [27]. It generalizes the classical result of Marshall, Olkin and Proschan [41], whose simple proof was later given by Milman and Pajor [43].
Lemma 9.2 ([27], Lemma 17).
Let $\omega:[0,\infty)\to[0,\infty)$ be decreasing with
$0<\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt<\infty$
for $\alpha>0$ and
$0<\int_{0}^{\infty}t^{\alpha-1}(\omega(0)-\omega(t))dt<\infty$
for $\alpha\in(-1,0)$. If $\varphi:[0,\infty)\to[0,\infty)$ is non-zero, with $\varphi(0)=0$, and such that $t\mapsto\varphi(t)$ and $t\mapsto{\varphi(t)}/{t}$ are increasing on $(0,\infty)$, then
$\zeta(\alpha)=\left\{\begin{aligned} &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}\omega(\varphi(t))dt}{\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt}\right)^{1/\alpha}&&\text{for}~~\alpha>0,\\ &\exp\left(\int_{0}^{\infty}\frac{\omega(\varphi(t))-\omega(t)}{t\omega(0)}dt\right)&&\text{for}~~\alpha=0,\\ &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}(\omega(\varphi(t))-\omega(0))dt}{\int_{0}^{\alpha-1}(\omega(t)-\omega(0))dt}\right)^{1/\alpha}&&\text{for}~~-1<\alpha<0.\end{aligned}\right$.
is a continuous, decreasing function of $\alpha$ on $(-1,\infty)$. Moreover, $\zeta$ is constant on $(-1,\infty)$ if $\varphi(t)=\lambda t$ on $0,\infty)$ for some $\lambda>0$.
In particular, taking $\omega(t)=e^{-t}$ and $\omega(t)=(1-st)_{+}^{1/s}$ with $s>0$ yields the classical analytic continuation formulas for the Gamma and Beta functions,
[…]
and
[…]
We note here, in the setting of Lemma 9.2, that if $\zeta(\alpha)$ is constant on $(-1,\infty)$, say $\zeta(\alpha)=\lambda$, then $\varphi(t)=\lambda t$ for all $t\geq 0$. When $\omega(t)$ is strictly decreasing, this conclusion follows from the proof of Milman and Pajor, but for completeness, we include a proof below.
Lemma 9.3.
Let $\omega$, $\varphi$, and $\zeta$ be as in Lemma 9.2. Suppose $\omega$ is strictly decreasing and $1<\alpha<\beta<\infty$. Then $\zeta(\alpha)=\zeta(\beta)$ implies $\varphi(t)=\lambda t$ with $\lambda=\zeta(\alpha)^{-1}$.
Proof.
Let $\lambda=\zeta(\alpha)^{-1}$. We first consider $0<\alpha<\beta<\infty$. For $\alpha>0$, we then have
$\int_{0}^{\infty}\omega(\lambda t)t^{\alpha-1}dt=\int_{0}^{\infty}\omega(\varphi(t))t^{\alpha-1}dt.$
We set $\Psi(r)=\int_{r}^{\infty}(\omega(\lambda t)-\omega(\varphi(t)))dt$, where $\Psi(0)=\Psi(\infty)=0$. Since $\varphi(t)/t$ is increasing while $\omega$ is decreasing, we have
$\omega(\varphi(t))>\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}<\lambda,$
and
$\omega(\varphi(t))\leq\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}\geq\lambda.$
This implies the sign of $\omega(\lambda r)-\omega(\varphi(r))=-\Psi^{\prime}(r)$. Therefore, $\Psi$ is first increasing and then decreasing, which yields that $\Psi(r)\geq 0$ for all $r\geq 0$. That is
$\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\geq\int_{r}^{\infty}\omega(\varphi(t))dt.$
For $\beta>\alpha>0$, we have
$\int_{0}^{\infty}\omega(\varphi(t))t^{\beta-1}dt$ $=(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\varphi(t))r^{\alpha-1}dt\Big)dr$
$\leq(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\Big)dr$
$=\int_{0}^{\infty}\omega(\lambda t)t^{\beta-1}dt=\frac{1}{\lambda^{\beta}}\int_{0}^{\infty}\omega(t)t^{\beta-1}dt,$
which implies $\zeta(\alpha)\geq\zeta(\beta)$. As we assumed, $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, equality holds in the inequality above, which implies that $\Psi(r)=0$ for all $r\geq 0$. Then we have
$\Psi^{\prime}(r)=\omega(\varphi(t))-\omega(\lambda t)=0.$
Since $\omega$ is strictly decreasing, we obtain $\varphi(t)=\lambda t$.
For $1<\alpha<\beta<0$, the proof is the same. Then we consider $1<\alpha<0<\beta<\infty$. Since $\zeta$ is decreasing, we have
$\zeta(\beta)\leq\zeta(0)\leq\zeta(\alpha).$
As assumed $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, we obtain that $\zeta$ is a constant on $(0,\beta)$, by the first case, we have $\varphi(t)=\lambda t$. ∎
9.1. Results for log-concave functions
We begin this subsection by recalling a result from [2], which will serve as a main tool in the equality characterization in Theorem 3.
Lemma 9.4.
[2, Lemma 4.2] Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave such that $f(o)=\|f\|_{\infty}$. Let
$g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy.$
Then for every $x\in\mathbb{R}^{n}$ and $\lambda\in[0,1]$, $g(\lambda x)=g(o)^{1-\lambda}g(x)^{\lambda}$ if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$ with $\triangle$ a simplex containing the origin.
For the reader’s convenience, we restate Theorem 4 below before giving the proof.
Theorem 9.5.
Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. Then
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$ (9.2)
for $1<\alpha<\beta<\infty$. In each inclusion, equality holds if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ a simplex containing the origin.
Proof.
By Lemma 2.3,
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
is an even, log-concave function and attains maximum at $t=0$.
For fixed $u\in\mathbb{S}^{n-1}$, we apply Lemma 9.2 to $\omega(t)=g(0)e^{-t}$ and $\varphi(t)=-\log(g(tu)/g(0))$. We observe that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)e^{-t}dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
for $\alpha>0$.
For $\alpha\in(-1,0)$, note that
$g(0)-g(tu)$ $=\|f\|_{1}-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ (9.3)
$=\frac{1}{2}\int_{\mathbb{R}^{n}}f(x)dx+\frac{1}{2}\int_{rn}f(x+tu)dx-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x)-f(x+tu))_{+}dx+\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x+tu)-f(x))_{+}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}|f(x)-f(x+tu)|dx,$
where $h(x)_{+}$ denotes the positive part of the function $h(x)$. Then we have
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)(1-e^{-t})dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
for $\alpha\in(-1,0)$.
Therefore Lemma 9.2 implies that
$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f.$ (9.4)
Letting $\alpha\rightarrow-1$, Lemma 8.2 gives the right inclusion. Since $\omega(t)=e^{-t}$ is strictly decreasing, by Lemma 9.3, equality holds in (9.4) if and only if $g(tu)=g(0)e^{-\lambda t}$ for some $\lambda>0$. Then Lemma 9.4 shows that $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ is a simplex containing the origin. ∎
In fact, if we set $f(x)=e^{-\|x\|_{K}}$ for a convex body $K\subset\mathbb{R}^{n}$, then inclusion (9.2) in Theorem 9.5 reduces to the geometric case stated in Theorem 9.1. However, in this functional extension, the endpoint ${\rm R}_{\infty}f$ is $\mathbb{R}^{n}$, so the inclusion no longer extends to that limit case.
Furthermore, combining Lemma 9.2 with (9.4), we obtain an explicit expression for $\zeta(0)$ by the continuity of $\zeta(\alpha)$, where
$\log\zeta(0)=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt.$ (9.5)
When $f=\operatorname{1}_{K}$, the quantity $\log\zeta(0)=\log\rho_{{\rm R}_{0}K}(u)$. The next result confirms that the formula (9.5) is consistent with the definition of ${\rm R}_{0}f$ (8.2), namely $\rho_{{\rm R}_{0}f}(u)=\zeta(0)$ with $u\in\mathbb{S}^{n-1}$ fixed.
Proposition 9.6.
For fixed $u\in\mathbb{S}^{n-1}$ and non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\rho_{{\rm R}_{0}f}(u)=\zeta(0).$
Proof.
By (9.5),
$\log\zeta(0)=$ $-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt$
$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\int_{0}^{\|f\|_{\infty}}|\frac{\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}d\mu_{f}(r)-e^{-t}\Big)dt$
$=\int_{0}^{\|f\|_{\infty}}\Big(-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt\Big)d\mu_{f}(r).$
Note that
$\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)$ $=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{|\{f\geq r\}|}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{\{f\geq r\}}(x),\operatorname{1}_{\{f\geq r\}}(x+tu)\}dx-e^{-t}\Big)dt$
$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt,$
where the first equation is from (9.5). Therefore, we have
$\log\zeta(0)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r)=\log\rho_{{\rm R}_{0}f}(u),$
which concludes the proof. ∎
9.2. Results for $s$-concave functions
In this subsection, we focus on the case of $s$-concave functions. The following lemma is a direct corollary from the Borell-Brascamp-Lieb inequality, which is also mentioned in [15, Theorem 11.2].
Lemma 9.7.
For $s$-concave $f\in L^{1}(\mathbb{R}^{n})$ and fixed $u\in\mathbb{S}^{n-1}$,
$t\mapsto g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu))\}dx$
is a $s/(ns+1)$-concave function on $\mathbb{R}$.
Proof.
We denote by
$N_{f}(x,t)=\min\{f(x),f(x+tu)\}.$
First, we prove that $N_{f}(x,t)$ is $s$-concave on $\mathbb{R}^{n}\times\mathbb{R}$. For $x_{0},x_{1}\in\mathbb{R}^{n}$, $t_{0},t_{1}\in\mathbb{R}$ and $\lambda\in[0,1]$, we denote $x_{\lambda}=(1-\lambda)x_{0}+\lambda x_{1}$ and $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$. Note that
$N_{f}(x_{\lambda},t_{\lambda})^{s}=\min\{f(x_{\lambda})^{s},f(x_{\lambda}+t_{\lambda}u)^{s}\}.$
Since $f$ is $s$-concave and $\min\{\cdot,\cdot\}$ is a concave function on $\mathbb{R}\times\mathbb{R}$, we have
$N_{f}(x_{\lambda},t_{\lambda})^{s}$ $\geq\min\{(1-\lambda)f(x_{0})^{s}+\lambda f(x_{1})^{s},(1-\lambda)f(x_{0}+t_{0}u)^{s}+\lambda f(x_{1}+t_{1}u)^{s}\}$
$\geq(1-\lambda)\min\{f(x_{0})^{s},f(x_{0}+t_{0}u)^{s}\}+\lambda\min\{f(x_{1})^{s},f(x_{1}+t_{1}u)^{s}\}$
$=(1-\lambda)N_{f}(x_{0},t_{0})^{s}+\lambda N_{f}(x_{1},t_{1})^{s},$
which implies the $s$-concavity of $N_{f}$.
For $\lambda\in[0,1]$, and fixed $t_{0},t_{1}\in\mathbb{R}$, we set
$F(x)=N_{f}(x,t_{0}),~~G(x)=N_{f}(x,t_{1}),~~H(x)=N_{f}(x,t_{\lambda}),$
where $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$. By the $s$-concavity of $F$, we have
$H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y)).$
The Borell-Brascamp-Lieb inequality (2.3) then implies that
$\int_{\mathbb{R}^{n}}H\geq\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big),$ (9.6)
that is, the $\frac{s}{ns+1}$-concavity of $g(tu)$ in $t$. ∎
We begin the proof of Theorem 5 by stating its inclusion result as the following theorem; the equality case will be addressed separately.
Theorem 9.8.
Let $f\in L^{1}(\mathbb{R}^{n})$ be $s$-concave with compact support. If $1<\beta<\alpha$, then
${\rm R}_{\infty}f\subset c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f\subset(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$ (9.7)
where $c_{n,\alpha}(s)=((n+\frac{1}{s})B(\alpha+1,n+\frac{1}{s}))^{-1/\alpha}$.
Proof.
By Lemma 9.7, the function $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ is $s^{\prime}$-concave with $s^{\prime}=s/(ns+1)$. We apply Lemma 9.2 with $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$.
Therefore, we obtain that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}(1-s^{\prime}t)_{+}^{1/s^{\prime}}dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
for $\alpha>0$ and
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}((1-s^{\prime}t)_{+}^{1/s^{\prime}}-1)dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
for $1<\alpha<0$, where we use (9.1) and (9.3). Note that
$\Big(\alpha B(\alpha,1+\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=\Big(nB(\alpha+1,\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=c_{n,\alpha}(s).$
Lemma 9.2 then implies that
$c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f,$ (9.8)
for $1<\alpha<\beta$.
Since Lemma 8.3 ensures that ${\rm R}_{\beta}f\to{\rm R}_{\infty}f$ and $c_{n,\beta}(s)\to 1$ as $\beta\rightarrow\infty$, the left inclusion follows directly from (9.8).
By Lemma 8.2 and $c_{n,\alpha}(s)\sim\frac{\alpha+1}{n+1/s}$ when $\alpha\rightarrow-1$, we have
$\lim_{\alpha\rightarrow-1^{+}}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=(n+\frac{1}{s})\|f\|_{1}\rho_{\Pi^{*}f}(u).$
which yields the right inclusion. ∎
When $s=\infty$, Theorem 9.8 reduces to the geometric case in Theorem 9.1 by Gardner and Zhang. In fact, following the argument of our proof, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, if one could show that $g(tu)^{1/n}$ is concave, then the following chain of inclusions would hold:
${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f,$
for $1<\alpha<\beta<\infty$.
This would yield precisely the functional counterpart of Theorem 9.1 that one would expect. However, $g(tu)^{1/n}$ is not necessarily concave for a general log-concave function $f$. A counterexample will be provided below.
Let $n=1$ and
[…]
It is clear that $\log f(x)=\log x+\log 2$ for $t\in[0,1]$, which is concave. Then we have
[…]
Note that
$\frac{1}{4}=g(\frac{1}{2})<\frac{1}{2}(g(0)+g(1))=\frac{1}{2},$
which shows that $g$ is not concave. In this case, we consider $\beta=1$ and $\alpha=1/2$ and obtain that
$c_{1,1}{\rm R}_{1}f=\Big[-\frac{2}{3},\frac{2}{3}\Big],\quad c_{1,\frac{1}{2}}{\rm R}_{\frac{1}{2}}f=\Big[-\frac{16}{25},\frac{16}{{25}}\Big].$
which is not consistent with the desired result $c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f$.
The following theorem describes the sharpness of the inclusion result in Theorem 5.
Theorem 9.9.
Equality holds in (9.7) if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $x_{0}\in\mathbb{R}^{n}$, $c>0$ and $\triangle$ is a simplex containing the origin.
Proof.
Note that ${\rm R}_{\alpha}f$ is invariant under scaling and translation of $f$. Without loss of generality, we may assume $f(x)=(1-\|x\|_{\triangle})_{+}^{1/s}$. A direct computation then gives
$\|f\|_{1}=B(\frac{1}{s},n+1)\frac{|\triangle|}{s}.$
We use the same notation for $g(tu)$ with a fixed $u\in\mathbb{S}^{n-1}$, and we have
$g(tu)$ $=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
$=\int_{0}^{1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dr$
$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}|(1-r)\triangle\cap((1-r)\triangle+tu)|dr$
$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}(1-r)^{n}|\triangle\cap(\triangle+\frac{t}{1-r}u)|dr.$
By [18, Lemma 3.4], $|\triangle\cap(\triangle+tu)|^{1/n}$ is linear in $t$. Therefore, there exists $\lambda>0$ such that
$|\triangle\cap(\triangle+\frac{t}{1-r})u|=|\triangle|(1-\lambda\frac{t}{1-r})_{+}^{n}.$
A straightforward, though somewhat tedious, calculation then shows that
$g(tu)=\frac{1}{s}B(\frac{1}{s},n+1)|\triangle|(1-\lambda t)^{\frac{ns+1}{s}}.$
We set $s^{\prime}=s/(ns+1)$. Then, for $1<\alpha<0$ and $\alpha>0$, by taking $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
$c_{n,\alpha}(s){\rm R}_{\alpha}f=\frac{1}{s^{\prime}}\zeta(\alpha)=\frac{1}{s^{\prime}}\Big(\frac{\int_{0}^{\infty}t^{\alpha-1}(1-\lambda t)_{+}^{1/s^{\prime}}dt}{\int_{0}^{\infty}t^{\alpha-1}(1-s^{\prime}t)^{1/s^{\prime}}dt}\Big)^{1/\alpha}=\frac{1}{\lambda},$
where the first equality is from thex calculation in Theorem 9.8.
This implies that
$c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,\quad\quad-1<\alpha<\beta<\infty.$
Letting $\alpha\rightarrow-1$ and $\beta\rightarrow\infty$, we have
${\rm R}_{\infty}f=c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f=(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$
which concludes the proof. ∎
We conclude the section by showing the necessary condition for the equality cases in Theorem 5. The following lemma is the key result for the equality characterization.
Lemma 9.10.
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ non-zero and $s$-concave with compact support. If
$g(tu)^{\frac{s}{ns+1}}=\Big(\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx\Big)^{\frac{s}{ns+1}}$
is linear in $t$ for any fixed $u\in\mathbb{S}^{n-1}$, then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
Proof.
We assume, without loss of generality, that $\|f\|_{\infty}=1$. Since
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x-tu)\}dx,$
for convenience of notations in the following proof, we set $N_{f}(x,t)=\min\{f(x),f(x-tu)\}$ here. The functions $F,G,H$ are as defined by
$F(x)=N_{f}(x,0),~~G(x)=N_{f}(x,t),~~H(x)=N_{f}(x,\lambda t),$
which satisfies that $H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y))$.
The linearity of $g(tu)^{\frac{s}{ns+1}}$ then implies that
$\int_{\mathbb{R}^{n}}H=\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big).$
That is, we have the equality in the Borell-Brascamp-Lieb inequality for $F,G,H$.
By Theorem 2.4, there exists $c_{0}=\Big(\frac{|{\rm supp}(G)|}{|{\rm supp}(F)|}\Big)^{1/n}$ and $y_{0}\in\mathbb{R}^{n}$ such that
${\rm supp}(G)=c_{0}\,{\rm supp}(F)+y_{0},$ (9.9)
while the definition of $G$ shows
${\rm supp}(G)={\rm supp}(F)\cap({\rm supp}(F)+tu).$ (9.10)
Since $u\in\mathbb{S}^{n-1}$ is arbitrary fixed, (9.9) and (9.10) shows that ${\rm supp}(F)\cap({\rm supp}(F)+tu)$ is homothetic to ${\rm supp}(F)$ for all $u\in\mathbb{S}^{n-1}$. By [18, Lemma 3.4], we have ${\rm supp}(F)$ is a simplex $\triangle$, where we assume that $|\triangle|$ is the same as the volume of the standard simplex $\triangle_{n}$ for convenience.
In fact, we only need to consider when $\triangle$ is the standard simplex, that is,
$\triangle=\triangle_{n}=\Big\{x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:x_{i}\geq 0,~~\sum_{i=1}^{n}x_{i}\leq 1\Big\}.$
Otherwise, there exists $\phi\in{\rm SL}(n)$ such that ${\rm supp}(f\circ\phi^{-1})=\triangle_{n}$ and we consider $f\circ\phi^{-1}$ instead of $f$. In this case, by (9.9) and (9.10), we have
$c_{0}(t)=1-t\sum_{i=1}^{n}u_{i},~~\text{and}~~y_{0}(t)=t(u_{1}^{+},\ldots,u_{n}^{+}),$
where $u_{i}^{+}=\max\{0,u_{i}\}$ and we denote by $a_{0}=\sum_{i=1}^{n}u_{i}$. Therefore, we can choose $u\in\mathbb{S}^{n-1}$ with $u_{i}\geq 0$, such that $y_{0}(t)=tu$.
Moreover, by Theorem 2.4, there exists a $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ satisfying,
$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{0}x+tu)=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda tu)=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$.
In fact, $\Phi(x)=F(x)=f(x)$ by the definition of $F$. Then we have
$G((1-a_{0}t)x+tu)^{s}=\min\{f((1-a_{0}t)x)^{s},f((1-a_{0}t)x+tu)^{s}\}=(1-a_{0}t)f(x)^{s}.$
Note that $(1-a_{0}t)f(x)^{s}\neq f((1-a_{0}t)x)^{s}$ since $f^{s}$ is concave. Therefore,
$f((1-a_{o}t)x+tu)^{s}=(1-a_{0}t)f(x)^{s}$ (9.11)
holds for $0\leq t\leq 1/\sum u_{i}$. Equation (9.11), together with $f\geq 0$ and the $s$-concavity of $f$ , imply that
$f\Big(\frac{u}{a_{0}}\Big)^{s}=0.$ (9.12)
We set $\bar{x}=(\bar{x}_{1},\ldots,\bar{x}_{n})\in\triangle_{n}$ such that $f(\bar{x})=\|f\|_{\infty}=1$. Note that if $\sum_{i}\bar{x}_{i}\in(0,1)$, we will have
$f\Big(\frac{\bar{x}}{\sum_{i}\bar{x}_{i}}\Big)^{s}=0\geq(1-\sum_{i}\bar{x}_{i})f(\bar{x})^{s}+\Big(\sum_{i}\bar{x}_{i}\Big)f(o)^{s}\geq 1-\sum_{i}\bar{x}_{i}>0,$ (9.13)
which cause a contradiction. Thus we have $\sum_{i}\bar{x}_{i}=1$ or $\bar{x}=o$. If $\sum_{i}\bar{x}_{i}=1$, equation (9.12) implies that $f(\bar{x})=0$ which contradicts to the choice of $\bar{x}$. Thus we have $\bar{x}=o$. Letting $x=\bar{x}=o$ in (9.11) then gives
$f(tu)^{s}=1-a_{0}t=1-\|tu\|_{\triangle_{n}}.$
Since ${\rm supp}(f)=\triangle_{n}$, we then have
$f(x)=(1-\|x\|_{\triangle_{n}})_{+}^{1/s},$
which concludes the proof. ∎
We conclude this section by the following theorem, which provides the full equality characterization in Theorem 5.
Theorem 9.11.
Let $s>0$ and $c_{n,\alpha}(s)$ be defined in Theorem 5. Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, and $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, if
$c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,$
then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
Proof.
Let $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ with fixed $u\in\mathbb{S}^{n-1}$. By Theorem 9.8, letting $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
$\zeta(\alpha)=s^{\prime}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u),$
where $s^{\prime}=s/(ns+1)$. Then $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f$ implies that $\zeta(\alpha)=\zeta(\beta)$. Since $\omega$ is strictly decreasing on its support, by Lemma 9.3, we obtain $\lambda=\zeta(\alpha)^{-1}$ such that
$\varphi(t)=\frac{1}{s^{\prime}}\Big(1-\Big(\frac{g(tu)}{g(0)}\Big)^{s^{\prime}}\Big)=\lambda t.$
This implies that $g(tu)^{s^{\prime}}$ is linear in $t$. Then we conclude the proof by Lemma 9.10. ∎
Acknowledgements: This research was funded in whole or in part by the Austrian Science Fund (FWF) doi/10.55776/37030. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v2

@@
-Affine Chord Sobolev Inequalities from Integral Geometry
+Affine chord Sobolev inequalities and radial mean bodies for functions
 Abstract.
-Functional analogues of isoperimetric inequalities for chord power integrals are established for $\alpha>0$, together with a complete characterization of the equality cases.
-For $\alpha\in(0,n)$, these inequalities—referred to as chord Sobolev inequalities— serves as a counterpart to the fractional Sobolev inequality of Almgren and Lieb, which corresponds to the functional isoperimetric inequality for chord power integrals with $\alpha\in(-1,0)$.
-Moreover, affine isoperimetric inequalities for chord power integrals are derived, that are significantly stronger than their Euclidean counterparts.
-Extending the framework to functional and affine settings, functional extensions of chord power integrals and radial mean bodies are introduced.
-The inclusion relations for classical radial mean bodies due to Gardner and Zhang are extended to both log-concave and $s$-concave function, and the extensions are shown to be sharp.
+Affine isoperimetric inequalities for the functional radial mean bodies are derived from the new affine chord Sobolev inequalities, which extend the recent affine isoperimetric inequalities of Haddad and Ludwig [24] from convex bodies to functions.
+The affine chord Sobolev inequalities further imply a strengthening of the Euclidean chord Sobolev inequalities introduced in [4].
+Moreover, for $s$-concave functions $f$ with compact support and $s>0$, a parameter-dependent monotonicity property of the functional radial mean body ${\rm R}_{\alpha}f$ is obtained: ${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f$ for $1<\alpha<\beta$, and, after suitable normalization, the reverse inclusion also holds.
+These sharp results generalize the corresponding monotonicity for geometric radial mean bodies established by Gardner and Zhang [18].
 1. Introduction
-For $\alpha>-1$, the chord power integral of a convex body $K\subset\mathbb{R}^{n}$ (a convex compact set with non-empty interior), is defined by
-$I_{\alpha+1}(K)=\int_{{\rm Aff}(n,1)}|K\cap l|^{\alpha+1}dl,$
-where ${\rm Aff}(n,1)$ is the affine Grassmannian consistent of $1$-dimensional lines in $\mathbb{R}^{n}$, and $dl$ is the Haar measure on it.
-The quantity $|K\cap l|$ represents the length of the intersection.
-The chord power integral has attracted increasing attention in recent years, mainly in connection with Minkowski-type problems (see for example [10, 17, 40, 48, 49]).
-In this paper, we study inequalities from the viewpoint of chord power integrals, while previous studies on related inequalities approached their problems from different perspectives.
-The isoperimetric inequalities for chord power integrals (cf. [46]) state that, given a convex body $K\subset\mathbb{R}^{n}$,
-$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.1)
-for $\alpha\in(-1,0)\cup(n,\infty)$, and
-$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{\frac{n+\alpha}{n}},$ (1.2)
-for $\alpha\in(0,n)$.
-Equality holds in both cases if and only if $K$ is a ball.
-Throughout, $|K|$ denotes the volume of $K$, that is, the $n$-dimensional Lebesuge measure of $K$, and $B^{n}$ stands for the Euclidean unit ball in $\mathbb{R}^{n}$ with volume $\omega_{n}$.
-For $\alpha\in(-1,0)$, the chord power integral $I_{\alpha+1}(K)$ is also referred to as the fractional perimeter, which is closely related to the fractional Sobolev norm (cf. [13, 32, 33]):
-$\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$
-where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a measurable function, and $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{n}$.
-Replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$, which takes value $1$ on $K$ and 0 elsewhere, yields
-$I_{\alpha+1}(K)=\frac{-\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K^{c}}\frac{1}{|x-y|^{n-\alpha}}dxdy,$ (1.3)
-which follows from the Blaschke-Petkantschin formula (cf. [46, Theorem 7.2.7]).
-Almgren and Lieb [1] established the Pólya–Szegö inequalities for fractional Sobolev norms, and the equality case was subsequently characterized by Frank and Seiringer [13].
-A direct consequence is the fractional Sobolev inequality: for $\alpha\in(-1,0)$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$,
-$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$ (1.4)
-where $\|f\|_{p}$ denotes the $L^{p}$ norm of $f$ and $W^{-\alpha,1}(\mathbb{R}^{n})$ denotes the fractional Sobolev space of $L^{1}$ functions with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$.
-Equality holds in (1.4) if and only if $f$ is a constant multiple of the characteristic function of a ball.
-The best constant $\sigma_{n,\alpha}$ is given by
-$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{|\alpha|(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}=\frac{n2^{\alpha}\pi^{\frac{n-\alpha-1}{2}}\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{n}{2}+1)^{\frac{\alpha}{n}}}{|\alpha|\Gamma(\frac{n+\alpha}{2}+1)},$ (1.5)
-where $\Gamma$ denotes the gamma function. The constant can be obtained from the results in [12] as described in [16].
-Given the sharp fractional Sobolev inequality introduced above—together with its best constant and equality cases—and the geometric isoperimetric inequality for chord power integrals, it becomes natural to regard (1.4) as a functional analogue of (1.1) for $\alpha\in(-1,0)$.
-More precisely, replacing $f$ with the characteristic function $\operatorname{1}_{K}$ of a convex body $K\subset\mathbb{R}^{n}$ reduces the inequality to the isoperimetric inequality for $I_{\alpha+1}(K)$ with $\alpha\in(-1,0)$.
-Building on this correspondence, it is natural to ask whether the isoperimetric inequalities established for $\alpha\in(0,n)$ and $\alpha>n$ also admit functional counterparts.
-Question.
-Does there exist a functional inequality for a suitable class of functions $f$ whose geometric instance, corresponding to $f=\operatorname{1}_{K}$, recovers the isoperimetric inequalities for chord power integrals with $\alpha>0$?
-Noting that for $\alpha>0$, the quantity $I_{\alpha+1}(K)$ also admits a double integral representation, as shown in [40, Lemma 3.3]:
-$I_{\alpha+1}(K)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{K}\int_{K}\frac{1}{|x-y|^{n-\alpha}}dxdy.$
-This representation suggests a connection with the Hardy-Littlewood-Sobolev inequality (HLS inequality) established by Lieb [30]: for $0<\alpha<n$ and non-negative $f\in L^{p}(\mathbb{R}^{n})$ with $p=2n/(n+\alpha)$,
-$\gamma_{n,\alpha}\|f\|_{\frac{2n}{n+\alpha}}^{2}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{f(x)f(y)}{|x-y|^{n-\alpha}}dxdy.$ (1.6)
-Equality in (1.6) occurs if and only if $f(x)=a(1+\lambda|x-x_{0}|^{2})^{-(n+\alpha)/2}$ for $x\in\mathbb{R}^{n}$ with $a\geq 0,\lambda>0$ and $x_{0}\in\mathbb{R}^{n}$.
-This observation leads to the unexpected conclusion that the HLS inequality is not the functional counterpart of (1.2).
-On further inspection, the essential obstruction appears to be the $2$-homogeneity of the HLS inequality in $f$.
-Motivated by the 1-homogeneous structure arising in the case $\alpha\in(-1,0)$, we turn to a formulation that reflects the intrinsic homogeneity underlying the geometric setting.
-This leads naturally to the following 1-homogeneous analogue of (1.6).
-Theorem 1 (Chord Sobolev inequalities).
-For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
-$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,$
-where the sharp constant $\sigma_{n,\alpha}$ is given by (1.5). Equality holds if and only if $f$ is a constant multiple of the characteristic function of a ball.
-With the double integral appearing in Theorem 1 and the fractional Sobolev norm (1.3), we can naturally extend the notion of the chord power integral to functions $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ by setting
-$I_{\alpha+1}(f)=\frac{-\alpha(\alpha+1)}{2n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha\in(-1,0),$
-and
-$I_{\alpha+1}(f)=\frac{\alpha(\alpha+1)}{n\omega_{n}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,\quad\quad\alpha>n.$
-A more unified formulation can be written as
-$I_{\alpha+1}(f)=\int_{0}^{\infty}I_{\alpha+1}(\{f\geq r\})dr,$
-which follows from the coarea formula when $\alpha\in(-1,0)$ and from the layer-cake representation of $\min\{f(x),f(y)\}$ when $\alpha>n$.
-Here $f$ is typically assumed to be an integrable log-concave function, although the log-concavity assumption is not required in general.
-For instance, when $\alpha\in(-1,0)$, it suffices that $f$ is a non-negative function in the corresponding fractional Sobolev space.
-The inequality in Theorem 1 is a functional counterpart of the classical isoperimetric inequality (1.2).
-It is invariant under translations and rotations, but not under volume-preserving linear transformations.
-Affine inequalities, which remain invariant under translations and volume-preserving linear transformations, have proved to be very powerful, as they often strengthen the corresponding Euclidean inequalities.
-A famous example is the Petty projection inequality for convex bodies in $\mathbb{R}^{n}$.
-It strengthens the Euclidean isoperimetric inequality and also directly implies it.
-See, for example, [14, 45], and see [7, 19, 23, 29, 36, 37, 39, 47, 51] for extensions in the $L^{p}$ Brunn–Minkowski theory and for more general sets.
-Recently, Milman and Yehudayoff [42] established isoperimetric inequalities for affine quermassintegrals.
-These new inequalities are significantly stronger than the isoperimetric inequalities for the classic quermassintegrals, thereby settling a long-standing conjecture of Lutwak [35].
-More recently, Haddad and Ludwig [25] established the following sharp affine fractional Sobolev inequality, which is significantly stronger than the classical inequality (1.4).
-Affine fractional Sobolev inequalities: For $1<\alpha<0$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$, that is, $f$ with finite $\|f\|_{W^{-\alpha,1}(\mathbb{R}^{n})}$,
-$\begin{split}2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha}\Big\|\frac{f(\cdot+tu)-f(\cdot)}{t}\Big\|_{1}dt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}\\ \leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.\end{split}$ (1.7)
-There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid.
-There is equality in the second inequality for radially symmetric functions.
-In the limit as $\alpha\rightarrow-1^{+}$, the inequality (1.7) recovers the well-known affine Sobolev inequality established by Zhang [51], which is an affine version of the classic $L^{1}$ Sobolev inequality.
-For a Sobolev function $f$, Zhang [51] defined the polar projection body $\Pi^{*}f$ by
-$\|\xi\|_{\Pi^{*}f}=\frac{1}{2}\int_{\mathbb{R}^{n}}|\nabla f(x)\cdot\xi|dx,~~\xi\in\mathbb{S}^{n-1},$
-where $\|\cdot\|_{K}$ denotes the gauge function of $K$.
-Wang [47] extended the definition to functions of bounded variation and obtained the generalized Sobolev-Zhang inequality.
-The extended version corresponds to the functional form of the Petty projection inequality.
-Using (1.7), Haddad and Ludwig derived the fractional Petty projection inequalities by taking $f=\operatorname{1}_{K}$.
-Motivated by the same central ideas as in their earlier work [25], Haddad and Ludwig [26, 27] established the affine fractional $L^{p}$ Sobolev inequalities as well as affine HLS inequalities.
-For further results on affine functional inequalities, we refer the reader to [9, 11, 22].
-The following Theorem 2 presents an affine version of our functional isoperimetric inequality for chord power integrals, and thus strengthens Theorem 1.
+Affine inequalities, invariant under translations and volume-preserving linear transformations, often strengthen classical Euclidean inequalities.
+A seminal example is Zhang’s affine Sobolev inequality [48], which significantly improves the classical Sobolev inequality.
+Wang [44] later extended this inequality to functions of bounded variation, yielding a generalized Petty projection inequality (see also [48]), itself a strengthening of the classical isoperimetric inequality.
+Recently, Milman and Yehudayoff [39] established affine isoperimetric inequalities for affine quermassintegrals, settling a long-standing conjecture of Lutwak [33].
+For further developments on affine inequalities, see, for example, [7, 10, 12, 18, 20, 21, 22, 23, 34, 35, 36].
+More recently, Haddad and Ludwig [24, 25, 26] developed a new approach to study affine inequalities.
+They proved the following sharp affine fractional Sobolev inequalities in [24], which are stronger than their Euclidean counterparts by Almgren and Lieb [2] (see also [15]).
+Affine fractional Sobolev inequalities: For $\alpha\in(-1,0)$ and $f\in W^{-\alpha,1}(\mathbb{R}^{n})$,
+$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}$ $r^{\alpha-1}\|f(\cdot)-f(\cdot+ru)\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.1)
+$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy,$
+where $\mathbb{S}^{n-1}$ is the unit sphere in $\mathbb{R}^{n}$ and $\omega_{n}$ denotes the volume of the unit ball in $\mathbb{R}^{n}$. The sharp constant $\sigma_{n,\alpha}$ is given by
+$\sigma_{n,\alpha}=\frac{n2^{\alpha}\pi^{\frac{n-\alpha-1}{2}}\Gamma(\frac{\alpha+1}{2})\Gamma(\frac{n}{2}+1)^{\frac{\alpha}{n}}}{|\alpha|\Gamma(\frac{n+\alpha+1}{2})}.$ (1.2)
+Here $\Gamma$ denotes the gamma function, $\|f\|_{p}:=(\int_{\mathbb{R}^{n}}f(x)^{p}dx)^{1/p}$ denotes the $L^{p}$ norm of $f$ and $W^{-\alpha,1}(\mathbb{R}^{n})$ is the space of integrable functions $f$ such that the right-most side of (1.1) is finite.
+Eliminating the intermediate term in (1.1) gives the fractional Sobolev inequalities, which converge to the classical Sobolev inequality when $\alpha\rightarrow-1^{+}$, as shown by Bourgain, Brezis and Mironescu [8].
+Haddad and Ludwig [24] showed that (1.1) becomes the affine Sobolev inequality in the limit $\alpha\rightarrow-1^{+}$.
+The fractional Sobolev inequalities have been extensively studied over the past decades (see, e.g., [1, 8, 13, 27, 30, 31]).
+In very recent work by the authors [4], the following sharp chord Sobolev inequalities for positive $\alpha$ are established, where, for $E\subset\mathbb{R}^{n}$, the characteristic function $\operatorname{1}_{E}(x)$ takes value $1$ with $x\in E$ and 0 otherwise.
+Chord Sobolev inequalities for $\boldsymbol{\alpha>0}$: For $\alpha\in(0,n)$ and non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
+$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$ (1.3)
+For $\alpha>n$ and non-zero, non-negative $f\in L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$,
+$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy,$ (1.4)
+where $\sigma_{n,\alpha}$ is given by (1.2). In both cases, equality holds if and only if $f$ is a constant multiple of the characteristic function of a ball.
+In [4], fractional Sobolev inequalities, together with (1.3) and (1.4), form a complete series called the chord Sobolev inequalities.
+When $f=\operatorname{1}_{K}$ is the characteristic function of a convex body $K\subset\mathbb{R}^{n}$ (a convex compact set with non-empty interior), these inequalities reduce to the isoperimetric inequalities for chord power integrals; see [4] for details.
+Chord power integrals have been studied extensively in recent years [4, 11, 19, 37, 45, 46].
+One of the main aims of this paper is to prove the following affine chord Sobolev inequalities for $\alpha>0$, which strengthen their Euclidean counterparts (1.3) and (1.4).
+Here $f$ is radially symmetric if its value depends only on $|x|$.
 Theorem 1.
-For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
-$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}\Big)^{\frac{\alpha}{n}}$ (1.8)
+For $\alpha\in(0,n)$ and non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
+$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}$ $\big\|\min\{f(\cdot),f(\cdot+ru)\}\big\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.5)
 $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
@@
 There is equality in the second inequality if $f$ is radially symmetric.
-In Section 4, we introduce an anisotropic version of the right side of (1.8), which gives rise to a new star-shaped set ${\rm L}_{\alpha}f$, defined via its radial function by
-$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,~~u\in\mathbb{S}^{n-1}.$
-With the definition, the first inequality in Theorem 2 can be written as
-$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$ (1.9)
-Since both sides of (1.9) are invariant under translations of $f$ and
-${\rm L}_{\alpha}(f\circ\phi^{-1})=\phi{\rm L}_{\alpha}f$
-for every volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, it follows that inequality (1.9) is indeed affine invariant.
-According to (1.1) and (1.2) the isoperimetic inequality for $I_{\alpha+1}(K)$ reverses when $\alpha>n$.
-Theorem 1 establishes the functional analogue for $\alpha\in(0,n)$.
-However, the case $\alpha>n$ does not admit a straightforward reversal, which will be discussed in detail in Section 6.
-The next result presents the affine inequality we obtain for $\alpha>n$, where ${\rm L}_{\alpha}f$ has been introduced above and $L^{\infty}(\mathbb{R}^{n})$ is the space of essentially bounded measurable functions, equipped with the essential supremum norm $\|\cdot\|_{\infty}$.
-By eliminating the intermediate term, it serves as the functional analogue of (1.1) for $\alpha>n$.
 Theorem 2.
-Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
-$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
+For $\alpha>n$ and non-zero, non-negative $f\in L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$,
+$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}\Big(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}$ $\big\|\min\{f(\cdot),f(\cdot+ru)\}\big\|_{1}dr\Big)^{\frac{n}{\alpha}}du\Big)^{\frac{\alpha}{n}}$ (1.6)
+$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
 There is equality in the first inequality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid.
 There is equality in the second inequality if $f$ is radially symmetric.
-In Section 7 and Section 8, we relate the set ${\rm L}_{\alpha}f$ to radial mean bodies originally introduced by Gardner and Zhang [18].
-For $\alpha>-1$ and a convex body $K\subset\mathbb{R}^{n}$, the radial $\alpha$-th mean body ${\rm R}_{\alpha}K$ is defined via its radial function by
-$\rho_{{\rm R}_{\alpha}K}(u)=\Big(\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx\Big)^{1/\alpha},\quad\quad\alpha\neq 0,$
+The affine fractional Sobolev inequalities (1.1), together with (1.5) and (1.6), are called affine chord Sobolev inequalities.
+To prove Theorem 1 and Theorem 2, we introduce an anisotropic version of the right-most side of (1.5) (see, e.g., [24, 30, 31]), which gives rise to a new star-shaped set ${\rm L}_{\alpha}f$, defined via its radial function by
+$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr,~~u\in\mathbb{S}^{n-1},$
+(see Section 3 for details). The first inequality in Theorem 1 can then be written as
+$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},$ (1.7)
+where $|{\rm L}_{\alpha}f|$ denotes the volume of ${\rm L}_{\alpha}f$. Since both sides of (1.7) are invariant under translations of $f$ and
+${\rm L}_{\alpha}(f\circ\phi^{-1})=\phi{\rm L}_{\alpha}f$
+for every volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, it follows that inequality (1.7) is affine invariant.
+For a convex body $K\subset\mathbb{R}^{n}$, the star-shaped set ${\rm L}_{\alpha}\operatorname{1}_{K}$ is proportional to the radial mean body ${\rm R}_{\alpha}K$ of Gardner and Zhang [18], as will be discussed in Section 5.
+Haddad and Ludwig [24] proved affine isoperimetric inequalities for ${\rm R}_{\alpha}K$ with $\alpha>-1$, stating that
+$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$ (1.8)
+$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n.$
+Equality holds if and only if $K$ is an ellipsoid.
+In [4], the authors extended the notion ${\rm R}_{\alpha}K$ to log-concave functions $f\in L^{1}(\mathbb{R}^{n})$ by
+$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t),~\alpha\neq 0,$ (1.9)
 and
-$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx,\quad\quad\alpha=0.$
-For $\alpha\in(-1,0)$, Haddad and Ludwig [25] showed that the radial mean body ${\rm R}_{\alpha}K$ is proportional to $\Pi^{*,-\alpha}\operatorname{1}_{K}$, where $\Pi^{*,-\alpha}\operatorname{1}_{K}$ denotes the fractional polar projection body arsing from the affine fractional Sobolev inequality (1.7) (see Section 2.6 for details).
-Of particular interest, for $f$ with bounded variation, Haddad and Ludwig [25] proved that
-$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|u\|_{\Pi^{*,-\alpha}f}^{-\alpha}=\|u\|_{\Pi^{*}f}.$
-In Section 7, we show that ${\rm R}_{\alpha}K$ is proportional to ${\rm L}_{\alpha}\operatorname{1}_{K}$ when $\alpha>n$.
-Motivated by this relation, we introduce an explicit definition of the functional counterpart ${\rm R}_{\alpha}f$, based on the fractional polar projection body $\Pi^{*,-\alpha}f$ and the star-shaped set ${\rm L}_{\alpha}f$.
-We note that Haddad and Ludwig [27] also used this notation for radial mean bodies, which in their setting are defined via the fractional $L^{2}$ polar projection body and the star-shaped sets appearing in affine HLS inequalities.
-Here, we use the same notation ${\rm R}_{\alpha}f$ for an $L^{1}$- extension to keep the notation simple, even though the definition in [27] is slightly different.
-In our setting, in particular, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$, we relate ${\rm R}_{\alpha}f$ to the radial mean bodies of its superlevel sets $\{f\geq r\}$.
-More precisely, for $\alpha>-1$,
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),\quad\quad\quad\alpha\neq 0,$
-and
-$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r),\quad\quad\alpha=0,$
-where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
-This unexpected relation naturally leads to the definition of ${\rm R}_{\infty}f$,
-${\rm R}_{\infty}f={\rm R}_{\infty}~{\rm supp}(f)={\rm D}~{\rm supp}(f),$
-where $f\in L^{1}(\mathbb{R}^{n})$ is a log-concave function with compact support, denoted by ${\rm supp}(f)$.
-The notation ${\rm D}K$ is the difference body of $K$, defined by ${\rm D}K=\{x-y:x,y\in K\}$.
-This observation yields the following monotonicity property:
+$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)d\mu_{f}(t),$
+where $d\mu_{f}(t)=\frac{|\{f\geq t\}|}{\|f\|_{1}}dt$ is a probability measure on $\mathbb{R}$.
+We note that Haddad and Ludwig [26] also use the notation ${\rm R}_{\alpha}f$ for the $\alpha$-th radial mean body, but with a different definition.
+The present definition can be viewed as an $L^{1}$ extension to their definitions.
+It was shown in [4] that ${\rm R}_{\alpha}f$ is proportional to ${\rm L}_{\alpha}f$ for $\alpha>0$, while for $\alpha\in(-1,0)$ it is proportional to the fractional polar projection body $\Pi^{*,-\alpha}f$ arising from the affine fractional Sobolev inequalities (1.1).
+Consequently, the affine chord Sobolev inequalities yield the following affine isoperimetric inequalities for ${\rm R}_{\alpha}f$: for a log-concave function $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$,
+$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$ (1.10)
+$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{1}\|f\|_{\infty}^{-1}}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n,$
+with equality if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid.
+In particular, when $f=\operatorname{1}_{K}$ is the characteristic function of a convex body, (1.10) reduces to the geometric inequality (1.8).
+Of particular interest, for log-concave $f\in L^{1}(\mathbb{R}^{n})$, Haddad and Ludwig [24] proved
+$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1},$
+where $\Pi^{*}f$ is the polar projection body of $f$ introduced by Zhang [48] and extended by Wang [44] (see Section 2.5).
+For the other endpoint $\alpha\to\infty$, the monotonicity of $L^{p}$ norms with respect to $p$, together with (1.9), leads to the definition of ${\rm R}_{\infty}f$:
+${\rm R}_{\infty}f={\rm R}_{\infty}\,{\rm supp}(f)={\rm D}\,{\rm supp}(f),$
+where ${\rm D}K:=\{x-y:x,y\in K\}$ denotes the difference body, and we set ${\rm R}_{\infty}f=\mathbb{R}^{n}$ if ${\rm supp}(f)=\mathbb{R}^{n}$.
+In Section 5, we prove the following monotonicity for log-concave $f\in L^{1}(\mathbb{R}^{n})$:
 ${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f$
 and
-$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
-for $1<\alpha<\beta<\infty$. When $f=\operatorname{1}_{K}$, these relations reduce to the geometric case of Gardner and Zhang [18].
-Gardner and Zhang also established the reversed inclusion in [18, Theorem 5.5], which requires a more involved argument. For a convex body $K\subset\mathbb{R}^{n}$, they proved
+$\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$ (1.11)
+for $1<\alpha<\beta<\infty$. When $f=\operatorname{1}_{K}$, these reduce to the geometric case of Gardner and Zhang [18].
+Gardner and Zhang [18] also established the reverse inclusion for radial mean bodies, which is more involved. For a convex body $K\subset\mathbb{R}^{n}$, they proved
 ${\rm D}K\subset c_{n,\beta}{\rm R}_{\beta}K\subset c_{n,\alpha}{\rm R}_{\alpha}K\subset n|K|\Pi^{*}K,$ (1.12)
-for $1<\alpha<\beta<\infty$.
-In each inclusion, equality holds if and only if $K$ is a simplex, denoted by $\triangle$.
-Here $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$, where $B(\alpha+1,n)$ denotes the Beta function.
-In (1.10), when $\beta=n$, the left side inclusion implies the Rogers-Shephard inequality [44], which is
-$|{\rm D}K|\leq\binom{2n}{n}|K|.$
-When $\alpha=n$, the right side inclusion yields Zhang’s inequality [50] (see also Ball [5]):
-$n^{-n}\binom{2n}{n}=|\triangle|^{n-1}|\Pi^{*}\triangle|\leq|K|^{n-1}|\Pi^{*}K|,$
-which is a reversed form of Petty projection inequality.
-To extend these geometric inequalities to their functional analogues, there are two natural ways to associate a convex body with a log-concave function:
+for $1<\alpha<\beta<\infty$, where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$ and $B(\cdot,\cdot)$ is the Beta function.
+Equality holds in each inclusion if and only if $K$ is a simplex.
+When $\beta=n$, the left-most inclusion gives the Rogers–Shephard inequality [41], while for $\alpha=n$, the right-most inclusion yields Zhang’s inequality [47] (see also Ball [5]).
+Building on the functional radial mean body ${\rm R}_{\alpha}f$ and (1.11), we aim to extend the reverse inclusion (1.12) from sets to functions.
+There are two natural ways to associate a convex body with a log-concave function:
 $K\mapsto\operatorname{1}_{K}\quad\quad\text{and}\quad\quad K\mapsto e^{-\|\cdot\|_{K}}.$
-Using the second correspondence, Alonso-Gutiérrez, Bernués and González Merino [2] extended Zhang’s inequality to log-concave functions, showing that for a log-concave $f\in L^{1}(\mathbb{R}^{n})$,
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq 2^{n}n!\|f\|_{1}^{n+1}|\Pi^{*}f|.$ (1.11)
-If $f(0)=\|f\|_{\infty}$, equality holds if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$.
-For further results on Zhang’s inequality and on radial mean bodies, we refer the reader to [3, 4, 20, 21, 24, 28].
-In Section 9, we extend (1.10) to both log-concave and $s$-concave functions, which are stated below as the following two theorems.
+Using the second correspondence, Langharst, Marín Sola and Ulivelli [28] showed that for log-concave $f\in L^{1}(\mathbb{R}^{n})$,
+$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f.$ (1.13)
+Equality holds in each inclusion if and only if $f(x)=\|f\|_{\infty}e^{-\|x-x_{0}\|_{\triangle}}$, where $x_{0}$ is a maximizer of $f$ and $\triangle$ is a simplex containing the origin and $x_{0}$.
+When $\alpha=n$, the right-most inclusion recovers Zhang’s inequality for log-concave functions, as established by Alonso-Gutiérrez, Bernués and González Merino [3].
+However, in the limit $\beta\rightarrow\infty$, the body ${\rm R}_{\infty}f$ no longer appears in (1.13).
+This indicates a limitation of the second correspondence in capturing the endpoint behavior of radial mean bodies.
+On the other hand, the monotonicity property (1.11) naturally extends from convex bodies to functions via the embedding $K\mapsto\operatorname{1}_{K}$.
+These considerations motivate us to adopt the first correspondence and lead to the following question.
+Question. For a non-zero, log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, does
+${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f$
+hold for $1<\alpha<\beta<\infty$? Here $c_{n,\alpha}$ is the constant in (1.12).
+Unfortunately, this inclusion fails for general log-concave functions, as we show by a counterexample in Section 6. This failure leads us to consider $s$-concave functions with $s>0$.
+Let $f_{+}(x):=\max\{f(x),0\}$ denote its positive part. The following theorem extends (1.12) to $s$-concave functions, with fully characterized equality cases.
 Theorem 3.
-Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
-$\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$
-for $1<\alpha<\beta<\infty$. In each inclusion, there is equality if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$.
-When $\alpha=n$, the inclusion
-$\frac{1}{\Gamma(n+1)^{1/n}}{\rm R}_{n}f\subset 2\|f\|_{1}|\Pi^{*}f|,$
-recovers (1.11).
-The inclusion relations in Theorem 4 can be deduced from the results of Alonso-Gutiérrez, Bernués and González Merino [2, Lemma 3.2, Remark 5], but the equality cases were not addressed there.
-In Section 9, we employ a different approach that allows us to establish Theorem 4 together with a complete characterization of equality.
-In proving Theorem 4, we first establish an inclusion relation between ${\rm R}_{\alpha}f$ and ${\rm R}_{\beta}f$, and then consider the limits $\alpha\to-1$ and $\beta\to\infty$ to obtain inclusions between ${\rm R}_{\infty}f$ and $\Pi^{*}f$.
-However, in this limiting process $\beta\to\infty$ degenerates to the singular set $\{o\}$, so that ${\rm R}_{\infty}f$ does not appear in the general case.
-This motivates us to consider the embedding $K\mapsto\operatorname{1}_{K}$, leading to the following natural question:
-Open problem. For a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, does
-${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f$
-hold? Here $c_{n,\alpha}$ is the constant in (1.10).
-Unfortunately, this inclusion fails for general log-concave integrable functions with compact support, as we show by a counterexample in Section 9.
-This failure leads us naturally to consider $s$-concave functions.
-Our main result concerns $s$-concave functions $f\in L^{1}(\mathbb{R}^{n})$, where we denote the positive part of $f$ by $f_{+}(x):=\max\{f(x),0\}$.
-The equality cases in this setting are fully characterized as well.
-Theorem 5.
-Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$,
-${\rm R}_{\infty}f\subset c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f\subset(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$ (1.12)
-where $c_{n,\alpha}(s)=(nB(\alpha+1,n+\frac{1}{s}))^{-1/\alpha}$.
-In each inclusion, there is equality if and only if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
-Theorems 4 and 5 share the same general structure and rely on similar methods of proof.
-As noted above, in Theorem 4, $\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f$ degenerates to the singular set $\{o\}$ as $\beta\to\infty$, so that the set ${\rm R}_{\infty}f$ does not appear.
-In contrast, Theorem 5 avoids this degeneration, and in the limit $s\to\infty$, the inclusion (1.12) recovers (1.10).
-2. Preliminary
-We collect some basic notation and results on $L^{p}$ spaces, Schwarz symmetrals, star-shaped sets, log-concave functions, radial mean bodies and fractional polar projection bodies.
-The books of Gardner [14], Schneider [45], and Schneider and Weil [46] are good general references.
+Let $s>0$ and let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, the following chain of inclusions holds:
+${\rm R}_{\infty}f\subset c_{n,\beta}(s)\,{\rm R}_{\beta}f\subset c_{n,\alpha}(s)\,{\rm R}_{\alpha}f\subset\Big(n+\frac{1}{s}\Big)\|f\|_{1}\,\Pi^{*}f,$ (1.14)
+where $c_{n,\alpha}(s)=\bigl((n+\tfrac{1}{s})\,B(\alpha+1,n+\tfrac{1}{s})\bigr)^{-1/\alpha}$.
+In each inclusion, there is equality if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
+Letting $\beta=n$ and $\alpha=n$ in (1.14) yields the Rogers-Shephard inequality and Zhang inequality, respectively, for $s$-concave functions; see Section 6 for detailed statement.
+The optimizer in (1.14) admits a simple geometric interpretation.
+Let $\triangle$ be a simplex in $\mathbb{R}^{n}$ with vertices $v_{0},\ldots,v_{n}$, and identify $\mathbb{R}^{n}$ with the hyperplane $\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n+1}$.
+For $f^{s}(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}\quad\text{with }x_{0}\in\triangle$, its hypograph is given by the set
+$\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm supp}(f^{s}),~0\leq l\leq f^{s}(x)\}$
+and coincides exactly with the simplex in $\mathbb{R}^{n+1}$ having vertices $v_{0},\ldots,v_{n},(x_{0},a)$.
+The equality characterization of (1.14) relies on the equality case of the Borell–Brascamp–Lieb inequality.
+Its highly rigid structure accounts for the difficulty in the study of the equality cases.
+2. Preliminaries
+In this section, we collect basic notation and results on function spaces and Schwarz symmetrals, together with some basic geometric facts.
+Useful general references include the books by Gardner [16], Schneider [42], and Schneider and Weil [43].
 2.1. $L^{p}$ spaces
 Let p>1p>-1 and $(X,\mathcal{B},\mu)$ be a measure space.
-In this paper, we mainly consider the case $X=\mathbb{R}^{n}$ or $X\subset\mathbb{R}^{n}$ with $\mu$ being the $n$-dimensional Lebesgue measure.
+In this work, we primarily focus on the cases $X=\mathbb{R}^{n}$ or $X\subset\mathbb{R}^{n}$ with $\mu$ being the $n$-dimensional Lebesgue measure.
 For $p\neq 0$ and a measurable function $f:X\rightarrow\mathbb{R}$, let
-$\|f\|_{p}=\Big(\int_{X}|f(x)|^{p}d\mu(x)\Big)^{1/p}.$
-In particular, if $p\in(-1,0)$ and $\mu(\{x\in X:f(x)=0\})>0$, we set $\|f\|_{p}=+\infty$. The space $L^{p}(X)$ is then defined as
+$\|f\|_{p}=\Big(\int_{X}|f(x)|^{p}d\mu(x)\Big)^{1/p},$
+and
+$\|f\|_{\infty}=\inf\{C>0:|f(x)|\leq C~~\text{almost everywhere on}~X\}.$
+In particular, for $p\in(-1,0)$, we set $\|f\|_{p}=\infty$ if $\mu(\{x\in X:f(x)=0\})>0$. The associated space $L^{p}(X)$ is then defined as
 $L^{p}(X)=\{f:X\rightarrow\mathbb{R}:f~\text{is measurable},~\|f\|_{p}<\infty\},$
-where the $f=g$ in $L^{p}(X)$ means that they are equal almost everywhere.
-When $\mu(x)=1$, the $L^{p}$-norm is monotone in $p$, that is,
+where $f=g$ in $L^{p}(X)$ means that they are equal almost everywhere.
+When $\mu(X)=1$, the $L^{p}$ norm is monotone in $p$: for 1<p<q<0-1<p<q<0 or $0<p<q<\infty$,
 $\|f\|_{p}\leq\|f\|_{q},$
-for 1<p<q<0-1<p<q<0 or $0<p<q<\infty$. This follows from Jensen’s inequality. In fact, the monotonicity also holds for $p<0<q$. To see this, under the assumption $\mu(X)=1$, we first note the limit
+as a consequence of Jensen’s inequality. In fact, this monotonicity also extends for $p<0<q$. To see this, under the assumption $\mu(X)=1$, we first note the limit
 $\lim_{p\rightarrow 0}\log\|f\|_{p}=\int_{X}\log|f(x)|d\mu(x),$
@@
 2.2. Symmetrization
-For $E\subset\mathbb{R}^{n}$, the characteristic function $\operatorname{1}_{E}$ is defined by $\operatorname{1}_{E}(x)=1$ for $x\in E$, and $\operatorname{1}_{E}(x)=0$ for $x\notin E$.
-If $E$ is a Borel set of finite measure, the Schwarz symmetral of $E$, denoted by $E^{\star}$, is the closed, centered ball with the same volume as $E$.
-Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a non-negative measurable function.
-The superlevel set of $f$ is defined by $\{f\geq t\}=\{x\in X:f(x)\geq t\}$ for $t>0$.
+For a Borel set $E\subset\mathbb{R}^{n}$ with finite measure, the Schwarz symmetral of $E$, denoted by $E^{\star}$, is the closed, centered ball having the same volume as $E$.
+Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function.
+For $t>0$, the superlevel set of $f$ is defined by $\{f\geq t\}=\{x\in\mathbb{R}^{n}:f(x)\geq t\}$.
 We say $f$ is non-zero if $\{f\neq 0\}$ has positive measure.
-For $f$ with superlevel sets of finite measure, by the layer cake formula,
+If the superlevel sets of $f$ have finite measure, the layer cake formula gives
 $f(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}}(x)dt$
-for almost every $x\in\mathbb{R}^{n}$.
-The Schwarz symmetral of $f$, denoted by $f^{\star}$, is defined as
+for almost every $x\in\mathbb{R}^{n}$. The Schwarz symmetral of $f$, denoted by $f^{\star}$, is defined as
 $f^{\star}(x)=\int_{0}^{\infty}\operatorname{1}_{\{f\geq t\}^{\star}}(x)dt$
-for $x\in\mathbb{R}^{n}$. Hence $f^{\star}$ is radially symmetric and the superlevel set $\{f^{\star}\geq t\}$ has the same volume as $\{f\geq t\}$ for every $t>0$.
-One main technique in our proofs is the Riesz rearrangement inequality (see, for example [31, Theorem 3.7]) and the characterization of equality case due to Burchard [8].
-Theorem 2.1 (Riesz’s rearrangement inequality.).
+for $x\in\mathbb{R}^{n}$, so that $f^{\star}$ is radially symmetric and the superlevel sets satisfy
+$|\{f^{\star}\geq t\}|=|\{f\geq t\}|$
+for all $t>0$.
+A central tool in our proofs is the Riesz rearrangement inequality (see, for example [29, Theorem 3.7]) together with the characterization of equality by Burchard [9].
+Theorem 2.1 (Riesz’s rearrangement inequality).
 For measurable $f,g,k:\mathbb{R}^{n}\rightarrow[0,\infty)$ with superlevel sets of finite measure.
@@
 2.3. Star-shaped sets and dual mixed volumes
-We say $K\subset\mathbb{R}^{n}$ is a star-shaped set (with respect to the origin) if for any $x\in K$ the line segment $[0,x]$ is contained in $K$.
-The gauge function $\|\cdot\|_{K}:\mathbb{R}^{n}\rightarrow[0,+\infty]$ of a star-shaped set $K$ is defined by
-$\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\},$
+A set $K\subset\mathbb{R}^{n}$ is called star-shaped (with respect to the origin) if for every $x\in K$, the line segment $[o,x]$ is contained in $K$.
+For a star-shaped set $K$ containing the origin in its interior, the gauge function $\|\cdot\|_{K}:\mathbb{R}^{n}\rightarrow[0,\infty]$ is defined by
+$\|x\|_{K}=\inf\{\lambda>0:x\in\lambda K\}.$ (2.1)
+If $K$ contains the origin, possibly on its boundary, then for any point $x$, the gauge function $\|x\|_{K}$ is defined by (2.1) if there exists $r>0$ such that $rx\in K$; otherwise, $\|x\|_{K}=\infty$.
 […]
-$\rho_{K}(x)=\|x\|_{K}^{-1}=\sup\{\lambda\geq 0:\lambda x\in K\}.$
-In particular, when $K$ is the unit ball, $\|\cdot\|_{K}$ is the Euclidean norm $|\cdot|$.
-For a star-shaped set $K$ with measurable radial function, the $n$-dimensional volume, or the $n$-dimensional Lebesgue measure of $K$ is given by
+$\rho_{K}(x)=\|x\|_{K}^{-1}=\sup\{r\geq 0:rx\in K\}.$
+In particular, if $K$ is the unit Euclidean ball, $\|\cdot\|_{K}$ coincides with the Euclidean norm $|\cdot|$.
+For a star-shaped set $K$ with measurable radial function, the volume, or the $n$-dimensional Lebesgue measure of $K$ is given by
 $|K|=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n}du,$
-where $du$ is the spherical Lebesgue measure. The dual mixed volume for star-shaped sets $K$ and $L$ by Lutwak [34] is given by
+where $du$ denotes the spherical Lebesgue measure. Following Lutwak [32], the dual mixed volume for star-shaped sets $K$ and $L$ is defined by
 $\tilde{V}_{\alpha}(K,L)=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{L}(u)^{\alpha}du,$
 […]
-For $\alpha\in(0,n)$, and star-shaped sets $K$ and $L$ of finite volume, the dual mixed volume inequality states that
+$\tilde{V}_{\log}(K,L)=\lim_{\alpha\rightarrow 0}\log\Big(\frac{\tilde{V}_{\alpha}(K,L)}{|K|}\Big)^{1/\alpha}=\frac{1}{n|K|}\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n}\log\Big(\frac{\rho_{K}(u)}{\rho_{L}(u)}\Big)du.$
+If $\alpha\in(0,n)$ and $K,L$ are star-shaped sets with finite volume, the dual mixed volume inequality states that
 $\tilde{V}_{\alpha}(K,L)\leq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.2)
-where equality holds if and only if $K$ and $L$ are dilates, that is, $\rho_{K}(u)=c\rho_{L}(u)$ for some $c\geq 0$ and almost all $u\in\mathbb{S}^{n-1}$.
-For $\alpha<0$ and $\alpha>n$, the dual mixed volume inequality states that
+with equality holds if and only if $K$ and $L$ are dilates, i.e., $\rho_{K}(u)=c\rho_{L}(u)$ for some $c\geq 0$ and almost all $u\in\mathbb{S}^{n-1}$.
+For $\alpha<0$ or $\alpha>n$, the dual mixed volume inequality states that
 $\tilde{V}_{\alpha}(K,L)\geq|K|^{\frac{n-\alpha}{n}}|L|^{\frac{\alpha}{n}},$ (2.3)
 where equality holds if and only if $K$ and $L$ are dilates.
-2.4. Log-concave functions
+2.4. Log-concave and $s$-concave functions
 A function $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is log-concave if $\log f$ is a concave function on $\mathbb{R}^{n}$, with values in $[-\infty,\infty)$.
@@
 Lemma 2.3.
-[2, Lemma 2.1] Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a log-concave and integrable function. Then the function $g:\mathbb{R}^{n}\rightarrow[0,\infty)$ defined by
+[3, Lemma 2.1] Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a log-concave and integrable function. Then the function $g:\mathbb{R}^{n}\rightarrow[0,\infty)$ defined by
 $g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy$
@@
 We say $f$ is $(\lambda_{0},s)$-concave, if $f((1-\lambda_{0})x+\lambda_{0}y)\geq\mathcal{M}_{\lambda_{0}}^{s}(f(x),f(y))$ for fixed $\lambda_{0}\in(0,1)$.
-The following theorem from Balogh and Kristály [6] gives the equality characterization for Borell-Brascamp-Lieb inequality.
+The following theorem of Balogh and Kristály [6] gives the equality characterization of the Borell-Brascamp-Lieb inequality.
 Theorem 2.4.
-[6, Theorem 3.1] Let $\lambda\in(0,1)$ and $s>-\frac{1}{n}$. Let $f,g,h:\mathbb{R}^{n}\rightarrow[0,\infty)$ be three non-zero, compactly supported integrable functions satisfying
+[6, Theorem 3.1] Let $\lambda\in(0,1)$ and $s>-\frac{1}{n}$. Let $F,G,H:\mathbb{R}^{n}\rightarrow[0,\infty)$ be non-zero, compactly supported integrable functions satisfying
 $H((1-\lambda)x+\lambda y)\geq{\mathcal{M}}_{\lambda}^{s}(F(x),G(y)).$
 Then there is equality in
-$\int_{\mathbb{R}^{n}}h\geq{\mathcal{M}}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}f,\int_{\mathbb{R}^{n}}g\Big),$
+$\int_{\mathbb{R}^{n}}H\geq{\mathcal{M}}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big),$ (2.4)
 if and only if there is $x_{0}\in\mathbb{R}^{n}$, an $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ and $c_{0}=(\frac{|{\rm supp}(G)|}{|{\rm supp}(F)|})^{1/n}$ such that up to null measure sets,
@@
 $\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{0}x+x_{0})=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda x_{0})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$.
-(2.3)
-2.5. Convex bodies and radial mean bodies
-A convex body in $\mathbb{R}^{n}$ is a convex, compact set with non-empty interior.
-Let $K\subset\mathbb{R}^{n}$ be a convex body.
-For $\alpha>-1$ and $\alpha\neq 0$, Gardner and Zhang [18] defined the radial $\alpha$-th mean body of $K$ via its radial function by
-$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{1}{|K|}\int_{K}\rho_{K-x}(u)^{\alpha}dx.$
-For $\alpha=0$,
-$\log\rho_{{\rm R}_{0}f}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx.$
-For $x=x^{\prime}+tu$ with $x^{\prime}\in K|u^{\perp}$, we have
-$\int_{K}\rho_{K-x}(u)^{\alpha}dx=\int_{K|u^{\perp}}\int_{t_{1}}^{t_{2}}(t_{2}-t)^{\alpha}dtdx^{\prime},$
-where $t_{1}=\min\{t:x^{\prime}+tu\in K\}$ and $t_{2}=\max\{t:x^{\prime}+tu\in K\}$. Since $t_{2}-t_{1}=|K\cap(x^{\prime}+\mathbb{R}u)|$, we then obtain
-$\int_{K}\rho_{K-x}(u)^{\alpha}du=\frac{1}{\alpha+1}\int_{K|u^{\perp}}|K\cap(x^{\prime}+\mathbb{R}u)|^{\alpha+1}dx^{\prime}.$
 2.5. Fractional polar projection bodies
-For $\alpha\in(-1,0)$ and measurable $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$,
-$\rho_{\Pi^{*,-\alpha}f}(u)^{-\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$ (2.4)
-Then the affine fractional Sobolev inequality (1.7) can be written as
-$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy.$
-Following the calculation in [27], we obtain that
-$\rho_{\Pi^{*,-\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}|K\triangle(K+tu)|dt$ (2.5)
-$=\int_{0}^{\infty}t^{\alpha-1}\int_{K|u^{\perp}}2\min\{|K\cap(y+\mathbb{R}u),t|\}dydt$
-$=2\int_{K|u^{\perp}}\int_{0}^{\infty}t^{\alpha-1}\min\{|K\cap(y+\mathbb{R}u)|,t\}dtdy$
-$=2\int_{K|u^{\perp}}\Big(\int_{0}^{|K\cap(y+\mathbb{R}u)|}t^{\alpha}dt+\int_{|K\cap(y+\mathbb{R}u)|}^{\infty}|K\cap(y+\mathbb{R}u)|t^{\alpha-1}dt\Big)dy$
-$=-\frac{2}{\alpha(\alpha+1)}\int_{K|u^{\perp}}|K\cap(y+\mathbb{R}u)|^{\alpha+1}dy,$
-where $K\triangle L$ is the symmetric difference of $K,L\subset\mathbb{R}^{n}$. Hence,
-$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K,$ (2.6)
-for $\alpha\in(-1,0)$.
-3. Proof of Theorem 1
-In this section, we focus on the proof of Theorem 1.
-Before presenting the proof, we first establish the following lemma, which is based on the Minkowski inequality (see, for example, [31, Theorem 2.4]).
-Let $F$ be a non-negative, measurable function on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and $p\geq 1$.
-The Minkowski inequality states that
-$\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)^{p}dx\Big)^{1/p}dy\geq\Big(\int_{\mathbb{R}^{n}}\Big(\int_{\mathbb{R}^{n}}F(x,y)dy\Big)^{p}dx\Big)^{1/p}.$ (3.1)
-Equality and finiteness in (3.1) for $p\in(1,\infty)$ imply the existence of measurable functions $\varphi,\psi:\mathbb{R}^{n}\rightarrow[0,\infty)$ such that $F(x,y)=\varphi(x)\psi(y)$ almost everywhere.
-Lemma 3.1.
-For $\alpha\in(0,n)$ and non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
-$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
-Equality holds if and only if $f=c\operatorname{1}_{E}$ for some measurable $E$ of finite measure and some constant $c>0$.
-Proof.
-By the Minkowski inequality (3.1),
-$\int_{\mathbb{R}^{n}}f(x)^{\frac{n}{n+\alpha}}dx$ $=\int_{\mathbb{R}^{n}}\Big(\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}dx$ (3.2)
-$\geq\Big(\int_{0}^{\infty}\Big(\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx\Big)^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}}$
-$=\Big(\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\Big)^{\frac{n}{n+\alpha}},$
-that is
-$\|f\|_{\frac{n}{n+\alpha}}\geq\int_{0}^{\infty}|\{f>t\}|^{\frac{n+\alpha}{n}}dt.$
-It is clear that equality is attained when $f=c\operatorname{1}_{E}$. Conversely, if equality holds in (3.2), there are non-negative measurable functions $\varphi$ and $\psi$ such that
-$\operatorname{1}_{\{f\geq r\}}(x)=\varphi(x)\psi(t).$
-This implies that
-$\{(x,r)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)\geq r\}=\{x\in\mathbb{R}^{n}:\varphi(x)>0\}\times\{r\in\mathbb{R}:\psi(r)>0\}.$
-Therefore, for every fixed $r\in\mathbb{R}$ such that $r>0$ and $r\leq f(x)$ for some $x$, the superlevel set $\{f\geq r\}$ is a set independent of $r$, which implies that $f=c\operatorname{1}_{E}$ for some measurable set $E$ of finite measure and some constant $c>0$.
-∎
-We are now ready to prove Theorem 1.
-Proof of Theorem 1.
-Note that
-$\min\{f(x),f(y)\}=\int_{0}^{\infty}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)dr.$
-By Fubini’s theorem and Riesz rearrangement inequality, we have
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdydr$
-$\leq\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{\{f\geq r\}^{\star}}(y)}{|x-y|^{n-\alpha}}dxdydr$
-$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr,$
-where the last equation is from the definition of chord power integrals and $|E^{\star}|=|E|$ for any $E\subset\mathbb{R}^{n}$ with finite measure. The best constant $\sigma_{n,\alpha}$ is given by
-$\sigma_{n,\alpha}=\frac{nI_{\alpha+1}(B^{n})}{\alpha(\alpha+1)\omega_{n}^{\frac{\alpha}{n}}}.$
-Then Lemma 3.1 implies that
-$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
-When the equality holds, it follows from Lemma 3.1 that $f$ is a characteristic function.
-Then by the isoperimetric inequality of chord integrals, we have $f=c\operatorname{1}_{B}$, up to a set of measure zero.
-∎
+For $\alpha\in(-1,0)$ and measurable $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the fractional polar projection body $\Pi^{*,-\alpha}f$ is defined by
+$\rho_{\Pi^{*,-\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dxdr.$
+The affine fractional Sobolev inequalities (1.1) state that
+$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{|f(x)-f(y)|}{|x-y|^{n-\alpha}}dxdy$
+for $f\in W^{-\alpha,1}(\mathbb{R}^{n})$
+Let $\mathcal{B}(\mathbb{R}^{n})$ denotes the class of Borel set in $\mathbb{R}^{n}$.
+Let $C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$ denote the set of smooth vector fields with compact support.
+Moreover, ${\rm div}\,\nu$ denotes the divergence of $\nu\in C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$.
+We say that $f\in L^{1}(\mathbb{R}^{n})$ is a function of bounded variation on $\mathbb{R}^{n}$ if there is a finite vector-valued Radon measure $\mathcal{D}f:\mathcal{B}(\mathbb{R}^{n})\rightarrow\mathbb{R}^{n}$ such that
+$\int_{\mathbb{R}^{n}}\nu(x)\cdot n_{f}(x)d|\mathcal{D}f|(x)=-\int_{\mathbb{R}^{n}}f(x){\rm div}\,\nu(x)dx$
+for every $\nu\in C_{c}^{\infty}(\mathbb{R}^{n};\mathbb{R}^{n})$, where $|\mathcal{D}f|:\mathcal{B}(\mathbb{R}^{n})\rightarrow[0,\infty)$ denotes the variation measure of $\mathcal{D}f$ and $n_{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the Radon-Nikodym derivative of $\mathcal{D}f$ with respect to $|\mathcal{D}f|$.
+For $f\in L^{1}(\mathbb{R}^{n})$ with bounded variation, the polar projection body $\Pi^{*}f$ is defined by
+$\|u\|_{\Pi^{*}f}=\frac{1}{2}\int_{\mathbb{R}^{n}}|n_{f}(x)\cdot u|\,d|\mathcal{D}f|(x).$
+For more information about $\Pi^{*}f$, we refer the reader to [44].
 3. The star-shaped set ${\rm L}_{\alpha}f$
-Building on the approach of Haddad and Ludwig [25, 26, 27], we introduce the following anisotropic version of $I_{\alpha+1}(f)$.
-Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function.
-We consider
+Building on the approach of Haddad and Ludwig [24, 25, 26], we introduce the following anisotropic integral. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. We consider
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy,$
@@
 Here, we first recall the definition of ${\rm L}_{\alpha}f$:
-$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt.$
+$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr.$
 Using Fubini’s theorem and polar coordinate, we obtain that
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy$ $=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(y+z),f(y)\}}{\|z\|_{K}^{n-\alpha}}dzdy$
-$=\int_{\mathbb{R}^{n}}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\rho_{K}(tu)^{n-\alpha}t^{n-1}\min\{f(y),f(y+tu)\}dtdudy$
-$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(y),f(y+tu)\}dydtdu$
+$=\int_{\mathbb{R}^{n}}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\rho_{K}(ru)^{n-\alpha}r^{n-1}\min\{f(y),f(y+ru)\}drdudy$
+$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(y),f(y+ru)\}dydrdu$
 $=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}du.$
 That is,
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f),$ (4.1)
-for measurable $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ and star-shaped set $K\subset\mathbb{R}^{n}$ with measurable radial function. In particular,
+$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f).$ (3.1)
+In particular,
 $|{\rm L}_{n}f|$ $=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{{\rm L}_{n}f}(u)^{n}du$
-$=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}t^{n-1}\min\{f(x),f(x+tu)\}dxdtdu$
+$=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}r^{n-1}\min\{f(x),f(x+ru)\}dxdrdu$
 $=\frac{1}{n}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy.$
@@
 For $\alpha\in(0,n)$, the dual mixed volume inequality (2.2) together with (3.1) implies
-$\sup\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
+$\sup\Big\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\Big\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
 For $\alpha>n$, the dual mixed volume inequality (2.3) together with (3.1) yields
-$\inf\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
+$\inf\Big\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\Big\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
 Moreover, in either regime a suitable dilate of ${\rm L}_{\alpha}f$ attains the supremum (respectively, infimum), provided $|{\rm L}_{\alpha}f|$ is finite.
@@
 Proposition 3.1.
-If $f:\mathbb{R}^{n}\to[0,\infty)$ is log-concave and in $L^{1}(\mathbb{R}^{n})$, then ${\rm L}_{\alpha}f$ is a convex set for every $\alpha>0$.
+Let $\alpha>0$. If $f:\mathbb{R}^{n}\to[0,\infty)$ is a log-concave, integrable function with compact support, then ${\rm L}_{\alpha}f$ is a convex body.
 Proof.
 For $u\in\mathbb{S}^{n-1}$, we define
-$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx.$
+$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
 Since $f$ is log-concave, Lemma 2.3 implies that $g$ is also log-concave.
 Note that
-$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}g(tu)dt.$
-Applying a result of Ball [5, Theorem 5], we obtain that
-$u\mapsto\left(\int_{0}^{\infty}t^{\alpha-1}g(tu)dt\right)^{-\frac{1}{\alpha}}$
-defines a norm on $\mathbb{R}^{n}$, which concludes the proof. ∎
-5. Affine isoperimetric inequalities for $\alpha\in(0,n)$
-In this section, we aim to establish an affine version of Theorem 1.
-To this end, we first prove the following crucial lemma, which uses the Riesz rearrangement inequality.
-We say that $A\subset\mathbb{R}^{n}$ is homothetic to $B\subset\mathbb{R}^{n}$ if there exist $x\in\mathbb{R}^{n}$ and $a>0$ such that $A=x+aB$.
+$\rho_{{\rm L}_{\alpha}f}(u)=\Big(\int_{0}^{\infty}r^{\alpha-1}g(ru)dr\Big)^{1/\alpha}.$
+It follows from [18, Corollary 4.2] that ${\rm L}_{\alpha}f$} is a convex body. ∎
+4. Affine chord Sobolev inequalities
+In this section, we establish the affine chord Sobolev inequalities stated in Theorem 1 and Theorem 2. To this end, we first prove a crucial lemma based on the Riesz rearrangement inequality.
+Recall that a set $A\subset\mathbb{R}^{n}$ is homothetic to $B\subset\mathbb{R}^{n}$ if there exist $x\in\mathbb{R}^{n}$ and $a>0$ such that $A=x+aB$.
+Moreover, $A$ is said to be equivalent to $B$ if their symmetric difference $A\triangle B$ is a null set.
+The following lemma provides a general form of the Riesz rearrangement inequality; see also [10, 24, 25, 26].
 Lemma 4.1.
@@
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
-The equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost all $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to a set of measure zero.
+Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost all $t>0$, the level set $\{f\geq t\}$ has measure zero or is homothetic to $D$ up to a set of measure zero.
 Proof.
 […]
-$\|z\|_{K}^{-q}=\int_{0}^{\infty}\operatorname{1}_{t^{-1/q}K}(z)dt.$
+$\|z\|_{K}^{-q}=\int_{0}^{\infty}\operatorname{1}_{r^{-1/q}K}(z)dr.$
 Using the Fubini’s theorem, we have
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{\|x-y\|_{K}^{q}}dxdydr$
-$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydtdr.$
+$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{\{f\geq t\}}(y)}{\|x-y\|_{K}^{q}}dxdydt$
+$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdydrdt.$
 By the Riesz rearrangement inequality, Theorem 2.2, we have
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
-$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
-$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq r\}}(y)dxdy,$
+$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{-1/q}K}$ $(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdy$ (4.1)
+$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq t\}}(x)\operatorname{1}_{r^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq t\}}(y)dxdy,$
 which implies that
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{\|x-y\|_{K^{\star}}^{q}}dxdy.$
-If the equality holds, we have
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
-$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
-for almost all $(r,t)\in(0,\infty)^{2}$. For sufficiently large $t>0$, the assumptions of Theorem 2.2 are fulfilled. Therefore, up to sets of measure zero, we have
-$\{f\geq r\}=x+aD,~~t^{-1/q}K=bD,$
-where $D$ is an ellipsoid and $x\in\mathbb{R}^{n}$. Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is not dependent on $t$ and $r$, which concludes the proof. ∎
+If equality holds, then (4.1) is an equality for almost all $(t,r)\in(0,\infty)^{2}$.
+For sufficiently large $r>0$, the assumptions of Theorem 2.2 are satisfied.
+Consequently, up to sets of measure zero, we have
+$\{f\geq t\}=x+aD,\qquad r^{-1/q}K=bD,$
+where $a,b>0$, $x\in\mathbb{R}^{n}$, and $D$ is a centered ellipsoid with the same volume as $K$.
+It follows that $K=\big(\frac{|K|}{|D|}\big)^{1/n}D$, showing that $D$ is independent of $t$ and $r$, which completes the proof.
+∎
+The next lemma shows that, for $\alpha\in(0,n)$, the volume of ${\rm L}_{\alpha}f$ does not increase under Schwarz symmetrals.
 Lemma 4.2.
-Let $\alpha\in(0,n)$. For non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
+Let $\alpha\in(0,n)$. For non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
 $|{\rm L}_{\alpha}f|\leq|{\rm L}_{\alpha}f^{\star}|.$
-If $|{\rm L}_{\alpha}f^{\star}|<\infty$, there is equality if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi$.
+If $|{\rm L}_{\alpha}f^{\star}|<\infty$, there is equality if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
 Proof.
@@
 By the monotone convergence theorem,
-$\lim_{k\rightarrow\infty}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f_{k}(x),f_{k}(x+tu)\}dxdt=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
+$\lim_{k\rightarrow\infty}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f_{k}(x),f_{k}(x+ru)\}dxdr=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr,$
 with the convergence being monotone increasing. Applying the monotone convergence theorem once more then gives
-$\lim_{k\rightarrow\infty}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{$ $f_{k}(x),f_{k}(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du$
-$=\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du.$
+$\lim_{k\rightarrow\infty}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{$ $f_{k}(x),f_{k}(x+ru)\}dxdr\Big)^{\frac{n}{\alpha}}du$
+$=\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr\Big)^{\frac{n}{\alpha}}du.$
 That is,
 $\lim_{k\rightarrow\infty}|{\rm L}_{\alpha}f_{k}|=|{\rm L}_{\alpha}f|=\infty.$ (4.2)
-Since $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$ and $f_{k}$ has compact support, $|{\rm L}_{\alpha}f_{k}|<\infty$.
+Since $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$ and $f_{k}$ has compact support, $|{\rm L}_{\alpha}f_{k}|<\infty$.
 By the definition of the Schwarz symmetral, we have $(f_{k})^{\star}\leq f^{\star}$.
@@
 There is equality precisely when ${\rm L}_{\alpha}f$ is a ball, which is the case for radially symmetric functions.
-For the second inequality, noting that $\|f\|_{\frac{n}{n+\alpha}}=\|f^{\star}\|_{\frac{n}{n+\alpha}}$ and ${\rm L}_{\alpha}f^{\star}$ is a ball, we obtain from Theorem Theorem 1, (4.1), and Lemma 5.2 that
+For the second inequality, noting that $\|f\|_{\frac{n}{n+\alpha}}=\|f^{\star}\|_{\frac{n}{n+\alpha}}$ and that ${\rm L}_{\alpha}f^{\star}$ is a ball, we obtain from the chord Sobolev inequality (1.3), (3.1), and Lemma 4.2 that
 $\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$
@@
 $\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
-If equality holds throughout, it follows from Theorem Theorem 1 that $f=c\operatorname{1}_{B}$ with $B$ a ball.
-By Lemma 5.2, $f=f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
+If equality holds throughout, it follows from the chord Sobolev inequality that $f^{\star}=c\operatorname{1}_{B}$, where $B$ is a ball.
+By Lemma 4.2, we have $f=f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
 This characterizes the equality cases.
 ∎
-6. Affine isoperimetric inequalities for $\alpha>n$
-As mentioned in the introduction, Theorem 1 extends the classical isoperimetric inequality for chord power integrals to function.
-It also provides an $L^{1}$ analogue of the HLS inequalities, for which both inequalities reverse when $\alpha>n$.
-In our setting, however, this reversal does not appear in such a simple form here.
-Before addressing this nontrivial case, we first present the following lemma, which serves as a counterpart to Lemma 5.1
+For $\alpha>n$, we state the following analogue of Lemma 4.1, which also generalizes [4, Lemma 3.2]. The proof follows a similar approach and is included for completeness.
 Lemma 4.3.
@@
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$
-Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to sets of measure zero.
+Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost $t>0$, the level set $\{f\geq t\}$ has measure zero or is homothetic to $D$ up to sets of measure zero.
 Proof.
 […]
-$\|z\|_{K}^{q}=\int_{0}^{\infty}\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(z)dt.$
+$\|z\|_{K}^{q}=\int_{0}^{\infty}\operatorname{1}_{\mathbb{R}^{n}\setminus r^{1/q}K}(z)dr.$
 Using Fubini’s theorem, we obtain
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),$ $f(y)\}\|x-y\|_{K}^{q}dxdy$ (4.3)
-$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\|x-y\|_{K}^{q}\operatorname{1}_{\{f\geq r\}}(y)dxdydr$
-$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt$
-$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)(1-\operatorname{1}_{t^{1/q}K}(x-y))\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt.$
+$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\|x-y\|_{K}^{q}\operatorname{1}_{\{f\geq t\}}(y)dxdydt$
+$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{\mathbb{R}^{n}\setminus r^{1/q}K}(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdydtdr$
+$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)(1-\operatorname{1}_{r^{1/q}K}(x-y))\operatorname{1}_{\{f\geq t\}}(y)dxdydtdr.$
 The Riesz rearrangement inequality, Theorem 2.2, implies that
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{1/q}K}(x-y)\operatorname{1}$ ${}_{\{f\geq r\}}(y)dxdy$ (6.2)
-$\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy.$
-Note that $\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx<\infty$ for almost all $r>0$, as $f\in L^{1}(\mathbb{R}^{n})$. Combing (6.2) with the calculation in (6.1), we have
+$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{1/q}K}(x-y)$ $\operatorname{1}_{\{f\geq t\}}(y)dxdy$ (4.4)
+$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq t\}}(x)\operatorname{1}_{r^{1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq t\}}(y)dxdy.$
+Note that $\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)dx<\infty$ for almost all $t>0$, as $f\in L^{1}(\mathbb{R}^{n})$. Combing (4.4) with the calculation in (4.3), we have
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}\|x-y\|_{K}^{q}dxdy\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f^{\star}(x),f^{\star}(y)\}\|x-y\|_{K^{\star}}^{q}dxdy.$ (4.5)
-Moreover, if equality holds in (6.3), then there exists a null set $N\subset(0,\infty)^{2}$ such that equality holds in (6.2) for all $(r,t)\in(0,\infty)^{2}\setminus N$.
-For almost every $t\in(0,\infty)$, we have $(r,t)\in(0,\infty)^{2}\setminus N$ for almost every $r>0$.
-For $t$ sufficiently small, Theorem 2.2 guarantees the existence of a centered ellipsoid $D$ and $x\in\mathbb{R}^{n}$ such that
-$\{f\geq r\}=x+aD\quad\text{and }\quad t^{1/q}K=bD.$
+Moreover, if equality holds in (4.5), then there exists a null set $N\subset(0,\infty)^{2}$ such that equality holds in (4.4) for all $(t,r)\in(0,\infty)^{2}\setminus N$.
+For almost all $r\in(0,\infty)$, we have $(t,r)\in(0,\infty)^{2}\setminus N$ for almost all $t>0$.
+For $r$ sufficiently small, Theorem 2.2 guarantees the existence of a centered ellipsoid $D$ and $x\in\mathbb{R}^{n}$ such that
+$\{f\geq t\}=x+aD\quad\text{and }\quad r^{1/q}K=bD.$
 Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is independent of $r$ and $t$, which concludes the proof. ∎
-With the previous lemma established, for $\alpha>n$, we have
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$ (6.4)
-$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{\{f^{\star}\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdy$
-$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
-However, Lemma 3.2 is not applicable in this case because the inequality holds in the opposite direction. We therefore apply a different approach to obtain the result as follows.
-Theorem 6.2.
-Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
-$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
-Equality holds if and only if $f=c\operatorname{1}_{B}$ with $B$ a ball and $c>0$.
-Proof.
-As we calculated in (6.4),
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy\geq\sigma_{n,\alpha}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
-By Jensen’s inequality, we have
-$\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\geq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|dr\Big)^{\frac{n+\alpha}{n}}=\|f\|_{\infty}^{-\frac{n+\alpha}{n}}\|f\|_{1}^{\frac{n+\alpha}{n}},$ (6.5)
-and thus
-$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$ (6.6)
-If equality holds in (6.6), then equality must occur in the application of Jensen’s inequality in (6.5).
-This in turn implies that $|\{f\geq r\}|$ is constant for almost every $r\in(0,\|f\|_{\infty})$.
-Equivalently, this happens if and only if $f=c\operatorname{1}_{E}$, where $E$ is a measurable set of finite measure and $c>0$.
-When $f=c\operatorname{1}_{E}$, the inequality (6.6) reduces to the geometric case:
-$I_{\alpha+1}(E)\geq\Big(\frac{I_{\alpha+1}(B^{n})}{\omega_{n}^{(n+\alpha)/n}}\Big)|E|^{\frac{n+\alpha}{n}},$
-with equality if and only if $E$ is a ball. This then completes the characterization of the equality cases. ∎
+Similarly, for $\alpha>n$, we show that $|{\rm L}_{\alpha}f|$ does not decrease under Schwarz symmetrals.
 Lemma 4.4.
@@
 Proof.
-By letting $q=n-\alpha$ in Lemma 6.2 and combining it with (4.1), we obtain
-$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\geq n\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
+By letting $q=\alpha-n$ in Lemma 4.3 and combining it with (3.1), we obtain
+$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\geq n\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
 With $K={\rm L}_{\alpha}f$ and the dual mixed volume inequality (2.3), we have
-$|{\rm L}_{\alpha}f|$ $\geq\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
+$|{\rm L}_{\alpha}f|$ $=\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
 $\geq\tilde{V}_{\alpha}(({\rm L}_{\alpha}f)^{\star},{\rm L}_{\alpha}f^{\star})$
@@
 The equality cases follows from Lemma 4.3. ∎
-Building on the results established above, we proceed to the proof of Theorem 3.
+Building on the preceding lemmas, we now establish the affine chord Sobolev inequalities for $\alpha>n$.
 Proof of Theorem 2.
-For the first inequality, if the left side is $\infty$, the inequality holds trivially. We then assume that
+If the right-most side is $\infty$, the inequality holds trivially. We then assume that
 $\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy<\infty.$
@@
 By the equality case in the dual mixed volume inequality, ${\rm L}_{\alpha}f$ must be a ball, which occurs when $f$ is radially symmetric.
-For the second inequality, we may assume that $|{\rm L}_{\alpha}f|$ is finite. Note that ${\rm L}_{\alpha}f^{\star}$ is a ball. Then Lemma 6.3 and (4.1) implies that
+For the first inequality, we may assume that $|{\rm L}_{\alpha}f|$ is finite. Note that ${\rm L}_{\alpha}f^{\star}$ is a ball. Then Lemma 4.4 and (3.1) imply that
 $n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f^{\star}|^{\frac{\alpha}{n}}=n\tilde{V}_{\alpha}(B^{n},{\rm L}_{\alpha}f^{\star})=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy.$
-Combining Theorem 6.2 for the right-hand side with the facts $\|f\|_{1}=\|f^{\star}\|_{1}$ and $\|f\|_{\infty}=\|f^{\star}\|_{\infty}$, we deduce
+Combining the chord Sobolev inequality (1.4) with $\|f\|_{1}=\|f^{\star}\|_{1}$ and $\|f\|_{\infty}=\|f^{\star}\|_{\infty}$, we deduce
 $\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
-The equality characterization is given by Theorem 6.2 and Lemma 6.3. ∎
-7. Radial mean bodies
+The equality case follows from the optimizers of (1.4) together with Lemma 4.4. ∎
+5. Radial mean bodies for log-concave functions
 In this section, we begin by recalling some basic notation and properties of the classical radial mean bodies.
@@
 for $\alpha\neq 0$ and by
-$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx$
-for $\alpha=0$.
+$\log\rho_{{\rm R}_{0}K}(u)=\frac{1}{|K|}\int_{K}\log\rho_{K-x}(u)dx.$
 They further defined ${\rm R}_{\infty}K$ consistently by
 $\rho_{{\rm R}_{\infty}K}(u)=\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}K}(u)=\max_{x\in K}\rho_{K-x}(u),$
-so that ${\rm R}_{\infty}K$ coincides with the difference body of $K$, namely
-${\rm R}_{\infty}K=DK=K-K.$
-As introduced in Section 2, $\rho_{{\rm R}_{\alpha}K}(u)$ can be expressed as the integral of the chord length $|K\cap(x^{\prime}+\mathbb{R}u)|$.
-It is also closely related to the covariogram of $K$, that is
-$g_{K}(x)=|K\cap(K+x)|.$
-In fact
-$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\rho_{DK}(u)}t^{\alpha-1}g_{K}(tu)dt$
+so that ${\rm R}_{\infty}K$ coincides with the difference body of $K$, namely,
+${\rm R}_{\infty}K={\rm D}K=K-K.$
+5.1. Affine isoperimetric inequalities for $\boldsymbol{{\rm R}_{\alpha}f}$
+Haddad and Ludwig [24] proved the following affine isoperimetric inequality for ${\rm R}_{\alpha}K$, with full equality characterization.
+Theorem 5.1.
+[24, Theorem 20] Let $K\subset\mathbb{R}^{n}$ be a convex body. Then
+$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha\in(-1,n),$
+$\frac{|{\rm R}_{\alpha}K|}{|K|}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},$ $\alpha>n,$
+with equality if and only if $K$ is an ellipsoid.
+For $\alpha>-1$, the authors [4] defined the radial mean body ${\rm R}_{\alpha}f$ for a non-zero, log-concave function $f\in L^{1}(\mathbb{R}^{n})$ by
+$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t),~~\alpha\neq 0,$ (5.1)
+and
+$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)d\mu_{f}(t),$ (5.2)
+where $d\mu_{f}(t)=\frac{|\{f\geq t\}|}{\|f\|_{1}}dt$ is a probability measure.
+The following result in [4] shows that ${\rm R}_{\alpha}f$ is proportional to the fractional polar projection body $\Pi^{*,-\alpha}f$ and the star-shaped set ${\rm L}_{\alpha}f$.
+Moreover, by the continuity of $\rho_{{\rm R}_{\alpha}f}(u)$ with respect to $\alpha$, the authors also gave an explicit formula for $\rho_{{\rm R}_{0}f}$.
+Lemma 5.2.
+[4, Proposition 6.2, Lemma 6.4] For non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
+$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dxdr$ (5.3)
+for $\alpha\in(-1,0)$ and
+$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr$ (5.4)
 for $\alpha>0$.
-Note that, for $\alpha>0$,
-$\rho_{{\rm L}_{\alpha}\operatorname{1}_{K}}(u)^{\alpha}$ $=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{K}(x),\operatorname{1}_{K}(x+tu)\}dxdt$
-$=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\operatorname{1}_{K}(x)\operatorname{1}_{K}(x+tu)dxdt$
-$=\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt$
-$=\int_{0}^{\infty}t^{\alpha-1}g_{K}(tu)dt.$
-Hence,
-${\rm L}_{\alpha}\operatorname{1}_{K}=\Big(\frac{|K|}{\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.1)
-for $\alpha>0$. Recall that in Section 2.5,
-$\Pi^{*,-\alpha}\operatorname{1}_{K}=\Big(\frac{2|K|}{-\alpha}\Big)^{1/\alpha}{\rm R}_{\alpha}K$ (7.2)
-for $\alpha\in(-1,0)$.
-Combining (7.1) and (7.2) with Theorem 2 and Theorem 3, we obtain the following Pólya–Szegö inequality:
-$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\leq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\leq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
-for $\alpha\in(0,n)$ and
-$\frac{I_{\alpha+1}(K)}{I_{\alpha+1}(B^{n})}\geq\Big(\frac{|{\rm R}_{\alpha}K|}{|{\rm R}_{\alpha}{B^{n}}|}\Big)^{\alpha/n}\geq\Big(\frac{|K|}{|B^{n}|}\Big)^{(n+\alpha)/n},$
-for $\alpha>n$. These two formulas can also be regarded as the affine isoperimetric inequalities for chord power integrals.
-8. Radial mean bodies for log-concave functions
-By (7.1) and (7.2), we define the radial mean bodies for log-concave functions, by
-${\rm R}_{\alpha}f=\Big(\frac{-\alpha}{2\|f\|_{1}}\Big)^{1/\alpha}\Pi^{*,-\alpha}f$ (8.1)
-for $\alpha\in(-1,0)$ and
-${\rm R}_{\alpha}f=\Big(\frac{\alpha}{\|f\|_{1}}\Big)^{1/\alpha}{\rm L}_{\alpha}f$
-for $\alpha>0$.
-The following proposition describes the relation between ${\rm R}_{\alpha}f$ and the radial $\alpha$-th mean body of the superlevel set $\{f\geq r\}$, for $\alpha\in(-1,0)\cup(0,\infty)$.
-Proposition 8.1.
-Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. For $\alpha\in(-1,0)\cup(0,\infty)$,
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r),$
-where $d\mu_{f}(r)=\frac{|\{f\geq r\}|}{\|f\|_{1}}dr$.
+For $\alpha=0$,
+$\log\rho_{{\rm R}_{0}f}(u)=-\gamma+\int_{0}^{\infty}\frac{1}{r}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx-e^{-r}\Big)dr,$
+where $\gamma=\Gamma^{\prime}(1)$ is the Euler constant.
+The following lemma describes the affine covariance of ${\rm R}_{\alpha}f$, where ${\rm GL}(n)$ denotes the group of general linear transformations on $\mathbb{R}^{n}$.
+Lemma 5.3.
+Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
+${\rm R}_{\alpha}(cf)={\rm R}_{\alpha}f~~\text{and}~~{\rm R}_{\alpha}(f\circ\tau_{y})={\rm R}_{\alpha}f,$
+where $c>0$ is a constant, $y\in\mathbb{R}^{n}$ and $f\circ\tau_{y}(x)=f(x-y)$. Moreover, for $\phi\in{\rm GL}(n)$,
+${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f.$
 Proof.
-First we consider $\alpha>0$. Recall that
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
-where
-$\min\{f(x),f(x+tu)\}=\int_{0}^{\|f\|_{\infty}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(x+tu)dr.$
-By Fubini’s theorem, we then obtain
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{\alpha}{\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dtdr.$
-Note that for a convex body $K$,
-$\rho_{{\rm R}_{\alpha}K}(u)^{\alpha}=\frac{\alpha}{|K|}\int_{0}^{\infty}t^{\alpha-1}|K\cap(K+tu)|dt.$
-We have
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}\cdot\frac{|\{f\geq r\}|}{\|f\|_{1}}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
-Second, when $\alpha\in(-1,0)$, the definition (8.1) and (2.4) imply that
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}|f(x+tu)-f(x)|dxdt.$
-Since
-$|f(x+tu)-f(x)|=\int_{0}^{\infty}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|dr,$
-we obtain
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}$ $=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\int_{0}^{\|f\|_{\infty}}|\operatorname{1}_{\{f\geq r\}}(x+tu)-\operatorname{1}_{\{f\geq r\}}(x)|drdxdt$
-$=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dtdr.$
-By the calculation in (2.5) and (2.6), we have
-$\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}=\frac{-\alpha}{2\{f\geq r\}}\int_{0}^{\infty}t^{\alpha-1}|\{f\geq r\}\triangle(\{f\geq r\}+tu)|dt,$
-which implies
-$\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\frac{-\alpha}{2\|f\|_{1}}\int_{0}^{\|f\|_{\infty}}\frac{2|\{f\geq r\}|}{-\alpha}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr=\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r).$
-This completes the proof. ∎
-By the continuity of the $L^{p}$-norm in $p$, we define
-$\log\rho_{{\rm R}_{0}f}(u)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r).$ (8.2)
-A more explicit formula for ${\rm R}_{0}f$ will be presented in Section 9.
-We recall the following convergence result from [25], which shows that as $\alpha\rightarrow-1$, the radial mean body ${\rm R}_{\alpha}f$ converges to the polar projection body of $f$.
+By (5.3) and (5.4), the radial mean body ${\rm R}_{\alpha}f$ is invariant under scaling and translations. It suffices to show ${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f$.
+We treat the case $\alpha>n$. The case $\alpha\in(-1,0)$ can be handled similarly. For $\phi\in{\rm GL}(n)$, note that $\|f\circ\phi^{-1}\|_{1}=|\det\phi|\|f\|_{1}$. Since
+$\int_{\mathbb{R}^{n}}\min\{f(\phi^{-1}x),f(\phi^{-1}(x+ru))\}dx=|\det\phi|\int_{\mathbb{R}^{n}}\min\{f(x),f(x+r\phi^{-1}u)\}dx,$
+by (5.4), we have
+$\rho_{{\rm R}_{\alpha}(f\circ\phi^{-1})}(u)^{\alpha}=\rho_{{\rm R}_{\alpha}f}(\phi^{-1}u)^{\alpha}=\rho_{\phi{\rm R}_{\alpha}f}(u)^{\alpha},$
+which concludes the proof. ∎
+Recall that the affine chord Sobolev inequalities can be written as
+$2\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\leq n\omega_{n}^{\frac{n-\alpha}{n}}|\Pi^{*,-\alpha}f|^{\frac{\alpha}{n}},$ $\alpha\in(-1,0),$ (5.5)
+$\sigma_{n,\alpha}\|f\|_{\frac{n}{n+\alpha}}$ $\geq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},$ $\alpha\in(0,n)$
+and
+$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}},~~\alpha>n.$ (5.6)
+By (5.4) and the affine chord Sobolev inequalities, we establish the following affine isoperimetric inequalities for ${\rm R}_{\alpha}f$. We first focus on non-negative $\alpha$.
+Note that, when $\alpha=n$ and $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ is log-concave, the chord Sobolev inequalities become the following two trivial inequalities:
+$\frac{\|f\|_{1}^{2}}{\|f\|_{\infty}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq\|f\|_{\frac{1}{2}},$ (5.7)
+where equality holds if and only if $f=c\operatorname{1}_{K}$ for some convex body $K\subset\mathbb{R}^{n}$ and constant $c>0$. Since
+$|{\rm R}_{n}f|=\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy,$
+and $|{\rm R}_{n}K|=|K|$ for any convex body $K\subset\mathbb{R}^{n}$, both (5.8) and (5.9) hold for $\alpha=n$ and equality holds if and only if $f=c\operatorname{1}_{K}$.
+Theorem 5.4.
+Let $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ be log-concave and non-zero. Then
+$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}$ $\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},\quad\alpha\in(-1,0)\cup(0,n),$ (5.8)
+$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{1}\|f\|_{\infty}^{-1}}$ $\geq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|},\quad\alpha>n.$ (5.9)
+In both cases, equality holds if and only if $f$ is a constant multiple of the characteristic function of an ellipsoid.
+Proof.
+Inequality (5.9) and the equality characterization follow from (5.6) directly. We turn to $\alpha\in(-1,0)\cup(0,n)$. By the affine chord Sobolev inequality (5.5), we obtain
+$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{n}{n+\alpha}}^{n/\alpha}\|f\|_{1}^{-n/\alpha}}\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|}.$
+For brevity, we denote by $\bar{f}(x)=f(x)/\|f\|_{1}$.
+Then $\bar{f}(x)dx$ is a probability measure on $\mathbb{R}^{n}$.
+Combining this with the monotonicity of $L^{p}$ norms, for $\alpha\in(-1,n)\setminus\{0\}$, we have
+$\|f\|_{\frac{n}{n+\alpha}}^{n/\alpha}\|f\|_{1}^{-n/\alpha}$ $=\Big(\int_{\mathbb{R}^{n}}\bar{f}(x)^{\frac{n}{n+\alpha}}dx\Big)^{\frac{n+\alpha}{\alpha}}$ (5.10)
+$=\Big(\int_{\mathbb{R}^{n}}\big(\bar{f}(x)^{-1}\big)^{\frac{\alpha}{n+\alpha}}\bar{f}(x)dx\Big)^{\frac{n+\alpha}{\alpha}}\leq\Big(\int_{\mathbb{R}^{n}}\bar{f}(x)^{\frac{1}{2}}dx\Big)^{2}=\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1},$
+where equality holds if and only if $f=c\operatorname{1}_{K}$ for some $c>0$ and convex body $K\subset\mathbb{R}^{n}$. Therefore, we obtain
+$\frac{|{\rm R}_{\alpha}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{\alpha}B^{n}|}{|B^{n}|}.$ (5.11)
+If equality holds, then (5.10) implies that $f=c\operatorname{1}_{K}$ where $K\subset\mathbb{R}^{n}$ is a convex body and $c>0$.
+Hence (5.11) reduces to the geometric inequality (5.1), whose equality cases imply that $K$ is an ellipsoid.
+∎
+Next we consider the case $\alpha=0$.
+For non-zero, log-concave $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ and fixed $u\in\mathbb{S}^{n-1}$, the radial function $\rho_{{\rm R}_{\alpha}f}(u)$ is continuous in $\alpha>-1$, as shown in [4].
+Therefore, taking limits on both sides of (5.11) yields the affine isoperimetric inequality for ${\rm R}_{0}f$, but the equality characterization does not follow from this argument.
+In the following theorem, we use (5.1) together with the definition of ${\rm R}_{0}f$ to establish the inequality with its equality characterization.
+Theorem 5.5.
+Let $f\in L^{\frac{1}{2}}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
+$\frac{|{\rm R}_{0}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|},$
+where equality holds if and only if $f$ is the characteristic function of an ellipsoid.
+Proof.
+By (5.2),
+$\rho_{{\rm R}_{0}f}(u)^{n}=\exp\Big(\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t)\Big).$
+Since $e^{t}$ is a convex function on $\mathbb{R}$ and $\mu_{f}$ is a probability on $\mathbb{R}$, it follows from Jensen’s inequality that
+$\rho_{{\rm R}_{0}f}(u)^{n}\leq\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t).$
+This, together with (5.1), implies that
+$|{\rm R}_{0}f|=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\rho_{{\rm R}_{0}f}(u)^{n}du$ $\leq\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\|f\|_{\infty}}\rho_{{\rm R}_{0}\{f\geq t\}}(u)^{n}d\mu_{f}(t)du$
+$=\int_{0}^{\|f\|_{\infty}}|{\rm R}_{0}\{f\geq t\}|d\mu_{f}(t)$
+$\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}\int_{0}^{\|f\|_{\infty}}\frac{|\{f\geq t\}|^{2}}{\|f\|_{1}}dt$
+$=\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|f\|_{1}}dxdy.$
+By (5.7), we obtain
+$\frac{|{\rm R}_{0}f|}{\|f\|_{\frac{1}{2}}\|f\|_{1}^{-1}}\leq\frac{|{\rm R}_{0}B^{n}|}{|B^{n}|}.$ (5.12)
+If equality holds, then
+$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy=\|f\|_{\frac{1}{2}},$
+which implies that $f=c\operatorname{1}_{K}$ for some convex body $K\subset\mathbb{R}^{n}$ and $c>0$.
+Hence (5.12) reduces to the geometric inequality in (5.1).
+By the equality characterization of (5.1), $K$ is an ellipsoid.
+∎
+5.2. Monotonicity of $\boldsymbol{{\rm R}_{\alpha}f}$
+To consider the monotonicity of radial mean bodies, we recall the following result by Haddad and Ludwig [24].
+Lemma 5.6.
+[24, Theorem 3] For $\alpha\in(-1,0)$ and $f\in L^{1}(\mathbb{R}^{n})$ with bounded variation,
+$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1}$
+for $u\in\mathbb{S}^{n-1}$.
+We mainly use Lemma 5.6 for log-concave functions $f\in L^{1}(\mathbb{R}^{n})$.
+By [4, Lemma 5.4], $\mathcal{H}^{n-1}(\partial\{f\geq t\})$ is a integrable function on $\mathbb{R}$.
+Hence the coarea formula [14, Theorem 5.9], implies that $f$ has bounded variation.
 The opposite endpoint $\alpha=\infty$ will be considered below.
 Lemma 5.7.
-[25, Theorem 3] For $\alpha\in(-1,0)$ and log-concave $f\in L^{1}(\mathbb{R}^{n})$,
-$\lim_{\alpha\rightarrow-1^{+}}(1+\alpha)\|f\|_{1}\rho_{{\rm R}_{\alpha}f}(u)^{\alpha}=\rho_{\Pi^{*}f}(u)^{-1}$
+Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave and non-zero. If $f$ has compact support, then
+$\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)=\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u),$
 for $u\in\mathbb{S}^{n-1}$.
-Lemma 8.3.
-Let $f\in L^{1}(\mathbb{R}^{n})$ be log-concave and non-zero. If $f$ has compact support, then
-$\lim_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)=\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u),$
-for $u\in\mathbb{S}^{n-1}$.
-Before proving the lemma below, we recall a result from [18, Theorem 5.5], which states that for every convex body $K$
-${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K,$
-where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. The complete statement of the theorem will be given in Section 9.
 Proof.
 Since Jensen’s inequality implies that ${\rm R}_{\alpha}K\subset{\rm D}K$ for $\alpha>-1$, we have
-$\rho_{{\rm R}_{\alpha}f}(u)\leq\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha}.$
+$\rho_{{\rm R}_{\alpha}f}(u)\leq\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t)\Big)^{1/\alpha}.$
 Letting $\alpha\rightarrow\infty$, we obtain
-$\limsup_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\leq\sup_{r\in(0,1)}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.3)
-On the other hand, by the reverse inclusion ${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K$,
-$\rho_{{\rm R}_{\alpha}f}(u)\geq c_{n,\alpha}^{-1}\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f>r\}}(u)^{\alpha}d\mu_{f}(r)\Big)^{1/\alpha},$
+$\limsup_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\leq\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u).$ (5.13)
+On the other hand, by the reversed inclusion ${\rm D}K\subset c_{n,\alpha}{\rm R}_{\alpha}K$ in (1.12),
+$\rho_{{\rm R}_{\alpha}f}(u)\geq c_{n,\alpha}^{-1}\Big(\int_{0}^{\infty}\rho_{{\rm D}\{f\geq t\}}(u)^{\alpha}d\mu_{f}(t)\Big)^{1/\alpha},$
 where $c_{n,\alpha}=(nB(\alpha+1,n))^{-1/\alpha}$. Note that $\lim_{\alpha\rightarrow\infty}c_{n,\alpha}=1$. Therefore we have
-$\liminf_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\geq\sup_{r}\rho_{{\rm D}\{f\geq r\}}(u)=\rho_{{\rm D}{\rm supp}(f)}(u).$ (8.4)
+$\liminf_{\alpha\rightarrow\infty}\rho_{{\rm R}_{\alpha}f}(u)\geq\sup_{t>0}\rho_{{\rm D}\{f\geq t\}}(u)=\rho_{{\rm D}\,{\rm supp}(f)}(u).$ (5.14)
 Combining (5.13) with (5.14), we then conclude the proof. ∎
+This yields the natural definition of ${\rm R}_{\infty}f$ given below.
 Definition 5.8.
@@
 for $1<\alpha<\beta$.
-Applying Jensen’s inequality together with (8.5), we immediately obtain
+Applying Jensen’s inequality, together with (5.15), we immediately obtain
 ${\rm R}_{\alpha}f\subset{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
-for $1<\alpha<\beta$. The following theorem provides a functional analogue of (8.6).
+for $1<\alpha<\beta<\infty$. The following theorem provides a functional analogue of (5.16).
 Theorem 5.9.
@@
 $\|f\|_{\infty}\Pi^{*}f\subset\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f\subset{\rm R}_{\infty}f,$
-for $1<\alpha<\beta$.
+for $1<\alpha<\beta<\infty$.
 Proof.
-If $f$ is not compactly supported, the right inclusion is trivial since ${\rm R}_{\infty}f=\mathbb{R}^{n}$. We now assume that $f$ is with compact support.
-By Proposition 8.1,
-$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}(1+\alpha)|\{f\geq r\}|\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
-$=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\alpha,r}{\rm R}_{\alpha}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha},$
-where $l_{\alpha,r}=((1+\alpha)|\{f\geq r\}|)^{1/\alpha}$ It follows from (8.6) that
-$l_{\alpha,r}\rho_{{\rm R}_{\alpha}\{f\geq r\}}(u)\leq l_{\beta,r}\rho_{{\rm R}_{\beta}\{f\geq r\}}(u),$
+If $f$ is not compactly supported, the right-most inclusion is trivial since ${\rm R}_{\infty}f=\mathbb{R}^{n}$. We now assume that $f$ has compact support.
+By the definition of ${\rm R}_{\alpha}f$ in (5.1),
+$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}(1+\alpha)|\{f\geq t\}|\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha}$
+$=\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\alpha,t}{\rm R}_{\alpha}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha},$
+where $l_{\alpha,t}=((1+\alpha)|\{f\geq t\}|)^{1/\alpha}$. It follows from (5.16) that
+$l_{\alpha,t}\rho_{{\rm R}_{\alpha}\{f\geq t\}}(u)\leq l_{\beta,t}\rho_{{\rm R}_{\beta}\{f\geq t\}}(u),$
 for $1<\alpha<\beta$.
 Jensen’s inequality then implies that
-$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\alpha}dr\Big)^{1/\alpha}$
-$\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,r}{\rm R}_{\beta}\{f\geq r\}}(u)^{\beta}dr\Big)^{1/\beta}$
+$\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}\rho_{{\rm R}_{\alpha}f}(u)$ $\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,t}{\rm R}_{\beta}\{f\geq t\}}(u)^{\alpha}dt\Big)^{1/\alpha}$
+$\leq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}\rho_{l_{\beta,t}{\rm R}_{\beta}\{f\geq t\}}(u)^{\beta}dt\Big)^{1/\beta}$
 $=\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}\rho_{{\rm R}_{\beta}f}(u),$
@@
 $\Big((1+\alpha)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\alpha}{\rm R}_{\alpha}f\subset\Big((1+\beta)\frac{\|f\|_{1}}{\|f\|_{\infty}}\Big)^{1/\beta}{\rm R}_{\beta}f.$
-Letting $\beta\rightarrow\infty$, together with $\lim_{\beta\rightharpoonup\infty}(1+\beta)^{1/\beta}=1$, we obtain the right inclusion.
-Letting $\alpha\rightarrow-1$, by Lemma 8.2, we obtain the left inclusion.
+Letting $\beta\rightarrow\infty$, together with $\lim_{\beta\rightarrow\infty}(1+\beta)^{1/\beta}=1$, we obtain the right-most inclusion.
+Letting $\alpha\rightarrow-1^{+}$, the left-most inclusion follows from Lemma 5.6.
 ∎
-9. Reversed inequalities
+6. Reverse inclusion for radial mean bodies
 In the previous section, we recalled the monotonicity relations (5.15) and (5.16) for radial mean bodies and extended them to their functional analogues.
 In this section, we turn to the reverse inclusion, which is considerably more delicate.
-We first recall the following theorem established by Gardner and Zhang [18]. In particular, for $\alpha=n$,
-$c_{n,n}{\rm R}_{n}K\subset n|K|\Pi^{*}K$
-which yields Zhang’s inequality [50].
+We begin by stating the following theorem established by Gardner and Zhang [18].
 Theorem 6.1.
@@
 In each inclusion equality holds if and only if $K$ is a simplex.
+When $\beta=n$, the left-most inclusion implies the Rogers-Shephard inequality:
+$|{\rm D}K|\leq\binom{2n}{n}|K|.$
+When $\alpha=n$, the right-most inclusion yields Zhang’s inequality:
+$n^{-n}\binom{2n}{n}=|\triangle|^{n-1}|\Pi^{*}\triangle|\leq|K|^{n-1}|\Pi^{*}K|,$
+where $\triangle$ denotes a simplex in $\mathbb{R}^{n}$.
 To extend Theorem 6.1 to the functional setting, we rely on the following result from [26].
-It generalizes the classical result of Marshall, Olkin and Proschan [41], whose simple proof was later given by Milman and Pajor [43].
-Lemma 9.2 ([27], Lemma 17).
-Let $\omega:[0,\infty)\to[0,\infty)$ be decreasing with
-$0<\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt<\infty$
+It generalizes the classical result of Marshall, Olkin and Proschan [38], whose proof was later simplified by Milman and Pajor [40].
+Lemma 6.2.
+[26, Lemma 17] Let $\omega:[0,\infty)\to[0,\infty)$ be decreasing with
+$0<\int_{0}^{\infty}r^{\alpha-1}\omega(r)dr<\infty$
 for $\alpha>0$ and
-$0<\int_{0}^{\infty}t^{\alpha-1}(\omega(0)-\omega(t))dt<\infty$
-for $\alpha\in(-1,0)$. If $\varphi:[0,\infty)\to[0,\infty)$ is non-zero, with $\varphi(0)=0$, and such that $t\mapsto\varphi(t)$ and $t\mapsto{\varphi(t)}/{t}$ are increasing on $(0,\infty)$, then
-$\zeta(\alpha)=\left\{\begin{aligned} &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}\omega(\varphi(t))dt}{\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt}\right)^{1/\alpha}&&\text{for}~~\alpha>0,\\ &\exp\left(\int_{0}^{\infty}\frac{\omega(\varphi(t))-\omega(t)}{t\omega(0)}dt\right)&&\text{for}~~\alpha=0,\\ &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}(\omega(\varphi(t))-\omega(0))dt}{\int_{0}^{\alpha-1}(\omega(t)-\omega(0))dt}\right)^{1/\alpha}&&\text{for}~~-1<\alpha<0.\end{aligned}\right$.
-is a continuous, decreasing function of $\alpha$ on $(-1,\infty)$. Moreover, $\zeta$ is constant on $(-1,\infty)$ if $\varphi(t)=\lambda t$ on $0,\infty)$ for some $\lambda>0$.
-In particular, taking $\omega(t)=e^{-t}$ and $\omega(t)=(1-st)_{+}^{1/s}$ with $s>0$ yields the classical analytic continuation formulas for the Gamma and Beta functions,
+$0<\int_{0}^{\infty}r^{\alpha-1}(\omega(0)-\omega(r))dr<\infty$
+for $\alpha\in(-1,0)$. If $\varphi:[0,\infty)\to[0,\infty)$ is non-zero, with $\varphi(0)=0$, and such that $r\mapsto\varphi(r)$ and $r\mapsto{\varphi(r)}/{r}$ are increasing on $(0,\infty)$, then
+$\zeta(\alpha)=\left\{\begin{aligned} &\left(\frac{\int_{0}^{\infty}r^{\alpha-1}\omega(\varphi(r))dr}{\int_{0}^{\infty}r^{\alpha-1}\omega(r)dr}\right)^{1/\alpha}&&\text{for}~~\alpha>0,\\ &\exp\left(\int_{0}^{\infty}\frac{\omega(\varphi(r))-\omega(r)}{r\omega(0)}dr\right)&&\text{for}~~\alpha=0,\\ &\left(\frac{\int_{0}^{\infty}r^{\alpha-1}(\omega(\varphi(r))-\omega(0))dr}{\int_{0}^{\alpha-1}(\omega(r)-\omega(0))dr}\right)^{1/\alpha}&&\text{for}~~-1<\alpha<0,\end{aligned}\right$.
+is a continuous, decreasing function of $\alpha$ on $(-1,\infty)$. Moreover, $\zeta$ is constant on $(-1,\infty)$ if $\varphi(r)=\lambda r$ on $0,\infty)$ for some $\lambda>0$.
+In particular, taking $\omega(r)=e^{-r}$ and $\omega(r)=(1-sr)_{+}^{1/s}$ with $s>0$ yields the classical analytic continuation formulas for the Gamma and Beta functions,
 […]
@@
 […]
-We note here, in the setting of Lemma 9.2, that if $\zeta(\alpha)$ is constant on $(-1,\infty)$, say $\zeta(\alpha)=\lambda$, then $\varphi(t)=\lambda t$ for all $t\geq 0$.
-When $\omega(t)$ is strictly decreasing, this conclusion follows from the proof of Milman and Pajor, but for completeness, we include a proof below.
+Under the assumptions of Lemma 6.2 with strictly decreasing $\omega(r)$, if $\zeta(\alpha)$ is constant on $(-1,\infty)$, say $\zeta(\alpha)=\lambda$, then $\varphi(r)=\lambda r$ for all $r\geq 0$.
+The result follows directly from Milman and Pajor [40], but we provide a proof for completeness.
 Lemma 6.3.
@@
 Suppose $\omega$ is strictly decreasing and $1<\alpha<\beta<\infty$.
-Then $\zeta(\alpha)=\zeta(\beta)$ implies $\varphi(t)=\lambda t$ with $\lambda=\zeta(\alpha)^{-1}$.
+Then $\zeta(\alpha)=\zeta(\beta)$ implies $\varphi(r)=\lambda r$ with $\lambda=\zeta(\alpha)^{-1}$.
 Proof.
 Let $\lambda=\zeta(\alpha)^{-1}$. We first consider $0<\alpha<\beta<\infty$. For $\alpha>0$, we then have
-$\int_{0}^{\infty}\omega(\lambda t)t^{\alpha-1}dt=\int_{0}^{\infty}\omega(\varphi(t))t^{\alpha-1}dt.$
-We set $\Psi(r)=\int_{r}^{\infty}(\omega(\lambda t)-\omega(\varphi(t)))dt$, where $\Psi(0)=\Psi(\infty)=0$. Since $\varphi(t)/t$ is increasing while $\omega$ is decreasing, we have
-$\omega(\varphi(t))>\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}<\lambda,$
+$\int_{0}^{\infty}\omega(\lambda r)r^{\alpha-1}dr=\int_{0}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr.$
+We set $\Psi(t)=\int_{t}^{\infty}(\omega(\lambda r)-\omega(\varphi(r)))dr$, where $\Psi(0)=\Psi(\infty)=0$. Since $\varphi(r)/r$ is increasing while $\omega$ is decreasing, we have
+$\omega(\varphi(r))>\omega(\lambda r),~~\text{if}~~{\varphi(r)}/{r}<\lambda,$
 and
-$\omega(\varphi(t))\leq\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}\geq\lambda.$
-This implies the sign of $\omega(\lambda r)-\omega(\varphi(r))=-\Psi^{\prime}(r)$.
-Therefore, $\Psi$ is first increasing and then decreasing, which yields that $\Psi(r)\geq 0$ for all $r\geq 0$.
+$\omega(\varphi(r))\leq\omega(\lambda r),~~\text{if}~~{\varphi(r)}/{r}\geq\lambda.$
+This implies the sign of $\omega(\lambda t)-\omega(\varphi(t))=-\Psi^{\prime}(t)$.
+Therefore, $\Psi$ is first increasing and then decreasing, which yields that $\Psi(t)\geq 0$ for all $t\geq 0$.
 That is
-$\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\geq\int_{r}^{\infty}\omega(\varphi(t))dt.$
+$\int_{t}^{\infty}\omega(\lambda r)r^{\alpha-1}dr\geq\int_{t}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr.$
 For $\beta>\alpha>0$, we have
-$\int_{0}^{\infty}\omega(\varphi(t))t^{\beta-1}dt$ $=(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\varphi(t))r^{\alpha-1}dt\Big)dr$
-$\leq(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\Big)dr$
-$=\int_{0}^{\infty}\omega(\lambda t)t^{\beta-1}dt=\frac{1}{\lambda^{\beta}}\int_{0}^{\infty}\omega(t)t^{\beta-1}dt,$
-which implies $\zeta(\alpha)\geq\zeta(\beta)$.
-As we assumed, $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, equality holds in the inequality above, which implies that $\Psi(r)=0$ for all $r\geq 0$.
-Then we have
-$\Psi^{\prime}(r)=\omega(\varphi(t))-\omega(\lambda t)=0.$
+$\int_{0}^{\infty}\omega(\varphi(r))r^{\beta-1}dr$ $=(\beta-\alpha)\int_{0}^{\infty}t^{\beta-\alpha-1}\Big(\int_{t}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr\Big)dt$
+$\leq(\beta-\alpha)\int_{0}^{\infty}t^{\beta-\alpha-1}\Big(\int_{t}^{\infty}\omega(\lambda r)r^{\alpha-1}dr\Big)dt$
+$=\int_{0}^{\infty}\omega(\lambda r)r^{\beta-1}dr=\frac{1}{\lambda^{\beta}}\int_{0}^{\infty}\omega(r)r^{\beta-1}dr,$
+which implies $\zeta(\alpha)\geq\zeta(\beta)$. If $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, equality holds in the inequality above, which yields that $\Psi(t)=0$ for all $t\geq 0$. Then we have
+$\Psi^{\prime}(t)=\omega(\varphi(t))-\omega(\lambda t)=0.$
 Since $\omega$ is strictly decreasing, we obtain $\varphi(t)=\lambda t$.
@@
 6.1. Results for log-concave functions
-We begin this subsection by recalling a result from [2], which will serve as a main tool in the equality characterization in Theorem 3.
+In this subsection, for the sake of completeness, we provide a proof of (1.13), which extends (1.12) to the class of log-concave functions and was recently established in [28, Theorem 1.6].
+The following result from [3] serves as a main tool in the equality characterization of (1.13).
 Lemma 6.4.
-[2, Lemma 4.2] Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave such that $f(o)=\|f\|_{\infty}$. Let
+[3, Lemma 4.2] Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave such that $f(o)=\|f\|_{\infty}$. Let
 $g(x)=\int_{\mathbb{R}^{n}}\min\{f(y),f(y+x)\}dy.$
-Then for every $x\in\mathbb{R}^{n}$ and $\lambda\in[0,1]$, $g(\lambda x)=g(o)^{1-\lambda}g(x)^{\lambda}$ if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$ with $\triangle$ a simplex containing the origin.
-For the reader’s convenience, we restate Theorem 4 below before giving the proof.
+Then for every $x\in\mathbb{R}^{n}$ and $\lambda\in(0,1)$, $g(\lambda x)=g(o)^{1-\lambda}g(x)^{\lambda}$ if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$ with $\triangle$ a simplex containing the origin.
+For the reader’s convenience, we restate (1.13) below and give a proof.
 Theorem 6.5.
-Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. Then
+Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
 $\frac{1}{\Gamma(\beta+1)^{1/\beta}}{\rm R}_{\beta}f\subset\frac{1}{\Gamma(\alpha+1)^{1/\alpha}}{\rm R}_{\alpha}f\subset 2\|f\|_{1}\Pi^{*}f$ (6.2)
-for $1<\alpha<\beta<\infty$.
-In each inclusion, equality holds if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ a simplex containing the origin.
+for $1<\alpha<\beta<\infty$. In each inclusion, there is equality if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ where $a=\|f\|_{\infty}$ and $\triangle$ is a simplex containing $x_{0}$.
 Proof.
 By Lemma 2.3,
-$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
-is an even, log-concave function and attains maximum at $t=0$.
-For fixed $u\in\mathbb{S}^{n-1}$, we apply Lemma 9.2 to $\omega(t)=g(0)e^{-t}$ and $\varphi(t)=-\log(g(tu)/g(0))$. We observe that
-$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)e^{-t}dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
+$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
+is an even, log-concave function and attains maximum at $r=0$.
+For fixed $u\in\mathbb{S}^{n-1}$, we apply Lemma 6.2 to $\omega(r)=g(o)e^{-r}$ and $\varphi(r)=-\log(g(ru)/g(o))$. We observe that
+$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}g(ru)dr}{\int_{0}^{\infty}r^{\alpha-1}g(o)e^{-r}dr}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
 for $\alpha>0$.
 For $\alpha\in(-1,0)$, note that
-$g(0)-g(tu)$ $=\|f\|_{1}-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ (9.3)
-$=\frac{1}{2}\int_{\mathbb{R}^{n}}f(x)dx+\frac{1}{2}\int_{rn}f(x+tu)dx-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
-$=\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x)-f(x+tu))_{+}dx+\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x+tu)-f(x))_{+}dx$
-$=\frac{1}{2}\int_{\mathbb{R}^{n}}|f(x)-f(x+tu)|dx,$
+$g(o)-g(ru)$ $=\|f\|_{1}-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ (6.3)
+$=\frac{1}{2}\int_{\mathbb{R}^{n}}f(x)dx+\frac{1}{2}\int_{\mathbb{R}^{n}}f(x+ru)dx-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
+$=\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x)-f(x+ru))_{+}dx+\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x+ru)-f(x))_{+}dx$
+$=\frac{1}{2}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dx,$
 where $h(x)_{+}$ denotes the positive part of the function $h(x)$. Then we have
-$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)(1-e^{-t})dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
+$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}(g(o)-g(ru))dr}{\int_{0}^{\infty}r^{\alpha-1}g(o)(1-e^{-r})dr}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
 for $\alpha\in(-1,0)$.
@@
 Letting $\alpha\rightarrow-1$, Lemma 5.6 gives the right inclusion.
-Since $\omega(t)=e^{-t}$ is strictly decreasing, by Lemma 9.3, equality holds in (9.4) if and only if $g(tu)=g(0)e^{-\lambda t}$ for some $\lambda>0$.
-Then Lemma 9.4 shows that $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ is a simplex containing the origin.
+Since $\omega(r)=e^{-r}$ is strictly decreasing, by Lemma 6.3, equality holds in (6.4) if and only if $g(ru)=g(o)e^{-\lambda r}$ for some $\lambda>0$.
+Then Lemma 6.4 shows that $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for $a=\|f\|_{\infty}$ and $x_{0}\in\triangle$, where $\triangle$ is a simplex containing the origin.
 ∎
-In fact, if we set $f(x)=e^{-\|x\|_{K}}$ for a convex body $K\subset\mathbb{R}^{n}$, then inclusion (9.2) in Theorem 9.5 reduces to the geometric case stated in Theorem 9.1.
-However, in this functional extension, the endpoint ${\rm R}_{\infty}f$ is $\mathbb{R}^{n}$, so the inclusion no longer extends to that limit case.
-Furthermore, combining Lemma 9.2 with (9.4), we obtain an explicit expression for $\zeta(0)$ by the continuity of $\zeta(\alpha)$, where
-$\log\zeta(0)=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt.$ (9.5)
-When $f=\operatorname{1}_{K}$, the quantity $\log\zeta(0)=\log\rho_{{\rm R}_{0}K}(u)$.
-The next result confirms that the formula (9.5) is consistent with the definition of ${\rm R}_{0}f$ (8.2), namely $\rho_{{\rm R}_{0}f}(u)=\zeta(0)$ with $u\in\mathbb{S}^{n-1}$ fixed.
-Proposition 9.6.
-For fixed $u\in\mathbb{S}^{n-1}$ and non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
-$\rho_{{\rm R}_{0}f}(u)=\zeta(0).$
-Proof.
-By (9.5),
-$\log\zeta(0)=$ $-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt$
-$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\int_{0}^{\|f\|_{\infty}}|\frac{\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}d\mu_{f}(r)-e^{-t}\Big)dt$
-$=\int_{0}^{\|f\|_{\infty}}\Big(-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt\Big)d\mu_{f}(r).$
-Note that
-$\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)$ $=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{|\{f\geq r\}|}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{\{f\geq r\}}(x),\operatorname{1}_{\{f\geq r\}}(x+tu)\}dx-e^{-t}\Big)dt$
-$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt,$
-where the first equation is from (9.5). Therefore, we have
-$\log\zeta(0)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r)=\log\rho_{{\rm R}_{0}f}(u),$
-which concludes the proof. ∎
+If we set $f(x)=e^{-\|x\|_{K}}$ for a convex body $K\subset\mathbb{R}^{n}$, then inclusion (6.2) in Theorem 6.5 reduces to the geometric case stated in Theorem 6.1.
+However, in this functional extension the endpoint ${\rm R}_{\infty}f$ is $\mathbb{R}^{n}$, so the inclusion no longer extends to that limit.
+This observation motivates the consideration of the embedding $K\mapsto\operatorname{1}_{K}$ in studying functional extensions of Theorem 6.1.
+The following counterexample shows, however, that this extension fails for general log-concave functions, which naturally leads to the study of $s$-concave functions in the next subsection.
+Example 6.6.
+Let $n=1$ and $f(x)=2x\operatorname{1}_{[0,1]}(x)$. It is clear that $\log f(x)=\log x+\log 2$ for $x\in[0,1]$, which is concave on ${\rm supp}(f)$. Then we have
+$g(r)=(1-|r|)_{+}^{2}.$
+In this case, we consider $\beta=1$ and $\alpha=\frac{1}{2}$ and obtain that
+$c_{1,1}{\rm R}_{1}f=\Big[-\frac{2}{3},\frac{2}{3}\Big],\quad c_{1,\frac{1}{2}}{\rm R}_{\frac{1}{2}}f=\Big[-\frac{16}{25},\frac{16}{{25}}\Big],$
+which is not consistent with the desired result $c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f$.
 6.2. Results for $s$-concave functions
-In this subsection, we focus on the case of $s$-concave functions. The following lemma is a direct corollary from the Borell-Brascamp-Lieb inequality, which is also mentioned in [15, Theorem 11.2].
+In this subsection, we focus on the case of $s$-concave functions with $s>0$. We continue to use the notation
+$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
+The following lemma is a direct corollary from the Borell-Brascamp-Lieb inequality, which is also mentioned in [17, Theorem 11.2].
 Lemma 6.7.
-For $s$-concave $f\in L^{1}(\mathbb{R}^{n})$ and fixed $u\in\mathbb{S}^{n-1}$,
-$t\mapsto g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu))\}dx$
-is a $s/(ns+1)$-concave function on $\mathbb{R}$.
+Let $s>0$ and $u\in\mathbb{S}^{n-1}$ be fixed. For an $s$-concave, integrable function $f$, the function $r\mapsto g(ru)$ is $s/(ns+1)$-concave on $\mathbb{R}$.
 Proof.
 We denote by
-$N_{f}(x,t)=\min\{f(x),f(x+tu)\}.$
-First, we prove that $N_{f}(x,t)$ is $s$-concave on $\mathbb{R}^{n}\times\mathbb{R}$.
-For $x_{0},x_{1}\in\mathbb{R}^{n}$, $t_{0},t_{1}\in\mathbb{R}$ and $\lambda\in[0,1]$, we denote $x_{\lambda}=(1-\lambda)x_{0}+\lambda x_{1}$ and $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$.
+$N_{f}(x,r)=\min\{f(x),f(x+ru)\}.$
+First, we prove that $N_{f}(x,r)$ is $s$-concave on $\mathbb{R}^{n}\times\mathbb{R}$.
+For $x_{0},x_{1}\in\mathbb{R}^{n}$, $r_{0},r_{1}\in\mathbb{R}$ and $\lambda\in[0,1]$, we denote $x_{\lambda}=(1-\lambda)x_{0}+\lambda x_{1}$ and $r_{\lambda}=(1-\lambda)r_{0}+\lambda r_{1}$.
 Note that
-$N_{f}(x_{\lambda},t_{\lambda})^{s}=\min\{f(x_{\lambda})^{s},f(x_{\lambda}+t_{\lambda}u)^{s}\}.$
+$N_{f}(x_{\lambda},r_{\lambda})^{s}=\min\{f(x_{\lambda})^{s},f(x_{\lambda}+r_{\lambda}u)^{s}\}.$
 Since $f$ is $s$-concave and $\min\{\cdot,\cdot\}$ is a concave function on $\mathbb{R}\times\mathbb{R}$, we have
-$N_{f}(x_{\lambda},t_{\lambda})^{s}$ $\geq\min\{(1-\lambda)f(x_{0})^{s}+\lambda f(x_{1})^{s},(1-\lambda)f(x_{0}+t_{0}u)^{s}+\lambda f(x_{1}+t_{1}u)^{s}\}$
-$\geq(1-\lambda)\min\{f(x_{0})^{s},f(x_{0}+t_{0}u)^{s}\}+\lambda\min\{f(x_{1})^{s},f(x_{1}+t_{1}u)^{s}\}$
-$=(1-\lambda)N_{f}(x_{0},t_{0})^{s}+\lambda N_{f}(x_{1},t_{1})^{s},$
+$N_{f}(x_{\lambda},r_{\lambda})^{s}$ $\geq\min\{(1-\lambda)f(x_{0})^{s}+\lambda f(x_{1})^{s},(1-\lambda)f(x_{0}+r_{0}u)^{s}+\lambda f(x_{1}+r_{1}u)^{s}\}$
+$\geq(1-\lambda)\min\{f(x_{0})^{s},f(x_{0}+r_{0}u)^{s}\}+\lambda\min\{f(x_{1})^{s},f(x_{1}+r_{1}u)^{s}\}$
+$=(1-\lambda)N_{f}(x_{0},r_{0})^{s}+\lambda N_{f}(x_{1},r_{1})^{s},$
 which implies the $s$-concavity of $N_{f}$.
-For $\lambda\in[0,1]$, and fixed $t_{0},t_{1}\in\mathbb{R}$, we set
-$F(x)=N_{f}(x,t_{0}),~~G(x)=N_{f}(x,t_{1}),~~H(x)=N_{f}(x,t_{\lambda}),$
-where $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$. By the $s$-concavity of $F$, we have
-$H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y)).$
-The Borell-Brascamp-Lieb inequality (2.3) then implies that
+For $\lambda\in[0,1]$, and fixed $r_{0},r_{1}\in\mathbb{R}$, we set
+$F(x)=N_{f}(x,r_{0}),~~G(x)=N_{f}(x,r_{1}),~~H(x)=N_{f}(x,r_{\lambda}),$
+where $r_{\lambda}=(1-\lambda)r_{0}+\lambda r_{1}$. By the $s$-concavity of $N_{f}$, we have
+$H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y)),$
+where $\mathcal{M}_{\lambda}^{s}$ denotes the $s$-mean (see Section 2.4). The Borell-Brascamp-Lieb inequality (2.4) then implies that
 $\int_{\mathbb{R}^{n}}H\geq\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big),$ (6.5)
-that is, the $\frac{s}{ns+1}$-concavity of $g(tu)$ in $t$. ∎
-We begin the proof of Theorem 5 by stating its inclusion result as the following theorem; the equality case will be addressed separately.
+that is, the $\frac{s}{ns+1}$-concavity of $g(ru)$ in $r$. ∎
+We begin the proof of Theorem 3 by stating the inclusion result as the following theorem, the equality case will be addressed separately.
 Theorem 6.8.
-Let $f\in L^{1}(\mathbb{R}^{n})$ be $s$-concave with compact support. If $1<\beta<\alpha$, then
+Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero $s$-concave function with compact support. If $1<\alpha<\beta$, then
 ${\rm R}_{\infty}f\subset c_{n,\beta}(s){\rm R}_{\beta}f\subset c_{n,\alpha}(s){\rm R}_{\alpha}f\subset(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$ (6.6)
@@
 Proof.
-By Lemma 9.7, the function $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ is $s^{\prime}$-concave with $s^{\prime}=s/(ns+1)$.
-We apply Lemma 9.2 with $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$.
+By Lemma 6.7, the function $g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ is $s^{\prime}$-concave with $s^{\prime}=s/(ns+1)$.
+We apply Lemma 6.2 with $\omega(r)=g(0)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(0))^{s^{\prime}})$.
 Therefore, we obtain that
-$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}(1-s^{\prime}t)_{+}^{1/s^{\prime}}dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
+$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}g(ru)dr}{\int_{0}^{\infty}g(0)r^{\alpha-1}(1-s^{\prime}r)_{+}^{1/s^{\prime}}dr}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
 for $\alpha>0$ and
-$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}((1-s^{\prime}t)_{+}^{1/s^{\prime}}-1)dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
+$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}(g(0)-g(ru))dr}{\int_{0}^{\infty}g(0)r^{\alpha-1}((1-s^{\prime}r)_{+}^{1/s^{\prime}}-1)dr}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
 for $1<\alpha<0$, where we use (6.1) and (6.3). Note that
-$\Big(\alpha B(\alpha,1+\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=\Big(nB(\alpha+1,\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=c_{n,\alpha}(s).$
+$\Big(\alpha B(\alpha,1+\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=\Big(\frac{1}{s^{\prime}}B(\alpha+1,\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=c_{n,\alpha}(s).$
 Lemma 6.2 then implies that
@@
 for $1<\alpha<\beta$.
-Since Lemma 8.3 ensures that ${\rm R}_{\beta}f\to{\rm R}_{\infty}f$ and $c_{n,\beta}(s)\to 1$ as $\beta\rightarrow\infty$, the left inclusion follows directly from (9.8).
+Since Lemma 5.7 ensures that ${\rm R}_{\beta}f\to{\rm R}_{\infty}f$ and $c_{n,\beta}(s)\to 1$ as $\beta\rightarrow\infty$, the left-most inclusion follows directly from (6.7) by letting $\beta\rightarrow\infty$.
 By Lemma 5.6 and $c_{n,\alpha}(s)\sim\frac{\alpha+1}{n+1/s}$ when $\alpha\rightarrow-1$, we have
-$\lim_{\alpha\rightarrow-1^{+}}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=(n+\frac{1}{s})\|f\|_{1}\rho_{\Pi^{*}f}(u).$
-which yields the right inclusion. ∎
-When $s=\infty$, Theorem 9.8 reduces to the geometric case in Theorem 9.1 by Gardner and Zhang.
-In fact, following the argument of our proof, for a log-concave function $f\in L^{1}(\mathbb{R}^{n})$ with compact support, if one could show that $g(tu)^{1/n}$ is concave, then the following chain of inclusions would hold:
-${\rm R}_{\infty}f\subset c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f\subset n\|f\|_{1}\Pi^{*}f,$
-for $1<\alpha<\beta<\infty$.
-This would yield precisely the functional counterpart of Theorem 9.1 that one would expect.
-However, $g(tu)^{1/n}$ is not necessarily concave for a general log-concave function $f$.
-A counterexample will be provided below.
-Let $n=1$ and
-[…]
-It is clear that $\log f(x)=\log x+\log 2$ for $t\in[0,1]$, which is concave. Then we have
-[…]
-Note that
-$\frac{1}{4}=g(\frac{1}{2})<\frac{1}{2}(g(0)+g(1))=\frac{1}{2},$
-which shows that $g$ is not concave. In this case, we consider $\beta=1$ and $\alpha=1/2$ and obtain that
-$c_{1,1}{\rm R}_{1}f=\Big[-\frac{2}{3},\frac{2}{3}\Big],\quad c_{1,\frac{1}{2}}{\rm R}_{\frac{1}{2}}f=\Big[-\frac{16}{25},\frac{16}{{25}}\Big].$
-which is not consistent with the desired result $c_{n,\beta}{\rm R}_{\beta}f\subset c_{n,\alpha}{\rm R}_{\alpha}f$.
-The following theorem describes the sharpness of the inclusion result in Theorem 5.
+$\lim_{\alpha\rightarrow-1^{+}}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=(n+\frac{1}{s})\|f\|_{1}\rho_{\Pi^{*}f}(u),$
+which yields the right-most inclusion. ∎
+The following theorem demonstrates the sharpness of (6.6).
 Theorem 6.9.
-Equality holds in (9.7) if $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $x_{0}\in\mathbb{R}^{n}$, $c>0$ and $\triangle$ is a simplex containing the origin.
+Let $s>0$. If $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $a=\|f\|_{\infty}>0$ and $\triangle\subset\mathbb{R}^{n}$ is a simplex containing $x_{0}$, then
+${\rm R}_{\infty}f=c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f=(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$
+where $1<\alpha<\beta<\infty$ and $c_{n,\alpha}(s)$ is given in Theorem 6.8.
 Proof.
-Note that ${\rm R}_{\alpha}f$ is invariant under scaling and translation of $f$. Without loss of generality, we may assume $f(x)=(1-\|x\|_{\triangle})_{+}^{1/s}$. A direct computation then gives
-$\|f\|_{1}=B(\frac{1}{s},n+1)\frac{|\triangle|}{s}.$
-We use the same notation for $g(tu)$ with a fixed $u\in\mathbb{S}^{n-1}$, and we have
-$g(tu)$ $=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
-$=\int_{0}^{1}|\{f\geq r\}\cap(\{f\geq r\}+tu)|dr$
-$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}|(1-r)\triangle\cap((1-r)\triangle+tu)|dr$
-$=\int_{0}^{1}\frac{1}{s}r^{\frac{1}{s}-1}(1-r)^{n}|\triangle\cap(\triangle+\frac{t}{1-r}u)|dr.$
-By [18, Lemma 3.4], $|\triangle\cap(\triangle+tu)|^{1/n}$ is linear in $t$. Therefore, there exists $\lambda>0$ such that
-$|\triangle\cap(\triangle+\frac{t}{1-r})u|=|\triangle|(1-\lambda\frac{t}{1-r})_{+}^{n}.$
-A straightforward, though somewhat tedious, calculation then shows that
-$g(tu)=\frac{1}{s}B(\frac{1}{s},n+1)|\triangle|(1-\lambda t)^{\frac{ns+1}{s}}.$
+By Lemma 5.3, ${\rm R}_{\alpha}f$ is invariant under scaling and translation. We may assume, without loss of generality, that $x_{0}=o$ and $a=1$. Then
+$\|f\|_{1}=\int_{0}^{1}|\{f\geq t\}|dt=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}|\{f^{s}\geq t\}|dt,$
+where $\{f^{s}\geq t\}=(1-t)\triangle$. Hence
+$\|f\|_{1}=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}(1-t)^{n}|\triangle|dt=B(\frac{1}{s},n+1)\frac{|\triangle|}{s}.$
+Moreover, we obtain
+$g(ru)$ $=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
+$=\int_{0}^{1}|\{f\geq t\}\cap(\{f\geq t\}+ru)|dt$
+$=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}|(1-t)\triangle\cap((1-t)\triangle+ru)|dt$
+$=\int_{0}^{1}\frac{1}{s}t^{\frac{1}{s}-1}(1-t)^{n}\Big|\triangle\cap(\triangle+\frac{r}{1-t}u)\Big|dt.$
+By [18, Lemma 3.4], $|\triangle\cap(\triangle+ru)|^{1/n}$ is an affine function with respect to $r>0$. Therefore, there exists $\lambda>0$ such that
+$\Big|\triangle\cap(\triangle+\frac{r}{1-t})u\Big|=|\triangle|\Big(1-\lambda\frac{r}{1-t}\Big)_{+}^{n},$
+where $\lambda$ depends on $u$.
+A straightforward calculation shows that
+$g(ru)=\frac{1}{s}B(\frac{1}{s},n+1)|\triangle|(1-\lambda r)_{+}^{\frac{ns+1}{s}}.$
 We set $s^{\prime}=s/(ns+1)$.
-Then, for $1<\alpha<0$ and $\alpha>0$, by taking $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
-$c_{n,\alpha}(s){\rm R}_{\alpha}f=\frac{1}{s^{\prime}}\zeta(\alpha)=\frac{1}{s^{\prime}}\Big(\frac{\int_{0}^{\infty}t^{\alpha-1}(1-\lambda t)_{+}^{1/s^{\prime}}dt}{\int_{0}^{\infty}t^{\alpha-1}(1-s^{\prime}t)^{1/s^{\prime}}dt}\Big)^{1/\alpha}=\frac{1}{\lambda},$
-where the first equality is from thex calculation in Theorem 9.8.
+For $1<\alpha<0$ and $\alpha>0$, by taking $\omega(r)=g(o)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(o))^{s^{\prime}})$ in Lemma 6.2, we obtain
+$c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u)=\frac{1}{s^{\prime}}\zeta(\alpha)=\frac{1}{s^{\prime}}\Big(\frac{\int_{0}^{\infty}r^{\alpha-1}(1-\lambda r)_{+}^{1/s^{\prime}}dr}{\int_{0}^{\infty}r^{\alpha-1}(1-s^{\prime}r)_{+}^{1/s^{\prime}}dr}\Big)^{1/\alpha}=\frac{1}{\lambda},$
+where the first equality is from the calculation in the proof of Theorem 6.8.
 This implies that
 $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,\quad\quad-1<\alpha<\beta<\infty.$
-Letting $\alpha\rightarrow-1$ and $\beta\rightarrow\infty$, we have
-${\rm R}_{\infty}f=c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f=(n+\frac{1}{s})\|f\|_{1}\Pi^{*}f,$
+Letting $\alpha\rightarrow-1$ and $\beta\rightarrow\infty$ completes the proof. ∎
+To establish the equality characterization, we first need an auxiliary lemma based on the equality case of the Borell–Brascamp–Lieb inequality (2.4).
+Before stating this result, we collect some preliminary facts that will be used later in the argument.
+Throughout, ${\rm cl}K$ denotes the closure of a set $K$.
+Lemma 6.10.
+Let $f:\mathbb{R}\rightarrow[0,\infty)$ be a concave function on $(0,c)$ with ${\rm supp}(f)=[0,c]$, and define
+$K={\rm cl}\{(x,y)\in\mathbb{R}^{2}:x\in(0,c)~\text{and}~0\leq y\leq f(x)\}.$
+Let $e_{1}=(1,0)\in\mathbb{R}^{2}$. For all $r\in(0,c)$, if either
+$K\cap(K-re_{1})=(1-\tfrac{r}{c})K\quad\text{or}\quad K\cap(K+re_{1})=(1-\tfrac{r}{c})K+re_{1},$
+then $K$ is a triangle.
+Proof.
+We assume that $K\cap(K-re_{1})=(1-\frac{r}{c})K$ and prove the statement, the other case is similar. Let
+$x_{0}=\inf\{x\in[0,c]:f(x)=\|f\|_{\infty}\}~\text{and}~x_{1}=\sup\{x\in[0,c]:f(x)=\|f\|_{\infty}\}.$
+We first aim to show that $x_{0}=x_{1}$.
+If $x_{1}>x_{0}$, we choose $r>0$ to be sufficiently small such that $K\cap(K-re_{1})\neq\emptyset$ and $r<x_{1}-x_{0}$.
+Then the identity $K\cap(K-re_{1})=(1-\frac{r}{c})K$ implies that
+$x_{1}-x_{0}-r=(1-\tfrac{r}{c})(x_{1}-x_{0}).$
+It follows that either $x_{1}-x_{0}=c$ or $x_{1}-x_{0}=0$.
+Since the case $x_{1}-x_{0}=c$ would imply that $K$ is a rectangle, contradicting the assumption $K\cap(K-re_{1})=(1-\frac{r}{c})K$, we conclude that $x_{0}=x_{1}$.
+Next we show that $f(0)=0$ if $x_{0}\neq 0$.
+Since $f$ is concave and hence continuous on $(0,c)$, we have $f(0^{-})=\lim_{x\rightarrow 0^{-}}f(x)=0$ and $f(0^{+})=\lim_{x\rightarrow 0^{+}}f(x)\geq f(0)\geq 0$.
+Thus, it suffices to show that $f(0^{+})=0$.
+Since $x_{0}$ is the unique maximizer, $f$ is monotone increasing on $(0,x_{0})$, and therefore $f(r)\geq f(0^{+})$ for $r>0$.
+Then
+$\big\{(0,y):y\in[0,f(0^{+})]\big\}\subset K\cap(K-re_{1})=(1-\tfrac{r}{c})K,$
+with $r$ sufficiently small. Consequently, $f(0^{+})=(1-\frac{r}{c})f(0^{+})$, which implies that $f(0)=0$. Similarly, one obtains $f(c)=0$ provided that $x_{0}\neq c$.
+We now prove that $K$ is a triangle provided that $x_{0}\in(0,c)$.
+In this case, we have $f(0)=f(c)=0$.
+Since $x_{0}$ is the unique maximizer, $f$ is monotone increasing on $(0,x_{0})$ and monotone decreasing on $(x_{0},c)$.
+The identity $K\cap(K-re_{1})=(1-\frac{r}{c})K$ then implies that
+$f\big((1-\tfrac{r}{c})x_{0}\big)=(1-\tfrac{r}{c})f(x_{0}),$
+and
+$f\big((1-\tfrac{r}{c})x_{0}+r\big)=(1-\tfrac{r}{c})f(x_{0}).$
+This shows that $f$ is linear on $0,x_{0}]$ and affine on $x_{0},c]$, which implies that $K$ is a triangle. For $x_{0}=0$ or $x_{0}=c$, the proof is similar. ∎
+Next we aim to extend Lemma 6.10 to concave functions on $\mathbb{R}^{n}$. For a non-negative, concave function $f\in L^{1}(\mathbb{R}^{n})$, we denote by
+${\rm H}(f)={\rm cl}\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}~{\rm supp}(f)~\text{and}~0\leq l\leq f(x)\}$ (6.8)
+the hypograph of $f$.
+Here ${\rm int}K$ denotes the interior of a set $K$.
+Let $e_{1},\ldots,e_{n}$ be an orthonormal basis of $\mathbb{R}^{n}$.
+To simplify the notation, we only consider the function whose support is the standard simplex defined by
+$\triangle_{n}=\Big\{(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:x_{i}\geq 0~\text{for}~i=1,\ldots,n,~\text{and}~\sum_{i=1}^{n}x_{n}\leq 1\Big\}.$ (6.9)
+Recall that $a^{+}=\max\{a,0\}$. Here we also denote $a^{-}=\max\{-a,0\}$.
+For $u=(u_{1},\ldots,u_{n})\in\mathbb{S}^{n-1}$, we denote by $u^{+}=(u_{1}^{+},\ldots,u_{n}^{+})$ and $u^{-}=(u_{1}^{-},\ldots,u_{n}^{-})$. Moreover, we set
+$a_{0}=\sum_{i=1}^{n}u_{i}^{+},~b_{0}=\sum_{i=1}^{n}u_{i}^{-},~\text{and}~c_{0}=\max\{a_{0},b_{0}\}.$ (6.10)
+One can verify geometrically that
+$\triangle_{n}\cap(\triangle_{n}+ru)=(1-c_{0}r)\triangle_{n}+ru^{+},$
+for any $u\in\mathbb{S}^{n-1}$.
+The next result extends Lemma 6.10 to higher dimensional cases.
+Since we identify $\mathbb{R}^{n}$ with the hyperplane $\mathbb{R}^{n}\times\{0\}\subset\mathbb{R}^{n+1}$, we sometimes write $x\in\mathbb{R}^{n}$ also for the corresponding point in $\mathbb{R}^{n+1}$, instead of $(x,0)$.
+Lemma 6.11.
+Let $u\in\mathbb{S}^{n-1}$ be fixed, and $c_{0}$ be defined in (6.10).
+Suppose $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ is concave and ${\rm supp}(f)=\triangle_{n}$, which is the standard simplex given by (6.9).
+For all $r\in(0,\frac{1}{c_{0}})$ and $u\in\mathbb{S}^{n-1}$, if
+${\rm H}(f)\cap\big({\rm H}(f)+ru\big)=(1-c_{0}r){\rm H}(f)+ru^{+},$ (6.11)
+then $f(x)=a(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+}$ with $a=\|f\|_{\infty}$ and $x_{0}\in\triangle_{n}$.
+Proof.
+We may assume $\|f\|_{\infty}=1$ without loss of generality.
+We aim to show that $f$ is linear on $o,x_{0}]$ and affine on $x_{0},e_{i}]$ for all $i\in\{1,\ldots,n\}$, where $[x,y]$ denotes the line segment between $x$ and $y$.
+The case where $x_{0}=o$ or $x_{0}=e_{i}$ is analogous.
+We first consider $f$ on $o,x_{0}]$.
+Let $u=-\frac{x_{0}}{|x_{0}|}$ and thus we have $u^{+}=o$.
+We choose the $2$-dimensional subspace $T$ spanned by $u$ and $(0,\ldots,0,1)\in\mathbb{R}^{n+1}$.
+Then (6.11) implies that
+$K\cap(K+ru)=(1-c_{0}r)K,$
+where $K={\rm H}(f)\cap T$. By Lemma 6.10, we obtain the linearity of $f$ on $o,x_{0}]$.
+Next we consider $f$ on $x_{0},e_{i}]$ for any fixed $i\in\{1,\ldots,n\}$. In this case, we let $u=\frac{x_{0}-e_{i}}{|x_{0}-e_{i}|}$ and then
+$u^{+}=\frac{1}{|e_{i}-x_{0}|}(x_{0}^{(1)},\ldots,x_{0}^{(i-1)},0,x_{0}^{(i+1)},\ldots,x_{0}^{(n)})~\text{and}~c_{0}=\frac{1-x_{0}^{(i)}}{|x_{0}-e_{i}|}.$
+We choose the $2$-dimensional subspace $T$ spanned by $u$ and $(x_{0},1)$. Restricting (6.11) to $T$, we obtain
+$\big({\rm H}(f)\cap T\big)\cap\big(({\rm H}(f)\cap T)+ru\big)=(1-c_{0}r){\rm H}(f)\cap T+ru^{+}.$ (6.12)
+Set $K={\rm H}(f)\cap T\subset T$. The factor $(1-c_{0}r)$ in (6.12) corresponds to a dilation in $T$, and hence
+$K\cap(K+ru)=(1-c_{0}r)\cdot K+ru.$
+Here, $(1-c_{0}r)\cdot K$ denotes the dilation of $K$ in $T$ with respect to a chosen point $e_{i}$ in $T$, which serves as the origin for the scaling in the affine subspace $T$.
+By Lemma 6.10, $f$ is affine on $x_{0},e_{i}]$.
+Therefore, we have that ${\rm H}(f)$ is a simplex in $\mathbb{R}^{n+1}$ with vertices $o,e_{1},\ldots,e_{n},(x_{0},1)$ and attains 0 on the boundary point, that is,
+$f(x)=(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+},$
 which concludes the proof. ∎
-We conclude the section by showing the necessary condition for the equality cases in Theorem 5. The following lemma is the key result for the equality characterization.
+We also need the following basic, but non-trivial result.
 Lemma 6.12.
-Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ non-zero and $s$-concave with compact support. If
-$g(tu)^{\frac{s}{ns+1}}=\Big(\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx\Big)^{\frac{s}{ns+1}}$
-is linear in $t$ for any fixed $u\in\mathbb{S}^{n-1}$, then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$, where $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
+Suppose $A,B\subset\mathbb{R}^{n}$ are convex sets with ${\rm int}A\cap{\rm int}B\neq\emptyset$. Then ${\rm cl}(A\cap B)={\rm cl}A\cap{\rm cl}B$.
 Proof.
-We assume, without loss of generality, that $\|f\|_{\infty}=1$. Since
-$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x-tu)\}dx,$
-for convenience of notations in the following proof, we set $N_{f}(x,t)=\min\{f(x),f(x-tu)\}$ here. The functions $F,G,H$ are as defined by
-$F(x)=N_{f}(x,0),~~G(x)=N_{f}(x,t),~~H(x)=N_{f}(x,\lambda t),$
-which satisfies that $H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y))$.
-The linearity of $g(tu)^{\frac{s}{ns+1}}$ then implies that
+Since $A\cap B\subset A$ and $A\cap B\subset B$, taking closures yields ${\rm cl}(A\cap B)\subset{\rm cl}A\cap{\rm cl}B$.
+Let $x\in{\rm cl}A\cap{\rm cl}B$.
+We choose $x_{0}\in{\rm int}A\cap{\rm int}B$ and let
+$x_{\lambda}=(1-\lambda)x_{0}+\lambda x,~\text{for}~\lambda\in[0,1).$
+By [42, Lemma 1.1.9], we have $x_{\lambda}\in{\rm int}A\cap{\rm int}B$.
+Since $x_{\lambda}\rightarrow x$ as $\lambda\rightarrow 1^{-}$, we obtain $x\in{\rm cl}({\rm int}A\cap{\rm int}B)\subset{\rm cl}(A\cap B)$.
+∎
+The following result associates the equality characterization of the Borell-Brascamp-Lieb inequality with (6.11) in Lemma 6.11.
+Lemma 6.13.
+Let $u\in\mathbb{S}^{n-1}$ be fixed, and $c_{0}$ be defined in (6.10). Suppose $f$ is a non-negative, non-zero, concave function on $\mathbb{R}^{n}$ with ${\rm supp}(f)=\triangle_{n}$. If
+$\min\{f((1-c_{0}r)x+ru^{+}),f((1-c_{0}r)x+ru^{-})\}=(1-c_{0}r)f(x),$ (6.13)
+for all $x\in{\rm int}\triangle_{n}$, then
+${\rm H}(f)\cap\big({\rm H}(f)+ru\big)=(1-c_{0}r){\rm H}(f)+ru^{+},$
+for $r\in(0,\frac{1}{c_{0}})$, where ${\rm H}(f)$ denotes the hypograph of $f$ as defined in (6.8).
+Proof.
+Let $c_{r}=1-c_{0}r$. For clarity of notation, we denote by
+$P=\{(x,l)\in{\mathbb{R}^{n+1}}:x\in{\rm int}\triangle_{n},~0\leq l\leq f(x)\},$
+and $m(x)=\min\{f((1-c_{0}r)x+ru^{+}),f((1-c_{0}r)x+ru^{-})\}$. Then ${\rm H}(f)={\rm cl}P$.
+We first have
+$\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq c_{r}f(x)\}=\phi_{r}P,$ (6.14)
+where $\phi_{r}={\rm diag}(1,\ldots,1,c_{r})\in{\rm GL}(n+1)$.
+To consider the hypograph of $m(x)$, we first denote
+$P_{1}$ $=\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq f((c_{r}x+ru^{+})\}$ (6.15)
+$~=\psi_{r}(\{(x-ru^{+},l):x\in{\rm int}\triangle_{n},~0\leq l\leq f(x)\})$
+$~=\psi_{r}(P-ru^{+}),$
+where $\psi_{r}={\rm diag}(c_{r}^{-1},\ldots,c_{r}^{-1},1)$. Similarly, we denote
+$P_{2}=\{(x,l)\in\mathbb{R}^{n+1}:x\in{\rm int}\triangle_{n},~0\leq l\leq f((c_{r}x+ru^{-})\}=\psi_{r}(P-ru^{-}).$ (6.16)
+Note that $\psi_{r}=c_{r}^{-1}\phi_{r}$ and $(A-x)\cap(A-y)=A\cap(A+(x-y))-x$ holds for any set AA and $x,y$. Combining this with (6.14), (6.15), (6.16) and the assumption (6.13), we obtain
+$P\cap(P+ru)=c_{r}P+ru^{+}.$ (6.17)
+Recall that $r\in(0,\frac{1}{c_{0}})$, which yields that ${\rm int}(P+ru)\cap{\rm int}P\neq\emptyset$. Then Lemma 6.12 implies that
+${\rm cl}\big(P\cap(P+ru)\big)={\rm cl}P\cap({\rm cl}P+ru).$
+Since ${\rm cl}P={\rm H}(f)$, we conclude the proof by taking closures on both sides of (6.17). ∎
+Combining Lemma 6.13 with Lemma 6.11, we obtain the following crucial corollary.
+Corollary 6.14.
+For $u\in\mathbb{S}^{n-1}$, let $c_{0}$ be defined in (6.10).
+Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a non-zero, concave function with ${\rm supp}(f)=\triangle_{n}$.
+For $r\in(0,\frac{1}{c_{0}})$, if
+$\min\{f((1-c_{0}r)x),f((1-c_{0}r)x+ru)\}=(1-c_{0}r)f(x),$
+for all $x\in{\rm int}\triangle_{n}$ and $u\in\mathbb{S}^{n-1}$, then $f(x)=a(1-\|x-x_{0}\|_{\triangle_{n}-x_{0}})_{+}$ with $a=\|f\|_{\infty}$ and $x_{0}\in\triangle_{n}$.
+We are now in a position to prove the following key result for the equality characterization in Theorem 3.
+Lemma 6.15.
+Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, $s$-concave function with compact support. For $u\in\mathbb{S}^{n-1}$, let
+$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
+If there is a constant $\lambda>0$ such that $g(ru)=\|f\|_{1}(1-\lambda r)_{+}^{\frac{ns+1}{s}}$ for $r>0$, then $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
+Proof.
+We assume $\|f\|_{\infty}=1$ without loss of generality. Since g(ru)=g(ru)g(ru)=g(-ru), we set
+$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x-ru)\}dx,$
+and $N_{f}(x,r)=\min\{f(x),f(x-ru)\}$ to simplify notation. Let $r\geq 0$ and $\lambda\in[0,1]$ be fixed, and define the functions $F,G,H:\mathbb{R}^{n}\to\mathbb{R}$ by
+$F(x)=N_{f}(x,0)=f(x),~~G(x)=N_{f}(x,r),~~H(x)=N_{f}(x,\lambda r),$
+which then satisfy
+$H((1-\lambda)x+\lambda y)\geq\Big((1-\lambda)F(x)^{s}+\lambda G(y)^{s}\Big)^{1/s}=\mathcal{M}_{\lambda}^{s}(F(x),G(y))$
+for all $x,y\in\mathbb{R}^{n}$ .
+Since $r\mapsto g(ru)^{\frac{s}{ns+1}}$ is an affine function with respect to $r$, then
 $\int_{\mathbb{R}^{n}}H=\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(\int_{\mathbb{R}^{n}}F,\int_{\mathbb{R}^{n}}G\Big).$
-That is, we have the equality in the Borell-Brascamp-Lieb inequality for $F,G,H$.
-By Theorem 2.4, there exists $c_{0}=\Big(\frac{|{\rm supp}(G)|}{|{\rm supp}(F)|}\Big)^{1/n}$ and $y_{0}\in\mathbb{R}^{n}$ such that
-${\rm supp}(G)=c_{0}\,{\rm supp}(F)+y_{0},$ (9.9)
-while the definition of $G$ shows
-${\rm supp}(G)={\rm supp}(F)\cap({\rm supp}(F)+tu).$ (9.10)
-Since $u\in\mathbb{S}^{n-1}$ is arbitrary fixed, (9.9) and (9.10) shows that ${\rm supp}(F)\cap({\rm supp}(F)+tu)$ is homothetic to ${\rm supp}(F)$ for all $u\in\mathbb{S}^{n-1}$.
-By [18, Lemma 3.4], we have ${\rm supp}(F)$ is a simplex $\triangle$, where we assume that $|\triangle|$ is the same as the volume of the standard simplex $\triangle_{n}$ for convenience.
-In fact, we only need to consider when $\triangle$ is the standard simplex, that is,
+Hence, equality holds in the Borell-Brascamp-Lieb inequality (2.4) for $F,G,H$.
+By Theorem 2.4, there exists $c_{r}=\Big(\frac{|{\rm supp}(G)|}{|{\rm supp}(f)|}\Big)^{1/n}$ and $y_{r}\in\mathbb{R}^{n}$ such that
+${\rm supp}(G)=c_{r}\,{\rm supp}(f)+y_{r},$ (6.18)
+while the definition of $G$ gives
+${\rm supp}(G)={\rm supp}(f)\cap({\rm supp}(f)+ru).$ (6.19)
+Since $u\in\mathbb{S}^{n-1}$ is arbitrarily fixed, (6.18) and (6.19) imply that ${\rm supp}(f)\cap({\rm supp}(f)+ru)$ is homothetic to ${\rm supp}(f)$ for all $u\in\mathbb{S}^{n-1}$.
+By [18, Lemma 3.4], ${\rm supp}(f)$ is a simplex denoted by $\triangle$.
+By Lemma 5.3, ${\rm R}_{\alpha}f$ is invariant under translations and ${\rm R}_{\alpha}(f\circ\phi^{-1})=\phi{\rm R}_{\alpha}f$ for all $\phi\in{\rm GL}(n)$.
+It suffices to consider the case where $\triangle$ is the standard simplex, that is,
 $\triangle=\triangle_{n}=\Big\{x=(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:x_{i}\geq 0,~~\sum_{i=1}^{n}x_{i}\leq 1\Big\}.$
-Otherwise, there exists $\phi\in{\rm SL}(n)$ such that ${\rm supp}(f\circ\phi^{-1})=\triangle_{n}$ and we consider $f\circ\phi^{-1}$ instead of $f$. In this case, by (9.9) and (9.10), we have
-$c_{0}(t)=1-t\sum_{i=1}^{n}u_{i},~~\text{and}~~y_{0}(t)=t(u_{1}^{+},\ldots,u_{n}^{+}),$
-where $u_{i}^{+}=\max\{0,u_{i}\}$ and we denote by $a_{0}=\sum_{i=1}^{n}u_{i}$. Therefore, we can choose $u\in\mathbb{S}^{n-1}$ with $u_{i}\geq 0$, such that $y_{0}(t)=tu$.
-Moreover, by Theorem 2.4, there exists a $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$ with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$ satisfying,
-$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{0}x+tu)=c_{0}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda tu)=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{0}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&.\end{aligned}\right$.
-In fact, $\Phi(x)=F(x)=f(x)$ by the definition of $F$. Then we have
-$G((1-a_{0}t)x+tu)^{s}=\min\{f((1-a_{0}t)x)^{s},f((1-a_{0}t)x+tu)^{s}\}=(1-a_{0}t)f(x)^{s}.$
-Note that $(1-a_{0}t)f(x)^{s}\neq f((1-a_{0}t)x)^{s}$ since $f^{s}$ is concave. Therefore,
-$f((1-a_{o}t)x+tu)^{s}=(1-a_{0}t)f(x)^{s}$ (9.11)
-holds for $0\leq t\leq 1/\sum u_{i}$. Equation (9.11), together with $f\geq 0$ and the $s$-concavity of $f$ , imply that
-$f\Big(\frac{u}{a_{0}}\Big)^{s}=0.$ (9.12)
-We set $\bar{x}=(\bar{x}_{1},\ldots,\bar{x}_{n})\in\triangle_{n}$ such that $f(\bar{x})=\|f\|_{\infty}=1$. Note that if $\sum_{i}\bar{x}_{i}\in(0,1)$, we will have
-$f\Big(\frac{\bar{x}}{\sum_{i}\bar{x}_{i}}\Big)^{s}=0\geq(1-\sum_{i}\bar{x}_{i})f(\bar{x})^{s}+\Big(\sum_{i}\bar{x}_{i}\Big)f(o)^{s}\geq 1-\sum_{i}\bar{x}_{i}>0,$ (9.13)
-which cause a contradiction.
-Thus we have $\sum_{i}\bar{x}_{i}=1$ or $\bar{x}=o$.
-If $\sum_{i}\bar{x}_{i}=1$, equation (9.12) implies that $f(\bar{x})=0$ which contradicts to the choice of $\bar{x}$.
-Thus we have $\bar{x}=o$.
-Letting $x=\bar{x}=o$ in (9.11) then gives
-$f(tu)^{s}=1-a_{0}t=1-\|tu\|_{\triangle_{n}}.$
-Since ${\rm supp}(f)=\triangle_{n}$, we then have
-$f(x)=(1-\|x\|_{\triangle_{n}})_{+}^{1/s},$
-which concludes the proof. ∎
-We conclude this section by the following theorem, which provides the full equality characterization in Theorem 5.
+By (6.18) and (6.19), we have
+$c_{r}=1-c_{0}r,~~\text{and}~~y_{r}=ru^{+},$
+where $c_{0}$ is defined by (6.10).
+Moreover, by Theorem 2.4, there exists a $(\lambda_{0},s)$-concave function $\Phi:{\rm supp}(F)\rightarrow[0,\infty)$, with $\lambda_{0}=\frac{\lambda c_{0}}{1-\lambda+\lambda c_{0}}$, satisfying
+$\left\{\begin{aligned} &F(x)=\Phi(x)&&,\\ &G(c_{r}x+y_{r})=c_{r}^{\frac{1}{s}}\Phi(x)&&,\\ &H((1-\lambda+\lambda c_{0})x+\lambda y_{r})=\Big[\mathcal{M}_{\lambda}^{\frac{s}{ns+1}}\Big(1,c_{r}^{\frac{ns+1}{s}}\Big)\Big]^{\frac{1}{ns+1}}\Phi(x)&&,\end{aligned}\right$.
+for almost all $x\in\triangle_{n}$. By the definition of $F$, we have $\Phi(x)=F(x)=f(x)$, and
+$G((1-c_{0}r)x+y_{r})^{s}=\min\{f((1-c_{0}r)x+ru^{-})^{s},f((1-c_{0}r)x+ru^{+})^{s}\}=(1-c_{0}r)f(x)^{s}.$ (6.20)
+Since $f^{s}$ is concave and hence continuous on ${\rm int}\triangle_{n}$, the equality (6.20) holds for all $x\in{\rm int}\triangle_{n}$. By applying Corollary 6.14, we complete the proof. ∎
+As a consequence, Lemma 6.15 derives the following theorem, which provides the full equality characterization in Theorem 3.
 Theorem 6.16.
-Let $s>0$ and $c_{n,\alpha}(s)$ be defined in Theorem 5. Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, and $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, if
+Let $s>0$ and $c_{n,\alpha}(s)$ be as defined in Theorem 3. Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. For $1<\alpha<\beta<\infty$, if
 $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f,$
-then $f(x)=c(1-\|x-x_{0}\|_{\triangle})_{+}^{1/s}$ where $c>0$, $x_{0}\in\mathbb{R}^{n}$ and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing the origin.
+then $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
 Proof.
-Let $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ with fixed $u\in\mathbb{S}^{n-1}$.
-By Theorem 9.8, letting $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
+Let $g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ for a fixed $u\in\mathbb{S}^{n-1}$ and $r\in\mathbb{R}$.
+Setting $\omega(r)=g(o)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(o))^{s^{\prime}})$ in Lemma 6.2, we obtain
 $\zeta(\alpha)=s^{\prime}c_{n,\alpha}(s)\rho_{{\rm R}_{\alpha}f}(u),$
 where $s^{\prime}=s/(ns+1)$.
-Then $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f$ implies that $\zeta(\alpha)=\zeta(\beta)$.
-Since $\omega$ is strictly decreasing on its support, by Lemma 9.3, we obtain $\lambda=\zeta(\alpha)^{-1}$ such that
-$\varphi(t)=\frac{1}{s^{\prime}}\Big(1-\Big(\frac{g(tu)}{g(0)}\Big)^{s^{\prime}}\Big)=\lambda t.$
-This implies that $g(tu)^{s^{\prime}}$ is linear in $t$. Then we conclude the proof by Lemma 9.10. ∎
+Then, the equality $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f$ implies that $\zeta(\alpha)=\zeta(\beta)$.
+Since $\omega$ is strictly decreasing on its support, Lemma 6.3 yields $\lambda=\zeta(\alpha)^{-1}$ such that
+$\varphi(r)=\frac{1}{s^{\prime}}\Big(1-\Big(\frac{g(ru)}{g(o)}\Big)^{s^{\prime}}\Big)=\lambda t.$
+This shows that $g(ru)^{s^{\prime}}$ is affine in $r$, and the conclusion follows from Lemma 6.15. ∎
+We conclude this section by showing the Rogers-Shephard inequality and Zhang inequality for $s$-concave functions, which are special cases of Theorem 3.
+Theorem 6.17 (Rogers-Shephard inequality for $s$-concave functions).
+Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. Then
+$|{\rm D}\,{\rm supp}(f)|\leq\frac{\Pi_{k=1}^{n}(k+\frac{1}{s})}{n\cdot n!\|f\|_{1}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy.$
+Equality holds if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
+Theorem 6.18 (Zhang’s inequality for $s$-concave functions).
+Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. Then
+$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq\frac{n\cdot n!(n+\frac{1}{s})^{n}}{\Pi_{k=1}^{n}(k+\frac{1}{s})}\|f\|_{1}^{n+1}|\Pi^{*}f|.$
+Equality holds if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
 Acknowledgements: This research was funded in whole or in part by the Austrian Science Fund (FWF) doi/10.55776/37030.